### v1

Abstract
We ask whether a stationary lattice in dimension $d$ whose points are shifted by identically distributed but possibly dependent perturbations remains hyperuniform. When $d=1$ or $2$, we show that it is the case when the perturbations have a finite $d$-moment, and that this condition is sharp. When $d\geq 3$, we construct arbitrarily small perturbations such that the resulting point process is not hyperuniform. As a side remark of independent interest, we exhibit hyperuniform processes with arbitrarily slow decay of their number variance.
1 Introduction
1.1 Setting and main results
Let $\mathsf{d}\geq 1$ be the ambient dimension. A point configuration is a locally finite collection of points in $\mathbb{R}^{\mathsf{d}}$, possibly with multiplicities, which we identify with a purely atomic Radon measure $\mathbf{X}$ (see Section 2.3 for technical details). For $r>0$, we let $\mathrm{B}_{r}$ be the closed Euclidean ball of radius $r$ centered at the origin and if $\mathbf{X}$ is a point configuration we denote by $|\mathbf{X}\cap\mathrm{B}_{r}|$ the number of its points in $\mathrm{B}_{r}$. The following definition has received a lot of interest in the past twenty years:
Let $\mathbf{X}$ be a random point configuration (i.e. a point process). We say that $\mathbf{X}$ is hyperuniform when:
${\sigma(r):=\frac{\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}\longrightarrow 0}\text{ as }r\to\infty,$ (1.1)
where $|\mathrm{B}_{r}|$ denotes the volume of a ball of radius $r$. The quantity $\mathrm{Var}[|\mathbf{X}\cap\mathrm{B}_{r}|]$ is called the number variance (of the random point configuration $\mathbf{X}$, in $\mathrm{B}_{r}$) and $\sigma$ the rescaled number variance. Of course, the volume $|\mathrm{B}_{r}|$ is proportional to $r^{\mathsf{d}}$ so when (1.1) holds one can wonder how slow the number variance growth actually is compared to $r^{\mathsf{d}}$. In the terminology of [Tor18], $\mathbf{X}$ is said to be hyperuniform of class I when in fact:
$\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]=\mathcal{O}(r^{\mathsf{d}-1})\text{ as }r\to\infty.$ (1.2)
The basic example of a class I-hyperuniform point process is the “stationary lattice” $\mathbf{L}:=\{x+U,\ x\in\mathcal{L}\}$, where $\mathcal{L}$ is some lattice in $\mathbb{R}^{\mathsf{d}}$ (i.e. $\mathbb{Z}^{\mathsf{d}}$ or the image of $\mathbb{Z}^{\mathsf{d}}$ by an affine transformation), and the random variable $U$ is uniformly distributed in a fundamental domain of $\mathcal{L}$. Since (1.2) is the slowest possible growth, stationary lattices are “as hyperuniform as it gets”. In this paper, we consider perturbed lattices of the form $\mathbf{PL}:=\{x+U+\mathrm{p}_{x},\ x\in\mathcal{L}\}$ where $U$ is as above while $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ is a family of random variables independent of $U$, thought of as “perturbations”, with values in $\mathbb{R}^{\mathsf{d}}$, and we discuss whether $\mathbf{PL}$ remains hyperuniform or not.
When the perturbations are independent and identically distributed, the answer is well-known: in any dimension, if $\mathrm{p}_{0}$ has a finite first moment, then $\mathbf{PL}$ remains hyperuniform of class I, see [GS75]. In fact, even if $\mathrm{p}_{0}$ has only a finite moment of order $10<s<$ 1, then $\mathbf{PL}$ remains hyperuniform (but not necessarily of class I).
Throughout this paper, we make the following assumption: the perturbations $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ are identically distributed, and moreover their joint distribution is $\mathcal{L}$-invariant in the sense that for all $x_{\star}$ in $\mathcal{L}$, $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ and $(\mathrm{p}_{x+x_{\star}})_{x\in\mathcal{L}}$ have the same distribution. It implies that the perturbed lattice $\mathbf{PL}$ itself is stationary.
We may now state our main results. Let us emphasize that we allow the perturbations to be dependent.
Theorem 1 ((Non)-hyperuniformity of perturbed lattices).
  • For all $\mathsf{d}\geq 1$, if $\mathbb{E}[|\mathrm{p}_{0}|^{\mathsf{d}}]$ is finite, then the number variance of $\mathbf{PL}$ in $\mathrm{B}_{r}$ is finite for every $r>0$.
  • For $\mathsf{d}=1$ or $2$, if $\mathbb{E}[|\mathrm{p}_{0}|^{\mathsf{d}}]$ is finite, then $\mathbf{PL}$ is hyperuniform in the sense of (1.1).
  • For all $\mathsf{d}\geq 1$, taking $\mathcal{L}=\mathbb{Z}^{\mathsf{d}}$, if $X$ is a positive random variable such that $\mathbb{E}[X^{\mathsf{d}}]=+\infty$, then there exist perturbations $(\mathrm{p}_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$, with $|\mathrm{p}_{0}|\leq X$ almost surely, such that the resulting perturbed lattice has infinite number variance in finite balls.
  • For all $\mathsf{d}\geq 3$, taking $\mathcal{L}=\mathbb{Z}^{\mathsf{d}}$, for all $\varepsilon>0$ there exist perturbations $(\mathrm{p}_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$ with $|\mathrm{p}_{0}|\leq\varepsilon$ almost surely, such that the resulting perturbed lattice satisfies $\lim_{r\to+\infty}\frac{\mathrm{Var}\left[|\mathbf{PL}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=+\infty$.
The third item shows that in dimension $1$ and $2$ the sufficient condition of the second item allowing to retain hyperuniformity is particularly sharp, because anything below that threshold can give rise to processes which not only are not hyperuniform, but have in fact infinite number variance in finite balls. In higher dimensions ($d\geq 3$) the last item shows that arbitrarily small perturbations can break hyperuniformity.
Theorem 2 (About class-I hyperuniformity).
  • For $\mathsf{d}=1$, if $\mathbb{E}[|\mathrm{p}_{0}|^{2}]$ is finite, then $\mathbf{PL}$ is class I hyperuniform (which, in dimension $1$, means that the number variance stays bounded as the size of the interval goes to +$\infty$+ ∞). This condition is sharp.
  • For $\mathsf{d}=2$, there are perturbations of $\mathbb{Z}^{2}$ which are almost surely bounded by an arbitrarily small constant, and for which the perturbed lattice is not class I hyperuniform.
We thus obtain with the first item a sharp sufficient condition for class-I hyperuniformity of perturbed lattices in dimension $1$ based only on the size of the perturbations. The second item indicates that in dimension $2$ (or higher) this is not possible - any sufficient condition should presumably involve the correlation structure of the perturbations, too (see Section 1.5).
Theorem 3 (Arbitrarily slow decay of the number variance).
Let $\tilde{\sigma}:(0,+\infty)\to(0,+\infty)$ be a non-increasing positive function with $\lim_{r\to+\infty}\tilde{\sigma}(r)=0$. For all $\mathsf{d}\geq 1$, there exist $c>0$ and a perturbed lattice $\mathbf{PL}$ such that:
  1. 1.
    $\mathbf{PL}$ is hyperuniform.
  2. 2.
    The rescaled number variance $\sigma$ of $\mathbf{PL}$ (see (1.1)) satisfies:
    $\sigma(r)\geq c\tilde{\sigma}(r)\text{ for all }r\geq 1.$ (1.3)
We emphasize that $\tilde{\sigma}$ can really be chosen tending to 00 as slow as one wants.
1.2 About hyperuniformity
General facts
In the statistical physics literature of the 1980’s (see e.g. [MY80, Leb83]) it has been observed that some classical systems of particles with long-range interactions exhibit cancellation of charge fluctuations, namely that the variance of the number of charges in a domain increases proportionally to the surface area of the domain’s boundary, and not to the volume. The notion of hyperuniformity, or equivalently of super-homogeneous distributions of points was introduced in [GJL02, TS03], with applications in theoretical chemistry and physics. We refer to the survey by S. Torquato [Tor18] for an overview of the considerable literature that has since been devoted to this topic.
One motivation to study hyperuniform processes is that they can exhibit a form of “order within disorder”: the fact that their rescaled number variance goes to 00 can be interpreted as a form of collective order yet their truncated correlation functions may all decay very fast at infinity hence there is no “long-range order” in the usual sense of statistical physics. It is worth mentioning that since the “number of points in big boxes” is a very natural observable to study, some “hyperuniformity” results can predate the coining of the term itself by decades.
Amusingly, the original definition for hyperuniformity was more stringent than (1.1) and corresponds to what is now called “class-I hyperuniformity” (see (1.2)), as it has been realized that many interesting processes satisfy (1.1) but not (1.2). The defining condition (1.1) should be understood by comparison to the fundamental example of a random point configuration given by the Poisson point process. Recall that for a Poisson point process $\mathbf{X}$ (with constant intensity $c$), the number of points $|\mathbf{X}\cap\mathrm{B}_{r}|$ follows a Poisson distribution of parameter $c|\mathrm{B}_{r}|$ for all $r>0$, and as a straightforward consequence the number variance is precisely proportional to the volume $|\mathrm{B}_{r}|$. For a random point configuration $\mathbf{X}$ to be hyperuniform thus means that its number variance (in large balls) is asymptotically “much smaller than for a Poisson point process”.
Hyperuniformity and structure factor
Assume that the (stationary) point process $\mathbf{X}$ has intensity $1$ (i.e. one point per unit volume on average) and is nice enough so that the following quantity:
$g_{2}(x):=\lim_{\varepsilon\to 0}\frac{1}{|\mathrm{B}_{\varepsilon}|^{2}}\mathbb{P}[\mathbf{X}\text{ has at least one point in }\mathrm{B}_{\varepsilon}(0)\text{ and in }\mathrm{B}_{\varepsilon}(x)]$
exists for all $x\in\mathbb{R}^{\mathsf{d}}$. The map $x\mapsto g_{2}(x)$ is usually called the two-point correlation function of $\mathbf{X}$. The structure factor $S$ is then defined as: $S:=1+\widehat{(g_{2}-1)}$, where $\widehat{\cdot}$ means the Fourier transform (see section 2.1 for definition and conventions).
Heuristically speaking, there is a correspondence between the small wave-number behavior of the structure factor (i.e. $S(k)$ for $|k|$ small) and the growth of number variance in large balls (see for instance [Tor18, Sec. 2]), which can be summarized as follows:
  • $S(0)=0$longleftrightarrow $\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]=o(r^{\mathsf{d}})$ ($\mathbf{X}$ is hyperuniform). This is sometimes phrased as the condition “$\lim_{k\to 0}S(k)=0$”, or as the sum rule $\int_{\mathbb{R}^{\mathsf{d}}}(g_{2}-1)=-1$.
  • $S(k)\sim|k|^{\alpha}$ with $\alpha>1$longleftrightarrow $\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]=\mathcal{O}(r^{\mathsf{d}-1})$ ($\mathbf{X}$ is class-I hyperuniform)
  • $S(k)\sim|k|^{\alpha}$ with $\alpha\in(0,1)$longleftrightarrow $\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]=\mathcal{O}(r^{\mathsf{d}-\alpha})$ ($\mathbf{X}$ is hyperuniform, but not Class-I)
On a mathematical level, there are several aspects to care about:
  1. 1.
    The two-point correlation function is in general not a proper function and has to be defined as a measure (we recall this in Section 2.3 below). The measure $g_{2}-\mathtt{d}x$ might have infinite mass, so its Fourier transform (and thus the structure factor itself) has to be defined in a distributional sense.
  2. 2.
    The value of $S$ at 00 might not be well-defined, or $S$ might not be continuous near 00. As a matter of fact, it can happen that the statement $\lim_{|k|\to 0}S(k)=0$ is false (because the limit does not exist) and yet the point process is hyperuniform. This is not a big issue in itself - in actual computations related to hyperuniformity, the structure factor always gets integrated against some “nice” functions.
  3. 3.
    In order to know that only the behavior of $S$ near/at 00 matters for hyperuniformity, one should place some integrability conditions on the covariance structure of the point process, as discussed in [Cos21, Sec. 2.2]. Those conditions are satisfied by many natural processes (e.g. statistical physics systems with fast decay of correlations) but not by lattices.
Such considerations are apparent in the mathematical literature on the topic, see e.g. [AGL21, Thm. 1] where an assumption is placed on the regularity of the structure factor near 00, and a certain integral condition on $S$ is stated, that takes into account the behavior away from 00.
1.3 Connection with the literature
Positive result
Our main positive result in Theorem 1 is the fact that in dimension $2$, perturbations with finite second moment preserve the hyperuniformity of lattices. To the best of our knowledge, this was not proven in the mathematical literature even for bounded perturbations. In the physics literature, this is discussed in several places including [GJL02, Sec. VI.C], [Gab04], [KT18, Sec. IV. A] - all of which do state that hyperuniformity of a two-dimensional lattice cannot be broken if the perturbations have finite variance. Their argument relies on computing a small wave-number approximation for the structure factor of $\mathbf{PL}$, in which the covariance structure of the perturbations plays an important role.
This is detailed in [Gab04], which includes the case of Gaussian perturbations (note that the paper also considers the perturbations of general hyperuniform processes, not only stationary lattices). Two difficulties arise when trying to make this argument mathematically rigorous:
  • A minor issue is that the “structure factor” for lattices or perturbed lattices is not well-defined as a function and has to be understood instead in some distributional sense as mentioned above.
  • Justifying the small wave-number expansion is fairly straightforward, but showing that the contribution of “non small” wave-numbers is indeed negligible requires a significant effort.
Examples of pathological processes
Our examples (see again Theorem 1) of non-hyperuniform perturbed lattices in dimension $\mathsf{d}=2$ (with perturbations failing to have a finite second moment) and $\mathsf{d}\geq 3$ (with arbitrarily small perturbations) seem to be the first mathematical constructions of such objects. Their existence was hinted on in the physics papers cited above (e.g. in [KT18] “In contrast to uncorrelated perturbed lattices that are always hyperuniform, correlated ones can be non-hyperuniform.”), but no explicit construction was given - rather, based on the small wave-number expansion for the structure factor mentioned above, they argue that if the spectral density of the perturbation field has a certain power-law near 00, then hyperuniformity will be broken. In dimension $2$, this does not suggest how to proceed (because in that case, the perturbations must have infinite variance, hence no properly defined spectral density). In higher dimensions, this suggests that certain Gaussian perturbation fields with well-chosen spectral densities can do the job, but one would still have to justify that the small wave-number expansion does actually govern the hyperuniformity status of perturbed lattices (which, as explained in the previous paragraph, is not straightforward), and - for bonus points - to produce counter-examples with perturbations that are bounded and not “only” with finite moments of all order like Gaussian variables.
Our constructions are elementary and allow us to prove lower bounds on the number variance directly, without relying on the structure factor.
Slow decay of the rescaled number variance
In Theorem 3 we give what seems to be the first examples of hyperuniform processes in any dimension with truly arbitrarily slow decay of the rescaled number variance (although this might have been “folklore” knowledge to some experts, the only general result that we are aware of is the fact that in dimension $1$, any power law decay can be realized by a well-chosen determinantal point process, see [Sos00, Sec. 3]). For his phenomenological classification, Torquato [Tor18] considers three asymptotic rates: class-I ($\sigma(r)\sim cr^{-1}$), class-II ($\sigma(r)\sim cr^{-1}\log r$) and class-III ($\sigma(r)\sim cr^{-\alpha}$ for some $\alpha\in(0,1)$), which does not include cases like $\sigma(r)\geq c\log^{-0.01}r$. We thus show that this classification does not quite cover the entire spectrum of possible behaviors. It is however fair to say that our examples are ad hoc constructions with little “physical” reality.
1.4 An optimal transport perspective
In this section, we reformulate our main result from the point of view of optimal transport between stationary point processes recently introduced in [EHJM23] (using their notation).
Let us recall the relevant notions. Denote the set of locally finite point configurations by $\Gamma$. For two distributions of stationary point processes $\mathsf{P}_{0}$ and $\mathsf{P}_{1}$ and a cost function $c:\Gamma\times\Gamma\to\mathbb{R}$ we are interested in the transport problem
$\inf_{Q\in\mathsf{Cpl}_{s}(\mathsf{P}_{0},\mathsf{P}_{1})}\int c\mathsf{d}Q$ (1.4)
where $\mathsf{Cpl}_{s}(\mathsf{P}_{0},\mathsf{P}_{1})$ is the set of all couplings between $\mathsf{P}_{0}$ and $\mathsf{P}_{1}$ that are invariant under the diagonal action of $\mathbb{R}^{d}$ on $\Gamma\times\Gamma$ given by $\theta_{x}(\xi,\eta)(\cdot,\cdot)=(\xi,\eta)(\cdot-x,\cdot-x)$ (for all $x\in\mathbb{R}^{d}$).
Note that in (1.4), $c$ is a cost function between point configurations. A particularly relevant choice of $c$ is the following: for $p\geq 1$ and for any two point configurations $\xi,\eta$ on $\mathbb{R}^{\mathsf{d}}$, we put:
$c_{p}(\xi,\eta):=\inf_{q\in\mathsf{cpl}(\xi,\eta)}\limsup_{n\to\infty}\frac{1}{n^{d}}\int_{\Lambda_{n}\times\mathbb{R}^{d}}|x-y|^{p}q(dx,dy),$
where $\mathsf{cpl}(\xi,\eta)$ denotes the set of all couplings between the point configurations $\xi$ and $\eta$, seen as measures on $\mathbb{R}^{\mathsf{d}}$. The quantity $c_{p}(\xi,\eta)$ can be thought of as a “$p$-Wasserstein distance per unit volume” between the measures $\xi$ and $\eta$.
Then, if $\mathsf{P}_{0},\mathsf{P}_{1}$ are the distributions of two random point configurations, we define their “$p$-Wasserstein distance” $\mathrm{Wass}_{p}$ by:
$\mathrm{Wass}_{p}^{p}(\mathsf{P}_{0},\mathsf{P}_{1}):=\inf_{Q\in\mathsf{Cpl}_{s}(\mathsf{P}_{0},\mathsf{P}_{1})}\int c_{p}\mathsf{d}Q.$
This defines a geodesic metric on the space of stationary point process with unit intensity.
In order to connect this definition of a Wasserstein distance between point processes with the result of the present paper, it is relevant to mention that there is a representation of $\mathrm{Wass}_{p}$ in terms of equivariant random matchings. Let $(\xi^{\bullet},\eta^{\bullet})$ be two random point configurations whose distribution is jointly invariant under the shifts $\theta_{x}$ (for $x\in\mathbb{R}^{d}$), denote their marginal distributions by $\mathsf{P}_{0}$, $\mathsf{P}_{1}$. Assume that $(\xi^{\bullet},\eta^{\bullet})$ are constructed on some probability space $(\Omega,\mathcal{F},\mathbb{P})$ equipped with a measurable flow $\theta:\Omega\to\Omega$ satisfying $\theta_{0}=id,\theta_{x}\circ\theta_{y}=\theta_{x+y}.$ Denote by $\mathsf{cpl}_{e}(\xi^{\bullet},\eta^{\bullet})$ the set of all equivariant couplings of $\xi^{\bullet}$ and $\eta^{\bullet}$, i.e. for each $\omega\in\Omega$ the measure $q^{\omega}$ is a coupling of $\xi^{\omega}$ and $\eta^{\omega}$ and $q^{\theta_{x}\omega}(A,B)=q^{\omega}(A-x,B-x)$. Then, we have [EHJM23, Prop. 1.2]
$\mathrm{Wass}_{p}^{p}(\mathsf{P}_{0},\mathsf{P}_{1})=\inf_{(\xi^{\bullet},\eta^{\bullet})}\inf_{q^{\bullet}\in\mathsf{cpl}_{e}(\xi^{\bullet},\eta^{\bullet})}\mathbb{E}\left[\int_{[0,1]^{\mathsf{d}}\times\mathbb{R}^{d}}|x-y|^{p}\mathsf{d}q(x.y)\right],$ (1.5)
where the first infimum runs over all probability spaces supporting jointly invariant random measures $\xi^{\bullet}\sim\mathsf{P}_{0},\eta^{\bullet}\sim\mathsf{P}_{1}.$ Furthermore, the infimum is attained by an invariant matching $\mathrm{m}^{\bullet}$, i.e. a coupling of the form $(id,T^{\bullet})_{\#}\xi^{\bullet}$, where $S_{\#}\mu$ denotes the push-forward of the measure $\mu$ by the measurable map $S$.
Now comes the important point: Note that $T^{\bullet}(x)-x$ is a stationary displacement field and any stationary displacement field induces a stationary matching. Moreover, if $\xi=\sum_{x\in\mathbb{Z}^{\mathsf{d}}}\delta_{x}$ the cost of a matching $m^{\bullet}=(id,T^{\bullet})_{\#}\xi^{\bullet}$ is given by $\mathbb{E}[|T(0)-0|^{p}]$, the $L^{p}$ norm of the displacement field.
Let us denote here by $\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}}$ the stationary version of the integer lattice $\mathbb{Z}^{\mathsf{d}}$, i.e. $\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}}:=\mathrm{Law}(\sum_{x\in\mathbb{Z}^{\mathsf{d}}}\delta_{x+U})$, with a random variable $U$ uniformly distributed in $[0,1]^{\mathsf{d}}$. It is equivalent to find:
  • A $c_{p}$-optimal equivariant matching between $\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}}$ and $\mathsf{P}$.
  • A stationary displacement field $(p_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$ with minimal $L^{p}$ norm $\mathbb{E}[|p(0)|^{{p}}]$ such that the law of the resulting perturbed lattice $\mathbf{PL}$ coincides with $\mathsf{P}$.
With this language, the results of Theorem 1 and Theorem 2 can be rephrased as follows:
Theorem.
Let $\mathsf{P}$ be the law of a stationary point process $\mathbf{X}$ on $\mathbb{R}^{\mathsf{d}}$. We have:
  1. 1.
    If $\mathsf{d}=1,2$ and if $\mathrm{Wass}_{\mathsf{d}}(\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}},\mathsf{P})$ is finite, then $\mathsf{P}$ is hyperuniform.
  2. 2.
    If $\mathsf{d}=1,2$, for all $\varepsilon>0$, there are examples with $\mathrm{Wass}_{\mathsf{d}-\varepsilon}(\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}},\mathsf{P})$ finite and yet $\mathsf{P}$ is not hyperuniform (in fact the number variance in finite balls is not even finite).
  3. 3.
    If $\mathsf{d}\geq 3$, for all $\varepsilon>0$, there are examples with $\mathrm{Wass}_{\infty}(\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}},\mathsf{P})\leq\varepsilon$ and yet $\mathsf{P}$ is not hyperuniform.
  4. 4.
    If $\mathsf{d}=1$ and if $\mathrm{Wass}_{2}(\mathsf{P}^{\mathbb{Z}},\mathsf{P})$ is finite, then $\mathsf{P}$ is class-I hyperuniform. If $\mathsf{d}=2$, for all $\varepsilon>0$, there are examples with $\mathrm{Wass}_{\infty}(\mathsf{P}^{\mathbb{Z}^{2}},\mathsf{P})\leq\varepsilon$ and yet $\mathsf{P}$ is not class-I hyperuniform.
Since $\mathrm{Wass}_{p}$ is a geodesic metric, it follows that in dimensions $\mathsf{d}=1,2$ the set of point processes with finite $\mathrm{Wass}_{p}$ distance to $\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}}$ is geodesically closed, and thus the geodesic interpolation between two hyperuniform point processes with finite $\mathrm{Wass}_{\mathsf{d}}$ distance to $\mathsf{P}^{\mathbb{Z}^{\mathsf{d}}}$ stays hyperuniform (resp. class-I hyperuniform for $\mathsf{d}=1$ and $\mathrm{Wass}_{2}$).
Remark 1.1.
One can more generally define the Wasserstein distance between Radon measures (not necessarily purely atomic ones). In that sense, a stationary lattice is always at finite $\mathrm{Wass}_{p}$ distance of the Lebesgue measure (seen as a deterministic Radon measure) for all $p$ (in fact for $p=+\infty$). It is thus equivalent to be “at finite distance of a stationary lattice” or “at finite distance of Lebesgue”.
1.5 Open questions and connection with recent results
Perturbations of general hyperuniform processes
Proving or disproving hyperuniformity of stationary perturbations of a general hyperuniform process (not only a stationary lattice) remains open. In fact, as we discuss below in Section 2.4, even the correct notion of “perturbations” is not completely clear in general: for example, do we allow perturbations to depend on the realization of the underlying process or not? In [Gab04], the case of perturbations that are independent from the process is considered.
For perturbations that can be correlated with the original point process, one can make the following observation: if $\mathbf{X}$ is hyperuniform and has finite $2$-Wasserstein distance to a lattice, then any point process $\mathbf{X}^{\prime}$ with finite $2$-Wasserstein distance to $\mathbf{X}$ is hyperuniform again, by the triangle inequality and our main result. In other words, any $L^{2}$ perturbation of $\mathbf{X}$ (independent of $\mathbf{X}$ or not) will yield an hyperuniform process. Note that this does not completely answer the question because, as observed in [HL24], not every hyperuniform process has finite $2$-Wasserstein distance to a lattice.
Converse statements in dimension $2$: the results of [LRY24, BDGZ24, HL24]
Using the language introduced just above, our main result in the present paper is that in dimension $2$, finite $2$-Wasserstein distance to a (stationary) lattice implies hyperuniformity. Several recent papers are dealing (among other things) with the converse statement.
In [LRY24], the converse statement (hyperuniformity $\implies$ finite $2$-Wasserstein distance) was shown to be true under a certain integrability condition on the reduced pair correlation function of the process which, they argue, proves that “most hyperuniform planar point processes are $L^{2}$-perturbed lattices”.
In [BDGZ24], finite $2$-Wasserstein distance is obtained under a quantitative assumption on the decay of the rescaled number variance which is just slightly stronger than hyperuniformity itself.
In a companion [HL24] to the present paper, the integrability condition of [LRY24] is observed to be closely related to some kind of “finiteness of the Coulomb energy” and it is shown that indeed:
[…]
(the second implication being the one from the present paper). However, simple examples are given in [HL24] of stationary point processes which are hyperuniform, but not at finite $2$-Wasserstein distance to a lattice - thus there is no perfect equivalence between those two notions. It is also proven that under a condition which is slightly stronger than hyperuniformity but slightly weaker than the one of [BDGZ24], one obtains finite $2$-Wasserstein distance to a lattice. Since this condition covers by far the vast majority of known examples, it completely agrees with the statement of [LRY24] cited above.
The decay condition of [HL24], called $\star$-hyperuniformity”, is shown there to be equivalent to finite Coulomb energy. In the present paper, as a corollary of the proof of Theorem 1, we obtain quantitative bounds on the decay of the rescaled number variance. In particular, for all $\varepsilon>0$ (see Corollary 3.8):
[…]
Sufficient condition for class-I hyperuniformity
In dimension $2$ we could not produce examples of perturbed lattices which are a) hyperuniform but not class-I hyperuniform and b) such that the covariance structure of the perturbations satisfy a summability condition of the type:
$\sum_{x\in\mathcal{L}}\left|\mathrm{Cov}(\mathrm{p}_{x},\mathrm{p}_{0})\right|<+\infty.$ (1.6)
It is well-known that (1.6) implies that the spectral density of the perturbations not only exists but is continuous, which eliminates certain “bad” behaviors, and could be enough to ensure that the resulting perturbed lattice is class-I hyperuniform.
To see why, recall that the basic plan to break hyperuniformity (see e.g. [KT18, eq. (34)]) is to choose perturbations $(\mathrm{p}_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$ whose spectral density $\rho$ satisfies $\rho(\omega)\sim C|\omega|^{-2}$ as $\omega\to 0$. Indeed, when doing the computations (see e.g. (3.19) below), the behavior near 00 of the spectral density of the perturbations turns out to play a major role for the asymptotic decay of the rescaled number variance, and $\rho(\omega)\sim C|\omega|^{-2}$ is in some sense the “critical” case separating hyperuniform and non-hyperuniform behaviors. Having (for the perturbations) $\rho(\omega)\sim C|\omega|^{-2+\delta}$ with $\delta>0$ produces a perturbed lattice which is hyperuniform, but for $\delta$ small it ought to be close to the “non-hyperuniform” case and in particular “not very hyperuniform” (type-I).
In dimension $2$, it is not possible to have $\rho(\omega)\sim C|\omega|^{-2}$ near 00 because $\rho$ would fail to be integrable, but one can certainly produce Gaussian perturbation fields with $\rho(\omega)\sim C|\omega|^{-2+\delta}$ for any $\delta>0$, and the resulting perturbed lattice will presumably not be class-I hyperuniform if $\delta\leq 1$. However, if the spectral density $\rho$ is known to be continuous near 00, such divergences are of course forbidden (one needs $\delta\geq 2$). We are thus left with the following question: in dimension $2$, if $\mathbb{E}[|\mathrm{p}_{0}|^{2}]<+\infty$ and (1.6) holds, is $\mathbf{PL}$ necessarily class-I hyperuniform?
1.6 Plan of the paper
In Section 2, we introduce our notation, the setting of perturbed lattices and point processes, and the initial results used for the computation of the number variance (such as correlation measures, structure factors, Bessel kernels, etc.).
Section 3 is devoted to the proof of the main result of the paper: in dimension $\mathsf{d}=2$, a finite second moment of perturbations ensures hyperuniformity. The proof is almost complete, except for the delicate Fourier computations needed for Proposition 3.7, which are postponed to Appendix A.
The analogous positive result in dimension $\mathsf{d}=1$ is simpler to prove and is provided in Appendix B, along with the proof of the second item of Theorem 2.
The proofs of other results, namely the first, third, and fourth items in Theorem 1, the second item in Theorem 2, and Theorem 3, are of a different nature, involving more probabilistic constructions. These proofs are presented in Section 4.
2 Preliminaries
2.1 Some notation
We write $A\preceq B$, or equivalently $A=\mathcal{O}(B)$, when there exists a constant $C$ depending only on the dimension $\mathsf{d}$, the lattice $\mathcal{L}$ and the law of the perturbations $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ such that $|A|\leq C|B|$. We write $f(x)=o(g(x))$ as $x\to a$ if $\lim_{x\to a}\frac{f(x)}{g(x)}=0$. If the value $a$ is clear from context, we do not mention it.
In the proof of Proposition 3.1, which is our main result, we will work with a large parameter $r$ and auxiliary parameters depending on $r$. In that context, we sometimes write $A\ll B$ to signify that $\frac{A}{B}\to 0$ as $r\to\infty$ ($A$, $B$ being two quantities depending explicitly or implicitly on $r$).
[…]
Fourier transforms
We choose the following convention for the Fourier transform: if $f$ is in $L^{1}(\mathbb{R}^{\mathsf{d}})$ we let:
$\hat{f}(s):=\int_{\mathbb{R}^{\mathsf{d}}}e^{-is\cdot t}f(t)\mathtt{d}t.$
We also define the characteristic function of a random variable $X$ as $\varphi_{X}(s):=\mathbb{E}[e^{-is\cdot X}]$. With this convention we have $\hat{\hat{f}}=(2\pi)f(-\cdot)$. We also have $\widehat{f\ast g}=\hat{f}\hat{g}$ but $\widehat{fg}=(2\pi)^{-1}\hat{f}\ast\hat{g}$.
2.2 Generalities on lattices
Some definitions
We recall some definitions about lattices, which can be found in most textbooks.
The lattices considered here are subgroups of $\mathbb{R}^{\mathsf{d}}$ generated by linear combinations with integer coefficients of the vectors of a certain basis of $\mathbb{R}^{\mathsf{d}}$ (we only consider “full-rank” lattices). The basic example is the so-called integer lattice $\mathbb{Z}^{\mathsf{d}}$ (or “square lattice” for $\mathsf{d}=2$) which is simply the subgroup of all points in $\mathbb{R}^{\mathsf{d}}$ with integer coefficients, or equivalently the set of all integer linear combinations of the vectors of the canonical basis of $\mathbb{R}^{\mathsf{d}}$. Any other lattice can be obtained as an affine transformation of $\mathbb{Z}^{\mathsf{d}}$.
Remark 2.1.
If we could show that any lattice $\mathcal{L}$ is itself a bounded stationary perturbation of $\mathbb{Z}^{\mathsf{d}}$, it would be enough to treat the case $\mathcal{L}=\mathbb{Z}^{2}$ for, say, the positive result of Theorem 1 in dimension $2$. This is indeed true, as they have both finite $\mathrm{Wass}_{2}$ distance to Lebesgue (see Remark 1.1). However, we did not see an “easy” explicit perturbation that one could write down in a stationary fashion to connect $\mathbb{Z}^{\mathsf{d}}$ in order to a general lattice, and the cost of working with general lattices is not too high, so we chose this option.
To a lattice $\mathcal{L}$ and a basis $(\lambda_{1},\dots,\lambda_{\mathsf{d}})$, we associate a “fundamental domain” of $\mathcal{L}$ which is the set:
$\mathcal{D}:=\left\{\sum_{i=1}^{\mathtt{d}}t_{i}\lambda_{i},\ (t_{1},\dots,t_{\mathsf{d}})\in[0,1)^{\mathsf{d}}\right\}.$
The fundamental domain is a compact parallelotope, whose volume does not depend on the choice of the basis and is called the “covolume” of $\mathcal{L}$. In this paper, we will always assume the covolume to be $1$. This is of course only for convenience and in no way a loss of generality. An interesting property is that every vector $t$ of $\mathbb{R}^{\mathsf{d}}$ can be uniquely written as: $t=x+u$, where $x\in\mathcal{L}$ and $u\in\mathcal{D}$. In other words, translating a fundamental domain by all the vectors in $\mathcal{L}$ makes a perfect tiling of $\mathbb{R}^{\mathsf{d}}$.
The dual lattice $\mathcal{L}^{\star}$ of $\mathcal{L}$ is defined as: :=$\mathcal{L}^{\star}:=\left\{k\in\mathbb{R}^{\mathsf{d}},\text{ such that }k\cdot x\in\mathbb{Z}\text{ for all }x\in\mathcal{L}\right\}$. The only important fact for us is that $\mathcal{L}^{\star}$ is indeed a lattice, and that it appears in Poisson’s summation formula, which we recall next.
Poisson’s summation formula
We refer to [SW71, Chap. VII]. If $f:\mathbb{R}^{\mathsf{d}}\to\mathbb{C}$ is in $L^{1}$ (so that $\hat{f}$ is well-defined) and if moreover there exists $\delta>0$ and a constant $C>0$ such that: $|f(x)|\leq\frac{C}{|x|^{\mathsf{d}+\delta}},\quad|\hat{f}(x)|\leq\frac{C}{|x|^{\mathsf{d}+\delta}}\text{ as }|x|\to\infty$, then Poisson’s summation formula holds, namely we have:
$\sum_{x\in\mathcal{L}}f(x)=\sum_{k\in\mathcal{L}^{\star}}\hat{f}(2\pi k).$ (2.1)
Note that this obviously includes the case of $f$ in Schwartz’s space (i.e. when $f$ is smooth and $f$ and all its derivatives decay faster than every polynomial at infinity).
Some lattice point counting
It is easy to see that for $R\geq 1$, the number of lattice points in any ball of radius $R$ is bounded by some constant times $R^{\mathsf{d}}$. We will also need more precise estimates, which we now record in dimension $2$.
Lemma 2.2 (Number of lattice points at a certain distance from $\partial\mathrm{B}_{R}$.).
Let $\mathcal{L}$ be a lattice in $\mathbb{R}^{2}$. For $R\geq 1$, and $d\geq 0$ we denote by $\mathcal{N}_{d,R}$ be the number of lattice points $x\in\mathcal{L}$ such that the distance $\mathrm{dist}(x,\partial\mathrm{B}_{R})\leq d$.
  1. 1.
    (Moderate distances) For $1\leq d\leq R$, we have $\mathcal{N}_{d+1,R}-\mathcal{N}_{d,R}=\mathcal{O}(R)$, hence there are $\mathcal{O}(R)$ lattice points whose distance to the circle $\partial\mathrm{B}_{R}$ is between $d$ and $d+1$.
  2. 2.
    (Small distances) For some $\gamma>0$, the following holds. Let $\tau$ be a number in $(0,1)$ depending on $R$. If $\tau=\tau(R)$ tends to 00 as $R\to\infty$, then $\mathcal{N}_{\tau,R}=\mathcal{O}(\tau R)+\mathcal{O}(R^{1-\gamma})$.
Proof of Lemma 2.2.
The problem of counting lattice points falling inside various shapes is very old. In the case of the so-called “integer” lattice, i.e. $\mathcal{L}=\mathbb{Z}^{2}$, this goes back at least to Gauss, who proved that:
$\mathcal{N}_{R}:=\left|\{x\in\mathbb{Z}^{2},|x|\leq R\}\right|=\pi R^{2}+\mathcal{O}(R)\text{ as $R\to\infty$.}$ (2.2)
The first item follows immediately.
Moreover, with our notation, $\mathcal{N}_{\tau,R}=\mathcal{N}_{R+\tau}-\mathcal{N}_{R-\tau}$ (we are counting points both inside and outside the disk $\mathrm{B}_{R}$) so Gauss’s bound gives $\mathcal{N}_{\tau,R}=4\pi R\tau+\mathcal{O}(R)$, which is not quite sufficient for the second item, but any improvement in the power of $R$ appearing in the error term in (2.2) will do. A $\mathcal{O}(R^{\frac{2}{3}})$ error term was indeed obtained in the early $20^{th}$ century for what had become “Gauss’ circle problem”, which settles our question for the case $\mathcal{L}=\mathbb{Z}^{2}$.
The case of a general lattice can be “mapped” onto a similar counting problem for the integer lattice as follows. Let $T$ be an affine transformation mapping $\mathcal{L}$ onto $\mathbb{Z}^{2}$ and let $\mathcal{E}$ be the ellipsis obtained as the image of the unit disk by $T$. More generally, let $\mathcal{E}_{R}$ be the image of $\mathrm{B}_{R}$ by $T$. Since $\mathrm{B}_{R}$ is the image of $\mathrm{B}_{1}$ by the homothety $R\mathrm{Id}$, we also have $\mathcal{E}_{R}=R\mathrm{Id}(\mathcal{E})$. The question of counting points falling in re-scaled versions of smooth domains has been studied by Van der Corput, Vinogradov and others (see e.g. [Hux96] for an extensive presentation). Their results apply to a re-scaled ellipsis and imply that for all $\varepsilon>0$:
$\mathcal{N}^{\prime}_{R}:=\left|\{x\in\mathbb{Z}^{2},x\in\mathcal{E}_{R}\}\right|=R^{2}\times|\mathcal{E}|+\mathcal{O}(R^{\frac{2}{3}+\varepsilon})\text{ as $R\to\infty$.}$ (2.3)
Of course the implicit multiplicative constant will depend on $\varepsilon$ and on the shape of $\mathcal{E}$ i.e. on the original lattice $\mathcal{L}$. Any fixed $\varepsilon<\frac{1}{3}$ works for us and it directly gives us the first item of the statement. For the second item, recall that we are interested in the following quantity:
$\mathcal{N}_{\tau,R}:=\left|\{x\in\mathcal{L},\mathrm{dist}(x,\partial\mathrm{B}_{R})\leq\tau\}\right|=\left|\{x\in\mathcal{L},x\in\mathrm{B}_{R+\tau}\setminus\mathrm{B}_{R-\tau}\}\right|.$
After applying the transformation $T$ (which is invertible), this is equivalent to counting
$\mathcal{N}^{\prime}_{\tau,R}:=\left|\{x\in\mathbb{Z}^{2},x\in\mathcal{E}_{R+\tau}\setminus\mathcal{E}_{R-\tau}\}\right|=\mathcal{N}^{\prime}_{R+\tau}-\mathcal{N}^{\prime}_{R-\tau}.$
Using (2.3) with any $\varepsilon<\frac{1}{3}$ yields $\mathcal{N}^{\prime}_{\tau,R}=|\mathcal{E}|\times 4\tau R+O(R^{1-\gamma})$ for some $\gamma>0$, which concludes the proof of the lemma. ∎
2.3 Generalities on point processes
Configuration of points
A deterministic configuration of points in $\mathbb{R}^{\mathsf{d}}$ is defined as a locally finite integer-valued measure on $\mathbb{R}^{\mathsf{d}}$, i.e. a measure of the form $\mathbf{x}=\sum_{i}\delta_{x_{i}}$ where $(x_{i})_{i}$ is a finite or infinite sequence of points in $\mathbb{R}^{\mathsf{d}}$ without accumulation points. Note that multiple points are allowed at the same location. The state space of all point configurations is endowed with the $\sigma$-field generated by the counting functions $\mathbf{x}\mapsto\mathbf{x}(B)$ for all Borel sets $B$ in $\mathbb{R}^{\mathsf{d}}$. A random configuration of points, or “point process”, is a random variable on this state space. We refer to [LP17, Sec. 2] for details about multiple points, labelling, etc.
Stationarity.
If $\mathbf{X}$ is a point configuration and $u$ is a vector in $\mathbb{R}^{\mathsf{d}}$, we denote by $\mathbf{X}+u$ the “shift by u” i.e. the effect of translating all the points of $\mathbf{X}$ by $u$. A point process is said to be stationary when its distribution is invariant with respect to all the translations i.e. when $\mathbf{X}+u$ and $\mathbf{X}$ have the same distribution for all $u\in\mathbb{R}^{\mathsf{d}}$.
When $\mathbf{X}$ is stationary, we define its intensity as the average number of points of $\mathbf{X}$ per unit volume. Here for convenience we always work with processes of intensity $1$.
Second order measures
When $\mathbf{X}$ is a random point configuration, we say that $\mathbf{X}$ is locally square integrable if $\mathbb{E}[|\mathbf{X}\cap\mathrm{B}_{r}|^{2}]$ is finite for all $r>0$. If $\mathbf{X}$ is locally square integrable, we can define its reduced second factorial moment measure, denoted by $\alpha_{\mathrm{red}}^{(2)}$. This measure plays a central role in the analysis of hyperuniformity for perturbed lattices because:
  1. 1.
    The measure $\alpha_{\mathrm{red}}^{(2)}$ of a perturbed lattice has a tractable expression, see (2.5).
  2. 2.
    The number variance is conveniently expressed in terms of $\alpha_{\mathrm{red}}^{(2)}$, see (2.10).
We refer to [LP17, Sections 4.4. & 8.2] for details.
Definition 2.3 (Second factorial moment measure).
If $\mathbf{X}$ is a point process, we define its second factorial moment measure as the Borel measure $\alpha^{(2)}$ on $\mathbb{R}^{\mathsf{d}}\times\mathbb{R}^{\mathsf{d}}$ such that for all Borel sets $B_{1},B_{2}$:
$\alpha^{(2)}(B_{1}\times B_{2}):=\mathbb{E}\sum_{x_{1},x_{2}\in\mathbf{X}}^{\neq}\mathbf{1}\left\{x_{1}\in B_{1},x_{2}\in B_{2}\right\}.$
If $\mathbf{X}$ is locally square integrable, then $\alpha^{(2)}$ is well-defined, and if moreover $\mathbf{X}$ is stationary, then we can define its reduced second factorial moment measure $\alpha^{(2)}_{red}$ (sometimes known as pair correlation measure), which is given, for $B$ a Borel set, by:
$\alpha_{red}^{(2)}(B)=\mathbb{E}\sum_{x,y\in\mathbf{X}}^{\neq}\mathbf{1}\{x\in[0,1)^{\mathsf{d}},y-x\in B\}.$ (2.4)
Note that in (2.4) we can replace $[0,1)^{\mathsf{d}}$ by any Borel set of volume $1$ (see e.g. the proof [LP17, Prop. 8.7]).
2.4 Basic properties of perturbed lattices
Discussion of the notion of “perturbations”
Let $\mathcal{L}$ be a lattice and $\mathcal{D}$ a fundamental domain of $\mathcal{L}$. Recall that we define the (unperturbed) stationary lattice as $\mathbf{L}:=\{x+U,x\in\mathcal{L}\}$ where $U$ is uniform in $\mathcal{D}$. One could think of at least two ways to define “perturbations” of $\mathbf{L}$:
  1. 1.
    Let $\mathrm{p}$ be a random field $\mathcal{L}\mapsto\mathbb{R}^{\mathsf{d}}$ the value of $\mathrm{p}_{x}$ for $x\in\mathcal{L}$ corresponds to the “perturbation” applied to $x$. Let $U$ be uniform in $\mathcal{D}$ and independent of $\mathrm{p}$. We define the perturbed stationary lattice $\mathbf{L}$ as the random collection of points given by $\mathbf{PL}:=\{(x+\mathrm{p}(x))+U,\ x\in\mathcal{L}\}$.
  2. 2.
    Let $\bar{\mathrm{p}}$ be a random vector field on $\mathbb{R}^{\mathsf{d}}$, thought of as a “displacement field” - a point placed at $x\in\mathbb{R}^{\mathsf{d}}$ gets displaced by $\bar{\mathrm{p}}(x)$. Let $U$ be uniform in $\mathcal{D}$ and independent of $\bar{\mathrm{p}}$. We define the perturbed stationary lattice $\mathbf{L}$ as the random collection of points given by $\mathbf{PL}^{\prime}:=\{(x+U)+\bar{\mathrm{p}}(x+U),\ x\in\mathcal{L}\}$.
It is straightforward to check that, in the stationary case, those two notions coincide.
Claim 2.4.
With $\mathbf{PL}$ and $\mathbf{PL}^{\prime}$ as defined above:
  1. 1.
    Assume that $\mathrm{p}$ is $\mathcal{L}$-invariant. For $t\in\mathbb{R}^{\mathsf{d}}$, write $t=x+U$ with $x\in\mathcal{L}$ and $U$ in $\mathcal{D}$, and set $\bar{\mathrm{p}}:=\mathrm{p}(t-U)$. Then $\bar{\mathrm{p}}$ is a $\mathbb{R}^{\mathsf{d}}$-invariant displacement field and $\mathbf{PL}^{\prime}=\mathbf{PL}$.
  2. 2.
    Assume that $\bar{\mathrm{p}}$ is $\mathbb{R}^{\mathsf{d}}$-invariant and set $\mathrm{p}:=\bar{\mathrm{p}}_{|\mathcal{L}}$. Then $\mathrm{p}$ is a $\mathcal{L}$-invariant perturbation field and $\mathbf{PL}=\mathbf{PL}^{\prime}$.
In the introduction, we chose the first option, but for our constructions in Section 4 it will be convenient to use the second option.
If one wanted to perturb hyperuniform point processes which are not lattices, the “displacement field” approach seems more tractable as it does not rely on an explicit labelling of the points. In that case however, the assumption of $\bar{\mathrm{p}}$ being independent of $\mathbf{L}$ could be seen as restrictive (it is not very much so for lattices).
Remark 2.5.
The properties of $\mathbf{PL}$ are of course invariant under a global shift, thus we can and will assume that the perturbations (which are identically distributed) are centered, i.e. $\mathbb{E}[\mathrm{p}_{0}]=0$.
Basic properties
Lemma 2.6.
If the perturbations are $\mathcal{L}$-invariant as assumed in Section 1, then:
  1. 1.
    $\mathbf{PL}$ is almost surely locally finite.
  2. 2.
    $\mathbf{PL}$ is a stationary point process.
Proof.
For the first item, take $r\geq 10$ and bound $\mathbb{E}[|\mathbf{PL}\cap\mathrm{B}_{r}|]$ by:
$\mathbb{E}[|\mathbf{PL}\cap\mathrm{B}_{r}|]\leq\mathcal{O}(r^{\mathsf{d}})+\sum_{x\in\mathcal{L},|x|\geq 100r}\mathbb{P}\left[\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\right].$
Let $C$ be an upper bound on the diameter of the fundamental domain $\mathcal{D}$. We see that
[$\mathbb{P}\left[\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\right]\leq\mathbb{P}\left[\{\mathrm{p}_{x}\in\mathrm{B}_{r+C}(-x)\}\right]=\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(-x)\}\right],$
using the fact that the perturbations are identically distributed. Finally, it remains to observe that:
$\sum_{x\in\mathcal{L},|x|\geq 100r}\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(-x)\}\right]\preceq r^{\mathsf{d}}$
is finite, thus $\mathbb{E}[|\mathbf{PL}\cap\mathrm{B}_{r}|]$ is finite, hence $\mathbf{PL}$ is almost surely locally finite.
For the second item, take $t\in\mathbb{R}^{\mathsf{d}}$ and decompose $U+t$ as $x_{\star}+v$ with $x_{\star}\in\mathcal{L}$ and $v\in\mathcal{D}$ (see Section 2.2). If $U$ is uniformly distributed in $\mathcal{D}$, then so is $v$. We then have, using the $\mathcal{L}$-invariance of $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ the following identities in law:
$\{x+U+\mathrm{p}_{x}+t\}_{x\in\mathcal{L}}=\{x+\mathrm{p}_{x}+x_{\star}+v\}_{x\in\mathcal{L}}=\{(x+x_{\star})+\mathrm{p}_{x+x_{\star}}+v\}_{x\in\mathcal{L}}=\{x+\mathrm{p}_{x}+v\}_{x\in\mathcal{L}}.$
First finiteness result
Lemma 2.7.
If $\mathbb{E}[|\mathrm{p}_{0}|^{\mathsf{d}}]$ is finite, then $\mathbf{PL}$ is locally square integrable in the sense of Section 2.3.
Proof of Lemma 2.7.
We write, with $U$ uniformly distributed over the fundamental domain of $\mathcal{L}$:
$\mathbb{E}[|\mathbf{PL}\cap\mathrm{B}_{r}|^{2}]=\mathbb{E}\left(\sum_{x\in\mathcal{L}}\mathbf{1}\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\right)^{2}=\mathbb{E}[\sum_{x,x^{\prime}\in\mathcal{L}}\mathbf{1}\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\times\mathbf{1}\{x^{\prime}+\mathrm{p}_{x^{\prime}}+U\in\mathrm{B}_{r}\}]\\ \leq\sum_{x,x^{\prime}\in\mathcal{L}}\min\left(\mathbb{P}\left[\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\right],\mathbb{P}\left[\{x^{\prime}+\mathrm{p}_{x^{\prime}}+U\in\mathrm{B}_{r}\}\right]\right).$
Let $C$ be an upper bound on the diameter of the fundamental domain $\mathcal{D}$. Since the indices $x,x^{\prime}$ play symmetrical roles we can write:
$\sum_{x,x^{\prime}\in\mathcal{L}}\min\left(\mathbb{P}\left[\{x+\mathrm{p}_{x}{+U}\in\mathrm{B}_{r}\}\right],\mathbb{P}\left[\{x^{\prime}+\mathrm{p}_{x^{\prime}}+U\in\mathrm{B}_{r}\}\right]\right)\leq 2\sum_{x,x^{\prime}\in\mathcal{L},|x^{\prime}|\leq|x|}\mathbb{P}\left[\{x+\mathrm{p}_{x}+U\in\mathrm{B}_{r}\}\right]\\ \preceq\sum_{x\in\mathcal{L}}|x|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{x}\in\mathrm{B}_{r+C}(-x)\}\right].$
Using the symmetry of $\mathcal{L}$ under $x\mapsto-x$ and the fact that the random variables $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ are identically distributed, we also have:
$\sum_{x\in\mathcal{L}}|x|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{x}\in\mathrm{B}_{r+C}(-x)\}\right]=\sum_{x\in\mathcal{L}}|x|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(x)\}\right].$
The contribution to the sum coming from $\{x\in\mathcal{L},|x|\leq 2r\}$ is $\mathcal{O}(r^{\mathsf{d}})\times\mathcal{O}(r^{\mathsf{d}})$ and the rest of the sum can be compared to an integral (here, without loss of generality, we assume $r$ large enough so that $|x|^{\mathsf{d}}\leq 10|x-r|^{\mathsf{d}}$ for $|x|\geq 2r$):
$\sum_{x\in\mathcal{L},|x|\geq 2r}|x|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(x)\}\right]\preceq\int_{t\in\mathbb{R}^{\mathsf{d}},|t|\geq r}|t|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(t)\}\right]\mathtt{d}t,$
which itself, using Fubini’s theorem, can be compared to an expression involving the law $P_{0}$ of $\mathrm{p}_{0}$, namely:
$\int_{t\in\mathbb{R}^{\mathsf{d}},|t|\geq r}|t|^{\mathsf{d}}\mathbb{P}\left[\{\mathrm{p}_{0}\in\mathrm{B}_{r+C}(t)\}\right]\preceq r^{\mathsf{d}}\int_{t\in\mathbb{R}^{\mathsf{d}}}|t|^{\mathsf{d}}\mathtt{d}P_{0}(t),$
and the last integral coincides with $\mathbb{E}[|\mathrm{p}_{0}|^{\mathsf{d}}]$, which is finite. ∎
In particular the reduced second factorial moment measure $\alpha_{\mathrm{red}}^{(2)}$ of $\mathbf{PL}$ is well-defined. We now give its expression.
The reduced covariance function $\alpha_{\mathrm{red}}^{(2)}$ of perturbed lattices
Proposition 2.8.
For all bounded Borel subsets $B$ of $\mathbb{R}^{\mathsf{d}}$, we have:
$\alpha_{\mathrm{red}}^{(2)}(B)=\mathbb{E}\sum_{x\in\mathcal{L}\setminus\{0\}}\mathbf{1}{\{x+\mathrm{p}_{x}-\mathrm{p}_{0}\in B\}}.$ (2.5)
Proof of Proposition 2.8.
We start from the definition (2.4) of $\alpha_{\mathrm{red}}^{(2)}$, using the fundamental domain $\mathcal{D}$ of $\mathcal{L}$ instead of $[0,1)^{\mathsf{d}}$ (this is valid because $\mathcal{D}$ has volume $1$):
$\alpha_{\mathrm{red}}^{(2)}(B)=\mathbb{E}\sum^{\neq}_{x\in\mathbf{PL},y\in\mathbf{PL}}\mathbf{1}\left\{x\in\mathcal{D},y-x\in B\right\}\\ =\mathbb{E}\sum^{\neq}_{x\in\mathcal{L},y\in\mathcal{L}}\mathbf{1}\left\{x+\mathrm{p}_{x}+u\in\mathcal{D},y+\mathrm{p}_{y}-x-\mathrm{p}_{x}\in B\right\}\\ =\mathbb{E}\sum^{\neq}_{x\in\mathcal{L},y\in\mathcal{L}}\mathbf{1}\left\{\mathrm{p}_{x}+u\in\mathcal{D}-x,y+\mathrm{p}_{y}-x-\mathrm{p}_{x}\in B\right\}\\ =\mathbb{E}\sum_{x\in\mathcal{L}}\sum_{y\in\mathcal{L}\setminus\{x\}}\mathbf{1}\left\{\mathrm{p}_{0}+u\in\mathcal{D}-x,y+\mathrm{p}_{y-x}-x-\mathrm{p}_{0}\in B\right\}\\ =\mathbb{E}\sum_{x\in\mathcal{L}}\sum_{y^{\prime}\in\mathcal{L}\setminus\{0\}}\mathbf{1}\left\{\mathrm{p}_{0}+u\in\mathcal{D}-x,y^{\prime}+\mathrm{p}_{y^{\prime}}-\mathrm{p}_{0}\in B\right\}\\ =\mathbb{E}\sum_{y^{\prime}\in\mathcal{L}\setminus\{0\}}\mathbf{1}\left\{\mathrm{p}_{0}+u\in\cup_{x\in\mathcal{L}}\{\mathcal{D}-x\},y^{\prime}+\mathrm{p}_{y^{\prime}}-\mathrm{p}_{0}\in B\right\}\\ =\mathbb{E}\sum_{y^{\prime}\in\mathcal{L}\setminus\{0\}}\mathbf{1}{\{y^{\prime}+\mathrm{p}_{y^{\prime}}-\mathrm{p}_{0}\in B\}}.$
In the first line we use the definition of $\mathbf{PL}$ (the symbol $\neq$ there refers to points in the perturbed lattice that are obtained as perturbations of two different lattice points (they could exceptionally end up at the same spot after perturbations), the second line is elementary, the third line uses the $\mathcal{L}$-invariance of the perturbations, the fourth line uses the fact that $\mathcal{L}$ is a group, the fifth line uses the fact that the translates $\{\mathcal{D}-x\}$ of the fundamental domain are disjoint and the sixth line makes use of the fact that they cover $\mathbb{R}^{\mathsf{d}}$ (see Section 2.2). ∎
Remainder estimates for $\alpha_{\mathrm{red}}^{(2)}$
Lemma 2.9.
Let $f:\mathbb{R}^{\mathsf{d}}\to\mathbb{R}$ be continuous and compactly supported in $\mathrm{B}_{m}$ for some $m\geq 1$. Assume that $M\geq 10m$. We have:
$\mathbb{E}\sum_{x\in\mathcal{L},|x|\geq M}f(x+\mathrm{p}_{x}-\mathrm{p}_{0})\preceq\|f\|_{\infty}\times\mathbb{E}\left[|\mathrm{p}_{0}|^{\mathsf{d}}\times\mathbf{1}\left\{|\mathrm{p}_{0}|\geq\frac{1}{4}M\right\}\right].$ (2.6)
Note that if $\mathbb{E}[|\mathrm{p}_{0}|^{\mathsf{d}}]<+\infty$, then the right-hand side of (2.6) tends to 00 as $M\to\infty$.
Proof of Lemma 2.9.
Without loss of generality, we can assume that $0\leq f\leq 1$. Let $|x|\geq M$. Since $f$ is supported on $\mathrm{B}_{m}$, then for $f(x+\mathrm{p}_{x}-\mathrm{p}_{0})$ to not be 00 we need $|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq|x|-m\geq\frac{1}{2}|x|$, so $\max(|\mathrm{p}_{x}|,|\mathrm{p}_{0}|)\geq\frac{1}{4}{|x|}$. We thus have, since $\mathrm{p}_{x}$ and $\mathrm{p}_{0}$ are identically distributed:
$\sum_{|x|\geq M}\mathbb{E}f(x+\mathrm{p}_{x}-\mathrm{p}_{0})\preceq\sum_{|x|\geq M}\mathbb{P}\left(|\mathrm{p}_{0}|\geq\frac{1}{4}|x|\right),$
and by elementary lattice counting this discrete sum is comparable to the integral:
$\int_{s\geq\frac{M}{4}}s^{\mathsf{d}-1}\mathbb{P}\left(|\mathrm{p}_{0}|\geq s\right)\preceq\mathbb{E}\left[|\mathrm{p}_{0}|^{\mathsf{d}}\mathbf{1}\left\{|\mathrm{p}_{0}|\geq\frac{M}{4}\right\}\right].$
2.5 Expressing the number variance in balls
An auxiliary function
For $r>0$, a classical computation shows that the Fourier transform of the indicator function of a ball of radius $r$ is given by: $\widehat{\mathsf{1}_{\mathrm{B}_{r}}}=\frac{\mathrm{J}_{\mathsf{d}/2}(r|t|)}{|t|^{\frac{\mathsf{d}}{2}}}r^{\frac{\mathsf{d}}{2}}$, where $\mathrm{J}_{\mathsf{d}/2}$ is a Bessel function of the first kind (whose definition is not needed here). As a consequence, we have $\frac{\widehat{\mathsf{1}_{\mathrm{B}_{r}}\ast\mathsf{1}_{\mathrm{B}_{r}}}}{|\mathrm{B}_{r}|}=C_{\mathsf{d}}\frac{\mathrm{J}_{\mathsf{d}/2}(r|t|)^{2}}{|t|^{\mathsf{d}}}$ for some constant $C_{\mathsf{d}}$. We define the function $j_{r}:\mathbb{R}^{\mathsf{d}}\to\mathbb{R}$ by: $j_{r}:t\mapsto C^{\prime}_{\mathsf{d}}\frac{\mathrm{J}_{\mathsf{d}/2}(r|t|)^{2}}{|t|^{\mathsf{d}}}$ with the right constant $C^{\prime}_{\mathsf{d}}$ so that its Fourier transform satisfies:
$\widehat{j_{r}}=\frac{\\ 1_{\mathrm{B}_{r}}\ast\mathsf{1}_{\mathrm{B}_{r}}}{|\mathrm{B}_{r}|}.$ (2.7)
For all $r>0$, we have $\widehat{j_{r}}(0)=\int_{\mathbb{R}^{\mathsf{d}}}j_{r}(t)\mathtt{d}t=1$ and $j_{r}(0)=\frac{|\mathrm{B}_{r}|}{2\pi}$. The family of functions $\{j_{r}\}_{r\geq 1}$ is an approximation of unity: the mass stays constant equal to $1$ and the functions converge pointwise to 00 on $\mathbb{R}^{\mathsf{d}}\setminus\{0\}$, while the value at 00 tends to +$\infty$+ ∞. Since $\widehat{j_{r}}$ is compactly supported (but not smooth), $j_{r}$ is analytic. We now record classical asymptotics for the Bessel function (see e.g. [AS68, Sec. 9.2]):
Asymptotics of $j_{r}$ and $\mathrm{J}_{1}$
For $\mathsf{d}=2$, the function $j_{r}:\mathbb{R}^{2}\to\mathbb{R}$ is given, for some constant $C^{\prime}_{2}$ whose value is irrelevant for us, by:
$t\mapsto j_{r}(t)=C^{\prime}_{2}\frac{\mathrm{J}_{1}(r|t|)^{2}}{|t|^{2}}.$
Lemma 2.10.
For $\mathsf{d}=2$, there exists $C>0$ such that:
  • (Large-$t$ asymptotics): for $|t|\geq r^{-1}$, $\mathrm{J}_{1}^{2}(rt)\leq C\frac{1}{r|t|}$ and thus:
    $|j_{r}(t)|\leq C\frac{1}{r|t|^{3}}\quad(\text{for }|t|\geq r^{-1}).$ (2.8)
    Moreover, for $|t|\leq r^{-1}$, $\mathrm{J}_{1}^{2}(rt)\leq Cr^{2}|t|^{2}$, so $\mathrm{J}_{1}(rt)$ is bounded uniformly in $r,t$.
  • (Precise large-$t$ asymptotics for $\mathrm{J}_{1}$): for $|t|\geq r^{-1}$, there exists constants $a,b$ such that:
    $\mathrm{J}_{1}^{2}(rt)=\frac{a}{r|t|}\cos^{2}\left({r|t|}-\frac{3\pi}{4}\right)+\frac{b}{r^{2}|t|^{2}}\cos\left(r|t|-\frac{3\pi}{4}\right)\sin\left(r|t|-\frac{3\pi}{4}\right)+\mathcal{O}\left(\frac{1}{r^{3}|t|^{3}}\right),$
    which, after some elementary trigonometry and up to changing the constant $b$, can be re-written as:
    $\mathrm{J}_{1}^{2}(rt)=\frac{a}{r|t|}\sin^{2}\left({r|t|}-\frac{\pi}{4}\right)+\frac{b}{r^{2}|t|^{2}}\cos\left(2r|t|\right)+\mathcal{O}\left(\frac{1}{r^{3}|t|^{3}}\right).$ (2.9)
Expressing the number variance
When $\mathbf{X}$ is a random point configuration, the following is well-known (see e.g. [Cos21, Sec. 1.2]):
Lemma 2.11.
Assume $\mathbf{X}$ is stationary, with intensity $1$ and locally square integrable. For all $r>0$:
$\frac{\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=\left\langle\delta_{0}+(\alpha_{\mathrm{red}}^{(2)}-1),\frac{\mathbf{1}_{\mathrm{B}_{r}}\ast\mathbf{1}_{\mathrm{B}_{r}}}{|\mathrm{B}_{r}|}\right\rangle=\left\langle\alpha_{\mathrm{red}}^{(2)},\frac{\mathbf{1}_{\mathrm{B}_{r}}\ast\mathbf{1}_{\mathrm{B}_{r}}}{|\mathrm{B}_{r}|}\right\rangle+1-|\mathrm{B}_{r}|.$ (2.10)
In view of (2.7) one can re-write (2.10) as:
$\frac{\mathrm{Var}\left[|\mathbf{X}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=\left\langle\alpha_{\mathrm{red}}^{(2)},\widehat{j_{r}}\right\rangle+1-|\mathrm{B}_{r}|,$ (2.11)
and the hyperuniformity of $\mathbf{X}$ is thus equivalent to having $\left\langle\alpha_{\mathrm{red}}^{(2)},\widehat{j_{r}}\right\rangle=|\mathrm{B}_{r}|-1+o(1)$, while class I hyperuniformity is equivalent to $\left\langle\alpha_{\mathrm{red}}^{(2)},\widehat{j_{r}}\right\rangle=|\mathrm{B}_{r}|-1+\mathcal{O}(r^{-1})$ as $r\to\infty$.
3 Positive result in dimension 2
This section is devoted to the proof of our main result in dimension $2$, namely:
Proposition 3.1.
If the perturbations have a finite second moment, then the perturbed lattice is hyperuniform.
3.1 Preliminaries
Recall Remark 2.5: without loss of generality, we assume that the perturbations are centered. In this section, we assume moreover that the perturbations have a finite second moment. In particular, by Proposition 2.8, the measure $\alpha_{\mathrm{red}}^{(2)}$ of $\mathbf{PL}$ exists and for all $r>0$, we have (see (2.11)):
$\frac{\mathrm{Var}\left[|\mathbf{PL}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=\left\langle\alpha_{\mathrm{red}}^{(2)},\widehat{j_{r}}\right\rangle+1-|\mathrm{B}_{r}|=\mathbb{E}\sum_{x\in\mathcal{L}\setminus\{0\}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})+1-|\mathrm{B}_{r}|=A+B,$ (3.1)
where $A$ is a deterministic contribution coming from the lattice and $B$ is the “error term” due to the perturbations:
$A:=\sum_{x\in\mathcal{L}\setminus\{0\}}\widehat{j_{r}}(x)+1-|\mathrm{B}_{r}|,\quad B:=\mathbb{E}\sum_{x\in\mathcal{L}\setminus\{0\}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x).$
The deterministic term A.
The fact that $A$ tends to zero as $r\to\infty$ is equivalent to the hyperuniformity of the (unperturbed) stationary lattice $\mathbf{L}$, and follows from a simple computation which we now recall.
Claim 3.2.
We have $A=\mathcal{O}(r^{-1})$, which corresponds to the fact that a stationary lattice is not only hyperuniform, but is even class I hyperuniform in the sense of (1.2).
Proof of Claim 3.2.
Write $A$ as $A=\sum_{x\in\mathcal{L}}\widehat{j_{r}}(x)-\widehat{j_{r}}(0)+1-|\mathrm{B}_{r}|$. On the one hand, applying Poisson’s formula as in (2.1) (recall that $\mathcal{L}^{\star}$ denotes the dual lattice), we have:
$\sum_{x\in\mathcal{L}}\widehat{j_{r}}(x)={2\pi}\sum_{k\in\mathcal{L}^{\star}}j_{r}(2\pi k)=2\pi j_{r}(0)+2\pi\sum_{k\in\mathcal{L}^{\star},k\neq 0}j_{r}(2\pi k).$
On the other hand we have $\widehat{j_{r}}(0)=1$ and ${2\pi}j_{r}(0)=|\mathrm{B}_{r}|$ (see Section 2.5). We are thus left with $A=2\pi\sum_{k\in\mathcal{L}^{\star},k\neq 0}j_{r}(2\pi k)$ and using the decay of $j_{r}$ stated in (2.8) we can easily bound that sum:
$\sum_{k\in\mathcal{L}^{\star},k\neq 0}j_{r}(2\pi k)\preceq\sum_{k\in\mathcal{L}^{\star},k\neq 0}\frac{1}{r|k|^{3}}=\mathcal{O}(r^{-1}).$
The rest of Section 3 is devoted to showing that the term $B$ in (3.1) tends to 00 as $r\to\infty$.
Taylor expansion of the characteristic function.
In order to control $B$, we first study the characteristic function $\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}$ of the random variable $\mathrm{p}_{x}-\mathrm{p}_{0}$ for $x\in\mathcal{L}$. Mainly, we will rely on a second order expansion of $\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)$ for $t$ small, which makes a crucial use of the existence of a second moment for the perturbations.
Lemma 3.3.
There exists $g:\mathbb{R}^{2}\to(0,+\infty)$ depending only on the distribution of $\mathrm{p}_{0}$ such that:
  1. 1.
    $\lim_{t\to 0}g(t)=0$, $g$ is bounded on $\mathbb{R}^{2}$. Moreover $g$ is radially symmetric and radially non-decreasing.
  2. 2.
    For all $x\in\mathcal{L}$, we control the error term in the second-order Taylor’s expansion of $\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}$ by:
    $\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\right)\right|\leq g(t)\min\left(|t|^{2},1\right),\text{ for all }t\in\mathbb{R}^{2}.$ (3.2)
Using such a “small wave-number expansion” is a key part in the argument of [Gab04].
Proof of Lemma 3.3.
An elementary estimate found in [Dur19, (3.3.3)] says that if $X$ is a random variable in $\mathbb{R}^{\mathsf{d}}$, for all $n\geq 0$ we have the following control on the error term in the Taylor’s expansion of the characteristic function of $X$ to order $n$:
$\left|\mathbb{E}[e^{-is\cdot X}]-\sum_{k=0}^{n}\frac{(-is)^{k}}{k!}\mathbb{E}[X^{k}]\right|\leq\mathbb{E}\left[\min(|sX|^{n+1},2|sX|^{n})\right]\text{ for all $s\in\mathbb{R}^{\mathsf{d}}$.}$ (3.3)
For $x\in\mathcal{L}$, we can use (3.3) with $X=\mathrm{p}_{x}-\mathrm{p}_{0}$ and $n=2$, which gives:
$\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\right)\right|\preceq|t|^{2}\mathbb{E}\left[\min\left(|t||\mathrm{p}_{x}-\mathrm{p}_{0}|^{3},|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right)\right].$
Using elementary inequalities and the fact that $\mathrm{p}_{x}$ and $\mathrm{p}_{0}$ are identically distributed, we can write
$|t|^{2}\mathbb{E}\left[\min(|t||\mathrm{p}_{x}-\mathrm{p}_{0}|^{3},|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2})\right]\preceq|t|^{2}\mathbb{E}\left[\min(|t||\mathrm{p}_{0}|^{3},|\mathrm{p}_{0}|^{2})\right].$ (3.4)
The right-hand side of (3.4) is radially symmetric, radially non-decreasing, does not depend on $x$ and is $o(|t|^{2})$ for $t\to 0$ by dominated convergence (the fact that the convergence is indeed dominated relies on the fact that $\mathrm{p}_{0}$ has a finite second moment). This shows that we can indeed take $g(t)\to 0$ as $t\to 0$ in (3.2). On the other hand, for large values of $t$, since $\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}$ is bounded by $1$, the error in (3.2) is exactly of order $|t|^{2}$ and we can take $g(t)$ of order $1$.
In general, we cannot control the “speed” for the decay of $g$. Note however that if we placed a stronger moment assumption on $\mathrm{p}_{0}$, e.g. by asking that $\mathbb{E}[|\mathrm{p}_{0}|^{2+\delta}]$ be finite, then we could take an explicit form for $g(t)$ as $|t|^{\delta}$ for $|t|$ small. ∎
In the sequel, we may thus use the following expansion for $|t|$ small:
$\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)=1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]+o(|t|^{2}),$ (3.5)
with an error term which may depend on $x$ but is bounded by $g(t)|t|^{2}$ uniformly in $x$.
Truncation near the origin.
We choose a pair of smooth functions $(\chi_{\ell},\widehat{\chi_{\ell}})$ which will serve as truncations in physical/Fourier space.
First, let $\chi$ be a smooth, compactly supported, radially symmetric function on $\mathbb{R}^{2}$ with values in $[0,1]$ such that $\chi(t)=1$ if $|t|\leq\frac{1}{2}$, $\chi(t)=0$ if $|t|\geq 1$. Both $\chi$ and $\widehat{\chi}$ are in Schwartz space and since $\chi$ is real-valued, $\widehat{\chi}$ is even.
Next, let $\ell$ be a length scale depending on $r$, to be chosen later, such that:
$\frac{100}{r}\leq\ell\leq\frac{1}{100}\text{ for all $r\gg 1$, and }{\lim_{r\to\infty}\ell=0,}\lim_{r\to\infty}r\ell=+\infty.$ (3.6)
As will appear below, our analysis requires a proper choice of the lengthscale, with $\ell r\to\infty$ in a possibly very slow fashion, depending on the common law of the perturbations. We let $\chi_{\ell}:=\chi\left(\cdot/\ell\right)$, so that $\widehat{\chi_{\ell}}=\ell^{2}\widehat{\chi}(\ell\cdot)$. We have:
  • $\chi_{\ell}$ is smooth, radially symmetric, non-negative, bounded on $\mathbb{R}^{2}$, equal to $1$ on $\mathrm{B}_{\ell/2}$ and compactly supported in $\mathrm{B}_{\ell}$.
  • $\widehat{\chi_{\ell}}$ is smooth, bounded by $\mathcal{O}(\ell^{2})$, with $\int_{\mathbb{R}^{2}}\widehat{\chi_{\ell}}=2\pi\chi_{\ell}(0)=2\pi$. Moreover, since $\widehat{\chi}$ is a Schwartz function, so is $\widehat{\chi_{\ell}}$ more precisely, for all $m\geq 0$, there exists $C_{m}$ (independent of $\ell$) such that:
    $|\widehat{\chi_{\ell}}(t)|\leq\frac{C_{m}\ell^{2}}{(1+\ell|t|)^{m}},\quad|\mathrm{Hess}\widehat{\chi_{\ell}}(t)|\leq\frac{C_{m}\ell^{4}}{(1+\ell|t|)^{m}}\text{ for all $t\in\mathbb{R}^{2}$.}$ (3.7)
Decomposing the term $B$.
Truncation.
Let us recall the expression of $B$, as coming from (3.1):
$B:=\mathbb{E}\sum_{x\in\mathcal{L}\setminus\{0\}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x).$
The function $\widehat{j_{r}}$ is bounded by $1$ and supported on $\mathrm{B}_{2r}$, thus by Lemma 2.9 we can restrict the sum to $\{|x|\leq 100r\}$ with a small error. We obtain:
$B=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\left(\mathbb{E}\left[\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})\right]-\widehat{j_{r}}(x)\right)+o_{r}(1).$ (3.8)
Note that the error term in (3.8) is in fact bounded, according to (2.6), by $\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq r}\right]$. If $|\mathrm{p}_{0}|$ has e.g. a finite moment of order $2+\varepsilon$ with $\varepsilon>0$, this is $\mathcal{O}(r^{-\varepsilon})$.
Decomposition.
For each $x\in\mathcal{L}$, using the characteristic function $\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}$ of $\mathrm{p}_{x}-\mathrm{p}_{0}$, we can write:
$\mathbb{E}\left[\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})\right]-\widehat{j_{r}}(x)=\mathbb{E}\left[\int_{t\in\mathbb{R}^{2}}e^{-it\cdot(x+\mathrm{p}_{x}-\mathrm{p}_{0})}j_{r}(t)\mathtt{d}t\right]-\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\mathtt{d}t\\ =\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\mathbb{E}\left[e^{-it\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})}\right]-1\right)\mathtt{d}t=\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)\mathtt{d}t.$ (3.9)
Since the mass of $j_{r}$ concentrates around 00, it is natural to introduce the cut-off $\chi_{\ell}$ (as chosen above) and the small-$t$ expansion (3.5) of the integrand. We obtain:
$\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)\chi_{\ell}(t)\mathtt{d}t=-\frac{1}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t\\ +\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)o(|t|^{2})\chi_{\ell}(t)\mathtt{d}t.$
In view of (3.8), we may thus write $B=B_{1}+B_{2}+B_{3}+o_{r}(1)$, where $B_{1}$ is the “main term”:
$B_{1}:=-\frac{1}{2}\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t,$ (3.10)
$B_{2}$ is the error due to the small wave-number Taylor’s expansion:
$B_{2}:=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)o(|t|^{2})\chi_{\ell}(t)\mathtt{d}t,$ (3.11)
and $B_{3}$ represents the contribution of higher wave-numbers:
$B_{3}:=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))\mathtt{d}t.$ (3.12)
We call $B_{1}$ the “main” term because it carries the most tractable information about the perturbations. However, our goal is to show that all three terms $B_{1},B_{2},B_{3}$ are going to 00 as $r\to\infty$.
3.2 The main term $B_{1}$
Recall that $B_{1}$ is given by (note that the restriction $x\neq 0$ is irrelevant):
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L},|x|\leq 100r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t.$ (3.13)
Lemma 3.4.
We have $B_{1}=o_{r}(1)$.
Proof of Lemma 3.4.
$B_{1}$ is written as a finite sum, but we would like to apply Poisson’s formula, so we start by re-introducing the missing terms.
Re-introducing some terms in the sum.
To do so, observe first that $\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t$ is proportional to $\widehat{j_{r}}\ast\Delta\widehat{\chi_{\ell}}(x)$, with $\widehat{j_{r}}$ supported on $\mathrm{B}_{2r}$ and bounded by $1$, while the function $\Delta\widehat{\chi_{\ell}}$ satisfies (see (3.7) with $m=5$):
$|\Delta\widehat{\chi_{\ell}}(t)|\preceq\frac{\ell^{4}}{\ell^{5}|t|^{5}}=\frac{1}{\ell|t|^{5}}\text{ for all $|t|\geq\ell^{-1}$.}$ (3.14)
If $|x|\geq 100r$ we have $|x-\cdot|\geq\frac{1}{2}|x|$ on $\mathrm{B}_{2r}$ and we can thus write, using (3.14) and $\|\widehat{j_{r}}\|_{\infty}\leq 1$:
$\widehat{j_{r}}\ast\Delta\widehat{\chi_{\ell}}(x)=\int_{\mathrm{B}_{2r}}\widehat{j_{r}}(u)\Delta\widehat{\chi_{\ell}}(x-u)\mathtt{d}u\preceq|\mathrm{B}_{2r}|\times 1\times\frac{1}{\ell|x|^{5}}\preceq r^{2}\ell^{-1}|x|^{-5}.$
Summing over $x\in\mathcal{L}$ with $|x|\geq 100r$ yields (we can compare $\sum_{x\in\mathcal{L},|x|\geq 100r}|x|^{-5}$ to an integral):
$\sum_{x\in\mathcal{L},|x|\geq 100r}\widehat{j_{r}}\ast\Delta\widehat{\chi_{\ell}}(x)\preceq\sum_{x\in\mathcal{L},|x|\geq 100r}r^{2}\ell^{-1}|x|^{-5}=\mathcal{O}((\ell r)^{-1})=o_{r}(1).$
We may thus re-introduce the “missing” terms in the sum defining $B_{1}$ and write:
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L}}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t+o_{r}(1),$ (3.15)
with an error term bounded by $\mathcal{O}((\ell r)^{-1})$.
The spectral measure.
Let $g_{\ell}$ be the following map, appearing in the integrand in (3.13):
$g_{\ell}:t\mapsto\chi_{\ell}(t)|t|^{2}.$
Introduce the autocorrelation function of the perturbation field $(\mathrm{p}_{x})_{x\in\mathcal{L}}$, namely the function $G:\mathcal{L}\to\mathbb{R}$ defined by (recall that the perturbations are assumed to be centered):
$G(x):=\mathbb{E}[\mathrm{p}_{x}\cdot\mathrm{p}_{0}].$ (3.16)
Since $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ is translation-invariant, it is classical that:
  • The autocorrelation function is bounded by $|G(x)|\leq G(0)$ for all $x\in\mathcal{L}$.
  • $G$ is a positive-definite function on $\mathcal{L}$. In particular, by Bochner’s theorem (see [Fol16, Thm. (4.18)]), there exists a “spectral measure” $\rho$ which is a non-negative finite measure on the Pontryagin dual $\Omega$ of the locally compact abelian group $\mathcal{L}$, such that:
    $G(x)=\int_{\Omega}e^{-2i\pi x\cdot\omega}\mathtt{d}\rho(\omega)\text{ for all }x\in\mathcal{L}.$ (3.17)
    Note that the Pontryagin dual of $\mathcal{L}$ is different from, although related to, the dual lattice $\mathcal{L}^{\star}$. When $\mathcal{L}=\mathbb{Z}^{2}$, its Pontryagin dual is the two-dimensional torus $\mathbb{T}^{2}$ which we can identify with $[-\frac{1}{2},\frac{1}{2})^{2}$. In general, we identify $\Omega$ with a copy of the fundamental domain $\mathcal{D}^{\star}$ of $\mathcal{L}^{\star}$ centered at the origin (this is convenient for the organization of our computation below, but in no way necessary).
Since $\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]=2(G(0)-G(x))$ for all $x\in\mathcal{L}$, we may return to (3.15), use our notation, and write $B_{1}$ as $B_{1}=B_{11}-B_{12}+o_{r}(1)$, with:
$B_{11}:=\sum_{x\in\mathcal{L}}G(x)\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t,\quad B_{12}:=G(0)\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t.$
Controlling $B_{11}.$
Using the spectral measure $\rho$ as in (3.17), we can re-write $B_{11}$ as:
$B_{11}=\sum_{x\in\mathcal{L}}\int_{\Omega}e^{-2i\pi x\cdot\omega}\left(\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t\right)\ \mathtt{d}\rho(\omega)=\int_{\Omega}\sum_{x\in\mathcal{L}}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}}(x)\mathtt{d}\rho(\omega).$
Applying Poisson’s formula (2.1) to the sum $\sum_{x\in\mathcal{L}}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}}(x)$ for all fixed $\omega\in\Omega$, we obtain:
$B_{11}=2\pi\sum_{k\in\mathcal{L}^{\star}}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}(2\pi k+2\pi\omega)\mathtt{d}\rho(\omega).$ (3.18)
Note that $g_{\ell}$ is a Schwartz function and thus so is $\widehat{g_{\ell}}$ and (since $\widehat{j_{r}}$ is $L^{1}$), so is $\widehat{j_{r}\times g_{\ell}}$ , which justifies our interversions of sums over $x\in\mathcal{L}$ and integrals against $\mathtt{d}\rho$, as well as the use of Poisson’s formula. We now distinguish between the lattice points $k=0$ and $k\in\mathcal{L}^{\star}\setminus\{0\}$.
For $k\neq 0$.
We chose $\Omega$ as a parallelotope (because it is the fundamental domain of some lattice) containing 00. For all $k\in\mathcal{L}^{\star}\setminus\{0\}$, the function $\omega\mapsto k+\omega$ stays away from 00 on $\Omega$ in the sense that:
$\min_{k\in\mathcal{L}^{\star}\setminus\{0\}}\min_{\omega\in\Omega}|k+\omega|\geq c:=\mathrm{dist}(0,\partial\Omega)>0.$
(In the case $\mathcal{L}=\mathcal{L}^{\star}=\mathbb{Z}^{2}$, we are simply saying that if $k\in\mathbb{Z}^{2}\setminus\{0\}$ and $u\in[\frac{1}{2},\frac{1}{2})^{2}$, then $|k+u|$ is bounded below by a positive constant). Thus, since $\chi_{\ell}$ is supported on $\mathrm{B}_{\ell}$, if $\ell$ is sufficiently small then we have $\chi_{\ell}(2\pi k+2\pi\omega)=0$ for all $k\in\mathcal{L}^{\star}\setminus\{0\}$ and all $\omega\in\Omega$, and in turn:
$\sum_{k\in\mathcal{L}^{\star}\setminus\{0\}}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}(2\pi k+2\pi\omega)\mathtt{d}\rho(\omega)=0.$
For $k=0$.
Denote by $C(r)$ the quantity:
$C(r):=\int_{\Omega}j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)\mathtt{d}\rho(\omega).$ (3.19)
Let us now justify that $C(r)$ tends to 00 as $r\to\infty$. The family of functions $\omega\mapsto j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)$ converges pointwise to 00 on $\Omega$ as $r\to\infty$ because:
  • $j_{r}$ converges pointwise to 00 on $\mathbb{R}^{2}\setminus\{0\}$ (see Section 2.5)
  • $g_{\ell}$ is bounded and $g_{\ell}(0)=0$.
Moreover, we have $j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)\preceq\frac{\mathrm{J}_{1}^{2}(2\pi r\omega)}{|\omega|^{2}}\times|\omega|^{2}=\mathrm{J}_{1}^{2}(2\pi r\omega)$, which is bounded uniformly in $r,\omega$ (see Lemma 2.10). Thus by the dominated convergence theorem, we have indeed:
$\lim_{r\to 0}\int_{\Omega}j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)\mathtt{d}\rho(\omega)=0.$ (3.20)
This is a point of our analysis where the speed of decay is more difficult to quantify. For example, imposing additional moment conditions on the perturbations does not help. We can however make the following observation:
Claim 3.5.
We have $\lim_{r\to\infty}C(r)=0$ in such a way that:
$\sum_{n\geq 0}C(2^{n})<+\infty.$
Proof of Claim 3.5.
Distinguishing between $|\omega|\leq\frac{1}{2\pi r}$ and $\frac{1}{2\pi r}\leq|\omega|\leq 1$, and using the asymptotics of Lemma 2.10, we can write:
$\int_{|\omega|\leq\frac{1}{2\pi r}}j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)\mathtt{d}\rho(\omega)\preceq\int_{|\omega|\leq\frac{1}{2\pi r}}r^{2}\times|\omega|^{2}\mathtt{d}\rho(\omega),$ (3.21)
as well as:
$\int_{\frac{1}{2\pi r}\leq|\omega|\leq 1}j_{r}(2\pi\omega)g_{\ell}(2\pi\omega)\mathtt{d}\rho(\omega)\preceq\int_{\frac{1}{2\pi r}\leq|\omega|\leq 1}\frac{1}{r|\omega|}\mathtt{d}\rho(\omega).$ (3.22)
For $j\geq 0$, let $\mathtt{A}(j)$ be the annulus $\{2^{-j-1}\leq|\omega|<2^{-j}\}$. In view of (3.21), we can compare $C(r)$ to the following sum:
$C(r)\preceq\frac{1}{r}\sum_{j=0}^{\log_{2}r-1}2^{j}\rho(\mathtt{A}(j))+r^{2}\sum_{j=\log_{2}r}^{+\infty}2^{-2j}\rho(\mathtt{A}(j)).$
In particular, if we take $r=2^{n}$ and sum over $n$, we obtain, since $\rho$ is a finite measure:
$\sum_{n\geq 0}C(2^{n})\preceq\sum_{n\geq 0}2^{-n}\sum_{j=0}^{n-1}2^{j}\rho(\mathtt{A}(j))+\sum_{n\geq 0}2^{2n}\sum_{j=n}^{+\infty}2^{-2j}\rho(\mathtt{A}(j))\\ =\sum_{j=0}^{+\infty}2^{j}\rho(\mathtt{A}(j))\sum_{n\geq j+1}2^{-n}+\sum_{j=0}^{+\infty}2^{-2j}\rho(\mathtt{A}(j))\sum_{n=0}^{j}2^{2n}\preceq\sum_{j=0}^{+\infty}\rho(\mathtt{A}(j))<+\infty.$
Handling $B_{12}.$
We are left with the term $B_{12}$. Using Poisson’s formula, we get for $r$ large enough:
$B_{12}:=G(0)\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t=G(0)\sum_{x\in\mathcal{L}}\widehat{j_{r}\times g_{l}}(x)=2\pi G(0)\sum_{k\in\mathcal{L}^{\star}}j_{r}(2\pi k)g_{l}(2\pi k)=0.$
Indeed $j_{r}(0)g_{l}(0)=0$, and since $g_{\ell}$ is supported on $\mathrm{B}_{\ell}$ the other summands vanish for $\ell$ small enough (depending on $\mathcal{L}$) thus for $r$ large enough. ∎
Conclusion.
We have obtained that $\lim_{r\to 0}B_{1}=0$. More precisely, for $r$ large enough, $B_{1}$ reduces to the sum of two error terms: the one from (3.15), which is $\mathcal{O}((r\ell)^{-1})$, and the one corresponding to $C(r)$ as in (3.19), which is such that $\sum_{n\geq 0}C(2^{n})<+\infty$ by Claim 3.5.
3.3 The error term $B_{2}$
Recall that:
$B_{2}:=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)o(|t|^{2})\chi_{\ell}(t)\mathtt{d}t.$
Lemma 3.6.
We have $B_{2}=\mathcal{O}\left(r\ell g(\ell)\right)$, with $g$ as in Lemma 3.3.
Proof of Lemma 3.6.
We use a rough triangular inequality and control $B_{2}$ by:
$B_{2}=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)o(|t|^{2})\mathtt{d}t\preceq r^{2}\int_{t\in\mathbb{R}^{2}}j_{r}(t)\chi_{\ell}(t)|t|^{2}g(t)\mathtt{d}t,$
where $g$ is as in Lemma 3.3. Recall that $\chi_{\ell}$ is supported on $\mathrm{B}_{\ell}$, and let us decompose the integral into two parts: first, since $\|j_{r}\|_{L^{1}}$ and $\|\chi_{\ell}\|_{\infty}$ are bounded, we have:
$\int_{t\in\mathbb{R}^{2},|t|\leq r^{-1}}j_{r}(t)\chi_{\ell}(t)|t|^{2}g(t)\mathtt{d}t\leq\|j_{r}\|_{L^{1}}\|\chi_{\ell}\|_{\infty}\times\sup_{|t|\leq r^{-1}}|t|^{2}g(t)=\mathcal{O}\left(r^{-2}g(r^{-1})\right),$
while on the other hand, using the decay (2.8) of $j_{r}$ (and the boundedness of $\chi_{\ell}$):
$\int_{t\in\mathbb{R}^{2},r^{-1}\leq|t|\leq\ell}j_{r}(t)\chi_{\ell}(t)|t|^{2}g(t)\mathtt{d}t\preceq\int_{t\in\mathbb{R}^{2},r^{-1}\leq|t|\leq\ell}\frac{1}{r|t|^{3}}|t|^{2}g(t)\mathtt{d}t\preceq\ell r^{-1}g(\ell).$
(In both inequalities, we have used the fact that $g$ is radially non-decreasing). Keeping only the dominant term and multiplying by $r^{2}$ proves the lemma. ∎
Recall that we are requiring $\ell$ to be larger than $\frac{1}{r}$, with $\ell r\to\infty$ (see (3.6)), and we now also want $\lim_{r\to}r\ell g(\ell)=0$. The choice of such an $\ell$ is not completely obvious in general, especially because we do not have an explicit form for $g$. If we knew (cf. Lemma 3.3) that $g(t)=\mathcal{O}(t^{\delta})$ as $t\to 0$, then we could take here $\ell$ as $r^{-1+\delta^{\prime}}$ for some $\delta^{\prime}>0$.
3.4 The error term $B_{3}$
Recall that:
$B_{3}:=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}{e^{-it\cdot x}}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))\mathtt{d}t.$ (3.23)
We will denote below by $\mathcal{B}_{3}(x)$ the summand corresponding to $x\in\mathcal{L}$.
Heuristically speaking, the function $j_{r}$ is an approximation of unity which lives “at scale $r^{-1}$” around the origin, and by construction $(1-\chi_{\ell})$ vanishes at scales smaller than $\ell$ (and we took $\ell\gg r^{-1}$), moreover $\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}-1\right)$ is bounded, so the integrand $j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))$ is “small” for all $t$’s and $x$’s and we can expect $B_{3}$ to be small. This seems to be implicitly taken for granted in physics papers, which discard such terms altogether in their analysis of the structure factor.
However, the function $t\mapsto j_{r}(t)$ only decays as $\frac{1}{r|t|^{3}}$ when $|t|\geq r^{-1}$ (see (2.8)), which is not very small (in terms of $r$), and the corresponding rough bound for each summand in $B_{3}$ is (using Lemma 3.3):
$\mathcal{B}_{3}(x):=\int_{t\in\mathbb{R}^{2}}{e^{-it\cdot x}}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))\mathtt{d}t\preceq\int_{t\in\mathbb{R}^{2},|t|\geq\ell}\frac{1}{r|t|^{3}}\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)\times\mathcal{O}(1)\mathtt{d}t\\ \preceq\int_{t\in\mathbb{R}^{2},\ell\leq|t|\leq 1}\frac{1}{r|t|^{3}}\times\mathcal{O}(|t|^{2})\mathtt{d}t+\int_{|t|\geq 1}\frac{1}{r|t|^{3}}\times\mathcal{O}(1)\mathtt{d}t=\mathcal{O}(r^{-1})$ (3.24)
and thus, by summing (3.24) over $\mathcal{O}(r^{2})$ terms, one would get $\mathcal{O}(r)$, which is far from $o_{r}(1)$. We thus need to abandon rough bounds on the integrand (at least for the vast majority of the summands) and work with the cancellations coming from the oscillating term.
Proposition 3.7.
There exists $\delta>0$ such that:
$B_{3}={\mathcal{O}\left(\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\ell^{-1}}\right]\right)+\mathcal{O}(r^{-\delta})}.$ (3.25)
Since $\mathrm{p}_{0}$ has a finite second moment, and $\ell\to 0$, the right-hand side of (3.25) is $o_{r}(1)$.
Studying $\mathcal{B}_{3}(x)$ as in (3.24) for $x\in\mathcal{L}$ quickly reveals that the circle $\partial\mathrm{B}_{2r}$ plays a specific role and that estimates tend to degenerate (or to simply get more difficult to obtain) when the lattice point $x$ is close to $\partial\mathrm{B}_{2r}$. A simple way to understand why is to observe that the definition of $\mathcal{B}_{3}(x)$ involves the Fourier transform $\widehat{j_{r}}$ of $j_{r}$ and that $\widehat{j_{r}}$ (which is proportional to the convolution of the indicator function of $\mathrm{B}_{r}$ with itself) fails to be smooth along $\partial\mathrm{B}_{2r}$. We perform three distinct analyses of $\mathcal{B}_{3}(x)$ depending on the position of $x\in\mathcal{L}$:
  • If $x$ is “far from the boundary”, we re-write $\mathcal{B}_{3}(x)$ in “physical” space and give it a simple probabilistic interpretation. We then rely on elementary estimates involving various Taylor’s expansions, the idea being that the functions at play are smooth away from $\partial\mathrm{B}_{2r}$.
  • If $x$ is “close, but not too close to the boundary”, we almost completely ignore the probabilistic interpretation and perform an elementary but somewhat tedious analysis of $\mathcal{B}_{3}(x)$ in Fourier space.
  • Finally, if $x$ is very close to, or even on, the boundary, we resort to the rough bound (3.24) and import a number-theoretic $o(r)$ estimate to count such lattice points.
We postpone the proof of Proposition 3.7 to the appendix, see Section A.
3.5 Conclusion
From (3.1) we found that the rescaled number variance $\sigma(r):=\frac{\mathrm{Var}\left[|\mathbf{PL}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}$ is given by $A+B$, where:
  • $A=o_{r}(1)$ by Claim 3.2.
  • $B=B_{1}+B_{2}+B_{3}+o_{r}(1)$ (see (3.8)).
  • $B_{1}$ is $o_{r}(1)$ by Lemma 3.4.
  • $B_{2}$ is $\mathcal{O}(r\ell g(\ell))$ by Lemma 3.6, with $g$ tending to 00 at 00.
  • $B_{3}$ is $o_{r}(1)$ by Proposition 3.7.
To prove Proposition 3.1, it remains to choose $\ell$ such that $\ell\to 0$, $r\ell\to\infty$ (to satisfy (3.6)) while $r\ell g(\ell)\to 0$. This is simply done by observing that the function $F:x\mapsto x\sqrt{g(x)}$ decreases to 00 as $x\to 0$ and that we may thus consider $\ell:=F^{-1}\left(\frac{1}{r}\right)$, which fits the bill.
Speed of decay.
One could ask about the speed at which $\sigma(r)$ tends to 00. In general, there is no upper bound - arbitrarily slow decays are possible, see Theorem 3, but we can say the following:
  • The term $A$ is in fact $\mathcal{O}(r^{-1})$ (see Claim 3.2), which corresponds to the fastest decay possible here.
  • The error term in the decomposition of $B$ as $B=B_{1}+B_{2}+B_{3}+o_{r}(1)$ is bounded by $\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq r}\right]$ (see (3.8)). If $|\mathrm{p}_{0}|$ has a moment of order $2+\alpha$ with $\alpha>0$, this is $\mathcal{O}(r^{-\alpha})$.
  • $B_{1}$ is (for $r$ large enough), the sum of two error terms: one is $\mathcal{O}((r\ell)^{-1})$, and the other is $C(r)$ as in (3.19), which satisfies $\sum_{n\geq 0}C(2^{n})<+\infty$ by Claim 3.5 (see the conclusion of Section 3.2).
  • $B_{2}$ is $\mathcal{O}(r\ell g(\ell))$ with $g$ coming from the Taylor’s expansion (3.2).
  • $B_{3}$ is ${\mathcal{O}\left(\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\ell^{-1}}\right]\right)+\mathcal{O}(r^{-\delta})}$ for some $\delta>0$ by Proposition 3.7.
If we assume that $|\mathrm{p}_{0}|$ has a moment of order $2+\alpha$ with $\alpha>0$, then:
  • The function $g$ from Lemma 3.3 can be taken as a $\mathcal{O}(|t|^{\alpha})$.
  • In particular, we can take $\ell=r^{-1+\alpha^{\prime}}$ for some $\alpha^{\prime}>0$ such that (3.6) is satisfied and $r\ell g(\ell)=\mathcal{O}(r^{-\alpha^{\prime\prime}})$ for some $\alpha^{\prime\prime}>0$.
  • The error term $\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\ell^{-1}}\right]$ becomes $\mathcal{O}(r^{-\alpha^{\prime\prime\prime}})$ for some $\alpha^{\prime\prime\prime}>0$.
In conclusion, $\sigma(r)=\mathcal{O}(r^{-\beta})+C(r)$ for some $\beta>0$, and in particular (using Claim 3.5) we have:
$\sum_{n\geq 0}\sigma(2^{n})<+\infty.$ (3.26)
Hyperuniform point processes whose rescaled number variance satisfies (3.26) are called $\star$-hyperuniform” in [HL24], where it is argued that they form an interesting sub-class (in fact, the overhelming majority). Moreover, $\star$-hyperuniformity is equivalent to finite (regularized) Coulomb energy (see [HL24, Theorem 4] for further details).
Corollary 3.8.
Let $\mathsf{d}=2$ and assume that $\mathbb{E}[|p_{0}|^{2+\alpha}]<\infty$ for some $\alpha>0$. Then, the resulting perturbed lattice $\mathbf{PL}$ is $\star$-hyperuniform. In particular, it has finite (regularized) Coulomb energy.
4 Constructions
For $N\geq 10$ we let $\Lambda_{N}=\Lambda_{N}^{(0)}$ be the (half-open) cube $\Lambda_{N}:=[-\frac{1}{2}N,\frac{1}{2}N)^{\mathsf{d}}$, and for all $k\in N\mathbb{Z}^{\mathsf{d}}$ we let $\Lambda^{(k)}_{N}:=k+\Lambda^{(0)}_{N}$, which is a cube of sidelength $N$ centered at $k$.
4.1 Infinite number variance without $\mathsf{d}$-th moment
We prove here the third item of Theorem 1.
Proposition 4.1.
Let $\mathsf{d}\geq 1$ and let $X$ be a non-negative random variable with $\mathbb{E}[X^{\mathsf{d}}]=+\infty$. Then there exists a $\mathbb{Z}^{\mathsf{d}}$-invariant perturbation field $(\mathrm{p}_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$ such that:
  1. 1.
    $|\mathrm{p}_{0}|\leq X$ almost surely.
  2. 2.
    The resulting perturbed lattice $\mathbf{PL}$ has infinite number variance in finite balls.
Proof of Proposition 4.1.
Let $(C^{(k)}_{N})_{k\in N\mathbb{Z}^{\mathsf{d}}}$ be independent random variables such that $C^{(k)}_{N}$ is uniformly distributed in $\Lambda^{(k)}_{N}$ for all $k\in N\mathbb{Z}^{\mathsf{d}}$. We first define a “displacement field” (as in Section 2.4) $\tilde{\mathrm{p}}_{N}:\mathbb{R}^{\mathsf{d}}\mapsto\mathbb{R}^{\mathsf{d}}$ by setting, for all $k\in N\mathbb{Z}^{\mathsf{d}}$:
$\tilde{\mathrm{p}}_{N}(x):=C^{(k)}_{N}-x,\text{ for all }x\in\Lambda^{(k)}_{N},$
hence applying $\tilde{\mathrm{p}}_{N}$ corresponds to “sending all the points of the cube $\Lambda^{(k)}_{N}$ to $C^{(k)}_{N}$. This displacement field is clearly $N\mathbb{Z}^{\mathsf{d}}$-stationary, we now turn it into a $\mathbb{R}^{\mathsf{d}}$-stationary field by choosing a random vector $\tau$ uniformly distributed in $\Lambda_{N}$ and independent of the $(C^{(k)}_{N})_{k\in N\mathbb{Z}^{\mathsf{d}}}$, and setting $\bar{\mathrm{p}}_{N}:=\tilde{\mathrm{p}}_{N}(\cdot+\tau)$. Note that by construction, the size of all displacements is bounded by $\sqrt{\mathsf{d}}N$. Finally, we set for a $[0,1]^{d}$-uniform random variable $U$ (independent of everything else):
$\mathbf{PL}^{\prime}_{N}:=\{x+U+\bar{\mathrm{p}}_{N}(x+U),\ x\in\mathbb{Z}^{\mathsf{d}}\}.$
Next, we count the number of points of the corresponding perturbed lattice $\mathbf{PL}^{\prime}_{N}$ falling in the unit ball.
With probability greater than $\frac{|\mathrm{B}_{1}|}{(2N)^{\mathsf{d}}}$ the random variable $C^{(0)}_{N}-\tau$ (which is the difference of two independent random vectors both uniformly distributed in $\Lambda_{N}^{(0)}$) is in $\mathrm{B}_{1}$, and under this event if $y$ is a point in $\Lambda^{(0)}_{N}-\tau$, then $\bar{\mathrm{p}}_{N}(y)=\tilde{\mathrm{p}}_{N}(y+\tau)=C^{(0)}_{N}-(y+\tau)$, consequently if $x$ is a lattice point in $\mathbb{Z}^{\mathsf{d}}$ such that $x+U\in\Lambda^{(0)}_{N}-\tau$, then we have:
$x+U+\bar{\mathrm{p}}_{N}(x+U)=x+U+\tilde{\mathrm{p}}_{N}(x+U+\tau)=x+U+{C^{(0)}_{N}}-(x+U+\tau)={C^{(0)}_{N}-\tau}\in\mathrm{B}_{1}.$
By a simple lattice point counting, and since we take $N\geq 10$, there exists a constant $c$ depending only on the dimension such that for all values of $\tau$ and $U$ there are at least $cN^{\mathsf{d}}$ points of $\mathbb{Z}^{\mathsf{d}}+U$ in the cube $\Lambda^{(0)}_{N}-\tau$. In conclusion, under the event $\{C^{(0)}_{N}-\tau\in\mathrm{B}_{1}\}$, the number of points of $\mathbf{PL}^{\prime}_{N}$ in $\mathrm{B}_{1}$ is larger than $cN^{\mathsf{d}}$. The second moment of the number of perturbed points falling in $\mathrm{B}_{1}$ is thus bounded below by:
$\mathbb{E}[|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{1}|^{2}]\geq(cN^{\mathsf{d}})^{2}\times\mathbb{P}(\{C^{(0)}_{N}-\tau\in\mathrm{B}_{1}\}\})\geq c^{\prime}N^{\mathsf{d}}.$ (4.1)
We are now ready to construct our example.
Infinite number variance without $\mathsf{d}$-th moment. If $X$ is a non-negative random variable, we let $\mathcal{N}:=\frac{1}{\sqrt{\mathsf{d}}}\lfloor X\rfloor$ and we define our perturbed lattice as $\mathbf{PL}:=\mathbf{PL}^{\prime}_{\mathcal{N}}$, so that $\mathbf{PL}$ is a mixture of the perturbed lattices $\mathbf{PL}^{\prime}_{N}$ defined above, the parameter $N$ being chosen at random according to $\frac{1}{\sqrt{\mathsf{d}}}\lfloor X\rfloor$.
On the one hand, since the perturbations in $\mathbf{PL}^{\prime}_{N}$ are bounded almost surely by $\sqrt{\mathsf{d}}N$ for all $N$, the size of a single perturbation $\mathrm{p}_{0}$ in $\mathbf{PL}$ is bounded by $X$ almost surely. On the other hand, using (4.1) we see that:
$\mathbb{E}[|\mathbf{PL}\cap\mathrm{B}_{1}|^{2}]\geq c^{\prime}\sum_{N\geq 10}N^{\mathsf{d}}\times\mathbb{P}[\mathcal{N}=N],\quad\text{ with }\mathcal{N}=\frac{1}{\sqrt{\mathsf{d}}}\lfloor X\rfloor,$ (4.2)
and thus if $X$ fails to have a finite moment of order $\mathsf{d}$ then certainly the variance of $|\mathbf{PL}\cap\mathrm{B}_{1}|$ is infinite. ∎
4.2 Arbitrarily slow decay of the rescaled number variance: proof of Theorem 3
The processes that we just used for the proof of Proposition 4.1 can also be used to construct hyperuniform processes with arbitrarily slow decay of the rescaled number variance.
Proposition 4.2.
Let $\tilde{\sigma}:(1,+\infty)\to(0,+\infty)$ be a non-increasing function such that $\lim_{r\to\infty}\tilde{\sigma}(r)=0$. There exists a $\mathbb{Z}^{\mathsf{d}}$-invariant perturbation field $(\mathrm{p}_{x})_{x\in\mathbb{Z}^{\mathsf{d}}}$ such that:
  1. 1.
    The resulting perturbed lattice $\mathbf{PL}$ is hyperuniform (i.e. $\lim_{r\to\infty}\sigma(r)=0$).
  2. 2.
    There exist $c>0,M\geq 1$ (depending on $\mathsf{d}$ and $\sigma$) such that $\mathbf{PL}$ satisfies, for $r\geq 1$:
    $\sigma(r):=\frac{\mathrm{Var}[|\mathbf{PL}\cap\mathrm{B}_{r}|]}{|\mathrm{B}_{r}|}\geq c\tilde{\sigma}(Mr).$ (4.3)
Proof of Proposition 4.2.
The letters $c,c^{\prime},c^{\prime\prime}$… denote small constants that may change from line to line.
We use the same family of processes $\{\mathbf{PL}^{\prime}_{N}\}_{N\geq 1}$ as in the proof of Proposition 4.1. With a similar argument as the one leading to (4.1), we find that if $r\geq 1$ and $N\geq 10r$, we have:
$\mathbb{E}[|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|^{2}]-|\mathrm{B}_{r}|^{2}\geq cN^{2\mathsf{d}}\times\mathbb{P}(\{C^{(0)}_{N}-\tau\in\mathrm{B}_{r}\}\})-|\mathrm{B}_{r}|^{2}\geq c^{\prime}N^{\mathsf{d}}r^{\mathsf{d}}-|\mathrm{B}_{1}|^{2}r^{2\mathsf{d}},$
because ($\mathbb{P}(\{{C^{(0)}_{N}-\tau}\in\mathrm{B}_{r}\})$ is larger than $c\frac{r^{\mathsf{d}}}{N^{\mathsf{d}}}$ and under this event, all the $N^{\mathsf{d}}$ points of $\Lambda_{N}-\tau$ will “jump” inside $\mathrm{B}_{r}$. Taking $M$ large enough (depending only on $\mathsf{d}$) we thus see that for $N\geq Mr$, we have:
$\mathrm{Var}[|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|]\geq c^{\prime\prime}N^{\mathsf{d}}r^{\mathsf{d}}.$ (4.4)
Now let us consider a mixture of the $\{\mathbf{PL}^{\prime}_{N}\}_{N\geq 1}$ of the form $\mathbf{PL}:=\sum_{N\geq 1}\alpha_{N}\mathbf{PL}^{\prime}_{N}$, with coefficients $\alpha_{N}$ chosen as:
$\alpha_{N}:=\frac{1}{S}\frac{-\tilde{\sigma}^{\prime}_{N}}{N^{\mathsf{d}}},\quad\tilde{\sigma}^{\prime}_{N}:=\tilde{\sigma}(N+1)-\tilde{\sigma}(N),\quad S:=\sum_{N\geq 1}\frac{-\tilde{\sigma}^{\prime}_{N}}{N^{\mathsf{d}}}.$
(Note that thanks to a summation by parts, the series $\sum_{N\geq 1}\frac{-\tilde{\sigma}^{\prime}_{N}}{N^{\mathsf{d}}}$ has the same nature as $\sum_{N\geq 1}\frac{\tilde{\sigma}_{N}}{N^{\mathsf{d}+1}}$, which converges for all $\mathsf{d}\geq 1$ because $(\tilde{\sigma}_{N})_{N}$ is bounded, and thus $S$ is well-defined.) Using (4.4) and computing a telescopic sum, we obtain:
$\frac{\mathrm{Var}[|\mathbf{PL}\cap\mathrm{B}_{r}|]}{|\mathrm{B}_{r}|}\geq c\sum_{N\geq Mr}\alpha_{N}N^{\mathsf{d}}=\frac{c}{S}\sum_{N\geq Mr}-\tilde{\sigma}^{\prime}_{N}\geq c\tilde{\sigma}(Mr),$
for some positive constant $c$, which gives (4.3). It remains to prove that $\mathbf{PL}$ is hyperuniform.
For $\mathsf{d}=1,2$ we can compute, using again a telescopic sum, the $\mathsf{d}$-th moment of the perturbations:
$\sum_{N\geq 1}\alpha_{N}\times N^{\mathsf{d}}=\frac{1}{S}\times\sum_{N\geq 1}-\tilde{\sigma}^{\prime}_{N}<+\infty,$
and thus (by our main result) $\mathbf{PL}$ is indeed hyperuniform. This condition is not sufficient in dimension $\mathsf{d}\geq 3$, but we can instead directly find an upper bound on the number variance. Let us return to the “building block” $\mathbf{PL}^{\prime}_{N}$ and observe that:
  • If $N\geq r$, then $\mathrm{Var}[|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|]=\mathcal{O}(r^{\mathsf{d}}N^{\mathsf{d}})$. Indeed, with probability $\mathcal{O}\left(\frac{r^{\mathsf{d}}}{N^{\mathsf{d}}}\right)$ the ball $\mathrm{B}_{r}$ receives $N^{\mathsf{d}}$ points, and it loses all of its $\mathcal{O}(r^{\mathsf{d}})$ points with probability $1-\mathcal{O}\left(\frac{r^{\mathsf{d}}}{N^{\mathsf{d}}}\right)$.
  • If $r\geq N$ we have:
    $\mathrm{Var}[|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|]\preceq\frac{r^{\mathsf{d}-1}}{N^{\mathsf{d}-1}}\times N^{2\mathsf{d}}=\mathcal{O}\left(r^{\mathsf{d}-1}N^{\mathsf{d}+1}\right).$
    Indeed, there are $\mathcal{O}\left(\frac{r^{\mathsf{d}-1}}{N^{\mathsf{d}-1}}\right)$ hypercubes of sidelength $N$ intersecting the hypersphere $\partial\mathrm{B}_{r}$, and those are the only ones contributing (independently) to the variance, each one with a contribution $\mathcal{O}(N^{2\mathsf{d}})$.
Summing over $N$, we obtain:
$\frac{\mathrm{Var}[|\mathbf{PL}\cap\mathrm{B}_{r}|]}{|\mathrm{B}_{r}|}\preceq\frac{1}{r}\sum_{N\leq r}\alpha_{N}N^{\mathsf{d}+1}+\sum_{N\geq r}\alpha_{N}N^{\mathsf{d}}=\frac{1}{r}\sum_{N\leq r}-\tilde{\sigma}^{\prime}_{N}N+\sum_{N\geq r}-\tilde{\sigma}^{\prime}_{N}\preceq\frac{1}{r}\sum_{N\leq r}\tilde{\sigma}_{N}+\tilde{\sigma}(r)\to 0,$
so the process that we constructed is indeed hyperuniform. ∎
4.3 Arbitrarily small perturbations with large number variance
We now prove the fourth item of Theorem 1 as well as the second item of Theorem 2.
Proposition 4.3.
Let $\varepsilon>0$.
  • If $\mathsf{d}\geq 3$, there are perturbations of $\mathbb{Z}^{\mathsf{d}}$ that are almost surely bounded by $\varepsilon$ and such that
    $\lim_{r\to+\infty}\frac{\mathrm{Var}\left[|\mathbf{PL}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=+\infty$
  • If $\mathsf{d}=2$, there are perturbations of $\mathbb{Z}^{2}$ that are almost surely bounded by $\varepsilon$ and such that $\mathbf{PL}$ is not hyperuniform of class I.
Proof.
Let $N\geq 10$ and let $(C^{(k)}_{N})_{k\in N\mathbb{Z}^{\mathsf{d}}}$ be independent random variables such that $C^{(k)}_{N}$ is uniformly distributed in $\Lambda^{(k)}_{N}$ for all $k\in N\mathbb{Z}^{\mathsf{d}}$. We define a first “displacement field” $\tilde{\mathrm{p}}_{N}:\mathbb{R}^{\mathsf{d}}\mapsto\mathbb{R}^{\mathsf{d}}$ by setting:
$\tilde{\mathrm{p}}_{N}(x):=\varepsilon\frac{x-C^{(k)}_{N}}{|x-C^{(k)}_{N}|}\text{ for all $k\in N\mathbb{Z}^{\mathsf{d}}$ and all $x$ in $\Lambda^{(k)}_{N}$. }$
Applying $\tilde{\mathrm{p}}_{N}$ corresponds to choosing a center $C^{(k)}_{N}$ uniformly at random in each cube $\Lambda^{(k)}_{N}$ and then push every point in $\Lambda^{(k)}_{N}$ by an amount $\varepsilon$ away from $C^{(k)}_{N}$. This displacement field is clearly $N\mathbb{Z}^{\mathsf{d}}$-stationary, and we can turn it into a $\mathbb{R}^{\mathsf{d}}$-stationary field by choosing a random vector $\tau$ uniformly distributed in $\Lambda_{N}$ and independent of $\tilde{\mathrm{p}}_{N}$ and setting $\bar{\mathrm{p}}_{N}:=\tilde{\mathrm{p}}_{N}(\cdot+\tau)$. Perturbations are bounded by $\varepsilon$. Finally, we set for a $[0,1]^{\mathsf{d}}$-uniform random variable $U$ (idependent of everything else):
$\mathbf{PL}^{\prime}_{N}:=\{x+U+\bar{\mathrm{p}}_{N}(x+U),\ x\in\mathbb{Z}^{\mathsf{d}}\}.$
Next, let $r$ be such that ${N\geq r}$. We want to count the number of points of $\mathbf{PL}^{\prime}_{N}$ falling in $\mathrm{B}_{r}$, and we will show that a significant amount of the original lattice points are “leaving” the ball with positive probability, creating a very large number variance if $\mathsf{d}\geq 3$ and a critical number variance if $\mathsf{d}=2$. We rely on the following elementary computation.
Claim 4.4.
Let $r>0,x,u,\tau,C\in\mathbb{R}^{\mathsf{d}}$ be such that:
$\frac{9}{10}r\leq|x+u|\leq\frac{11}{10}r,\quad|\tau-C|\leq\frac{1}{10}r.$
Then we have:
$\left|x+u+\varepsilon\frac{x+u+\tau-C}{|x+u+\tau-C|}\right|\geq|x+u|+\frac{\varepsilon}{3}.$ (4.5)
Proof of Claim 4.4.
We simply compute:
$\left|x+u+\varepsilon\frac{x+u+\tau-C}{|x+u+\tau-C|}\right|^{2}=|x+u|^{2}+\varepsilon^{2}+2\varepsilon(x+u)\cdot\frac{x+u+\tau-C}{|x+u+\tau-C|}.$
With our assumptions, we know that $\frac{8}{9}|x+u|\leq|x+u+\tau-C|\leq\frac{10}{9}|x+u|$, and thus:
$(x+u)\cdot\frac{x+u+\tau-C}{|x+u+\tau-C|}=|x+u|^{2}\frac{1}{|x+u+\tau-C|}+(x+u)\cdot(\tau-C)\frac{1}{|x+u+\tau-C|}\\ \geq|x+u|^{2}\times\frac{9}{10|x+u|}-|x+u|\times\frac{1}{10}r\times\frac{9}{8|x+u|}\geq\frac{9}{10}\times\frac{9}{10}r-\frac{9}{80}r\geq\frac{1}{2}r.$
As a consequence, we see that (using again $|x+u|\leq\frac{11}{10}r$):
$\left|x+u+\varepsilon\frac{x+u+\tau-C}{|x+u+\tau-C|}\right|^{2}\geq|x+u|^{2}+\varepsilon r+\varepsilon^{2}\geq|x+u|^{2}+2\frac{10}{2\times 11}|x+u|\varepsilon+\varepsilon^{2}\geq\left(|x+u|+\frac{\varepsilon}{3}\right)^{2},$
which proves the claim. ∎
Since $\tau$ and $C^{(0)}_{N}$ are uniformly distributed in $\Lambda_{N}$ and independent, there exists $c>0$ depending only on the dimension, such that:
($\mathbb{P}\left(\left\{|\tau|\leq\frac{r}{10}\right\}\cap\left\{\left|\tau-C^{(0)}_{N}\right|\leq\frac{r}{10}\right\}\right)\geq c{\frac{r^{2\mathsf{d}}}{N^{2\mathsf{d}}}}.$ (4.6)
This event is interesting for us due to the following observation:
Claim 4.5.
On the event $\left\{|\tau|\leq\frac{r}{10}\right\}\cap\left\{\left|\tau-C^{(0)}_{N}\right|\leq\frac{r}{10}\right\}$, for all $u\in[0,1]^{\mathsf{d}}$, if $\frac{9}{10}r\leq|x+u|\leq\frac{11}{10}r$ then $|x+u+\bar{\mathrm{p}}(x+u)|\geq|x+u|+\frac{\varepsilon}{3}$.
In particular, if $x$ is such that $|x+u|\geq r-\frac{\varepsilon}{4}$ then $x+u+\bar{\mathrm{p}}(x+u)\notin\mathrm{B}_{r}$.
Proof of Claim 4.5.
If $|x+u|\leq\frac{11}{10r}$ then $|x+u+\tau|\leq\frac{12}{10}r$, which means that $x+u+\tau\in\Lambda^{(0)}_{N}$ (because $N\geq 5r$) and thus, by construction of the displacement field:
$\bar{\mathrm{p}}(x+u)=\tilde{\mathrm{p}}_{N}(x+u+\tau)=\varepsilon\frac{x+u+\tau-C^{(0)}_{N}}{|x+u+\tau-C^{(0)}_{N}|}.$
We may then apply Claim 4.4 with $C=C^{(0)}_{N}$, which yields $|x+u+\bar{\mathrm{p}}(x+u)|\geq|x+u|+\frac{\varepsilon}{3}$. Next, let $x$ be such that $|x+u|\geq r-\frac{\varepsilon}{4}$.
  1. 1.
    If $|x+u|\geq\frac{11}{10}r$ then $x+u+\bar{\mathrm{p}}(x+u)$ cannot be in $\mathrm{B}_{r}$ because perturbations are bounded by $\varepsilon<\frac{1}{10}r$.
  2. 2.
    If $|x+u|\leq\frac{11}{10}r$, then since by assumption $|x+u|\geq r-\frac{\varepsilon}{4}\geq\frac{9}{10}r$ we get from above that:
    $|x+u+\bar{\mathrm{p}}(x+u)|\geq|x+u|+\frac{\varepsilon}{3}\geq r-\frac{\varepsilon}{4}+\frac{\varepsilon}{3}>r$
    and thus $x+u+\bar{\mathrm{p}}(x+u)$ is not in $\mathrm{B}_{r}$.
From Claim 4.5 we deduce that on the event $\left\{|\tau|\leq\frac{r}{10}\right\}\cap\left\{\left|\tau-C^{(0)}_{N}\right|\leq\frac{r}{10}\right\}$ we have:
[…]
We are thus left with finding an upper bound for the number of integer points in the ball $\mathrm{B}_{r-\frac{\varepsilon}{4}}(-U)$, which we expect from the continuous approximation to be of order $|\mathrm{B}_{r}|-Cr^{\mathsf{d}-1}\varepsilon$, and thus significantly smaller than $|\mathrm{B}_{r}|$. Although the quest for sharp bounds seems to still be an active area of research, “basic” estimates are sufficient for us and have thankfully been known for a long time - as the classical result of Landau [Lan15], which gives:
$\sup_{a\in\mathbb{R}^{\mathsf{d}}}\left|\left|\left\{x\in\mathbb{Z}^{\mathsf{d}},x\in\mathrm{B}_{r}(a)\right\}\right|-|\mathrm{B}_{r}|\right|\leq C_{\mathsf{d}}r^{\frac{\mathsf{d}(\mathsf{d}-1)}{\mathsf{d}+1}}\ll r^{\mathsf{d}-1},$ (4.8)
so in particular it ensures that for all $u\in[0,1]^{d}$ we have:
|$\left|\left\{x\in\mathbb{Z}^{\mathsf{d}},|x+u|\leq r-\frac{\varepsilon}{4}\right\}\right|=|\mathrm{B}_{r-\frac{\varepsilon}{4}}|+o(r^{\mathsf{d}-1})\leq|\mathrm{B}_{r}|-C_{\mathsf{d}}r^{\mathsf{d}-1}\varepsilon+o(r^{\mathsf{d}-1}),$ (4.9)
so in particular $|\mathrm{B}_{r}|-\left|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}\right|\geq C^{\prime}_{\mathsf{d}}r^{\mathsf{d}-1}\varepsilon$. From this and our lower bound (4.6) on the probability of $\left\{|\tau|\leq\frac{r}{10}\right\}\cap\left\{\left|\tau-C^{(0)}_{N}\right|\leq\frac{r}{10}\right\}$ we deduce:
For r,N large enough with $\text{For $r,N$ large enough {with $N\geq r$},}\quad\mathbb{E}\left[\left||\mathrm{B}_{r}|-|\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|\right|^{2}\right]\geq c_{\mathsf{d}}r^{2\mathsf{d}-2}{\times\frac{r^{2\mathsf{d}}}{N^{2\mathsf{d}}}},$ (4.10)
for some positive constant $c_{\mathsf{d}}$ depending on the dimension. We are now ready to construct our examples.
Non-hyperuniformity for $\mathsf{d}\geq 3$. If $\mathsf{d}\geq 3$, fix some integer-valued random variable $\mathcal{N}$ with $\mathbb{P}[\mathcal{N}=N]$ proportional to $\frac{1}{N^{1+\delta}}$ for $N\geq 1$ (where $\delta\in(0,1)$ is arbitrary), and set $\mathbf{PL}:=\mathbf{PL}^{\prime}_{\mathcal{N}}$ i.e. $\mathbf{PL}$ is a mixture of the perturbed lattices $\mathbf{PL}^{\prime}_{N}$ defined above, the parameter $N$ being chosen at random. For any $r$, a lower bound on the number variance in $\mathrm{B}_{r}$ is given by:
$\mathbb{E}\left[\left||\mathbf{PL}\cap\mathrm{B}_{r}|-|\mathrm{B}_{r}|\right|^{2}\right]\geq\sum_{N\geq r}\mathbb{E}\left[\left||\mathbf{PL}^{\prime}_{N}\cap\mathrm{B}_{r}|-|\mathrm{B}_{r}|\right|^{2}\right]\mathbb{P}[\mathcal{N}=N].$ (4.11)
From (4.10) we deduce that, for $r$ large enough (depending on $\mathsf{d}$):
$\frac{\mathrm{Var}[|\mathbf{PL}\cap\mathrm{B}_{r}|]}{|\mathrm{B}_{r}|}\geq c_{\mathsf{d}}\frac{r^{2\mathsf{d}-2}}{r^{\mathsf{d}}}\times\sum_{5r\geq N\geq r}\frac{1}{N^{1+\delta}},$
which after a simple computation is seen to be greater than $r^{2\mathsf{d}-2-\mathsf{d}-\delta}$, which is $\gg 1$ for all $\mathsf{d}\geq 3$.
Non-class-I hyperuniformity for $\mathsf{d}=2$. The same construction for $\mathsf{d}=2$ yields $\frac{\mathrm{Var}[|\mathbf{PL}\cap\mathrm{B}_{r}|]}{|\mathrm{B}_{r}|}\geq r^{-\delta}$ where $\delta>0$ can be arbitrary small, hence this process is not class I hyperuniform.
Note that one could go further and consider coefficients $\mathbb{P}[\mathcal{N}=N]$ proportional to $\frac{1}{N\log^{1+\delta}N}$, which would yield a decay of the rescaled number variance slower than any power law. ∎
Appendix A Control on the term $B_{3}$: proof of Proposition 3.7
In this section, we introduce two additional lengthscales: a very small one $\tau$ and a very large one $\bar{d}$, both depending on the parameter $r$ (and on the other lengthscale $\ell$ which satisfies $\ell r\to+\infty$), chosen as:
$\tau:=r^{-0.1}\ll 1,\quad\bar{d}:=\sqrt{r\ell^{-1}}\text{ so that }\ell^{-1}\ll\bar{d}\ll r.$ (A.1)
Note that since $\ell^{-1}\leq\frac{1}{100}r$ (see (3.6)) we have $\bar{d}\geq 10\ell^{-1}$.
Recall that $B_{3}=\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\mathcal{B}_{3}(x)$, with $\mathcal{B}_{3}(x)$ defined by:
$\mathcal{B}_{3}(x):=\int_{t\in\mathbb{R}^{2}}{e^{-it\cdot x}}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))\mathtt{d}t\\ =\mathbb{E}\left[\widehat{j_{r}(1-\chi_{\ell})}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}(1-\chi_{\ell})}(x)\right]$ (A.2)
We will control $\mathcal{B}_{3}(x)$ in various ways, depending on the distance $d=d(x)$ from $x$ to the circle $\partial\mathrm{B}_{2r}$ (which we refer to as “the boundary”). We distinguish three cases:
  • $x$ “far from the boundary” i.e. $d\geq\bar{d}$, which we deal with in Section A.1.
  • $x$ “close, but not too close to the boundary” i.e. $\tau\leq d\leq\bar{d}$, which we deal with in Sections A.2 to A.6.
  • $x$ “extremely close to the boundary” i.e. $d\leq\tau$, which will be a consequence of Lemma 2.2, see the end of Section A.7.
A.1 Far from the boundary
In this section, we focus on lattice points $x$, both inside and outside $\mathrm{B}_{2r}$, such that $d\geq\bar{d}$ and we prove the following result:
Lemma A.1.
$\sum_{x\in\mathcal{L}\setminus\{0\},d\geq\bar{d},|x|\leq 100r}\mathcal{B}_{3}(x)\preceq\frac{1}{\ell\bar{d}}+\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{\bar{d}}{2}}\right].$
Both terms go to zero as $r\to\infty$ because $\bar{d}\gg\ell^{-1}\gg 1$ and $\mathrm{p}_{0}$ has a finite second moment.
Proof of Lemma A.1.
Let $\mathcal{A}:=\widehat{j_{r}}-\frac{1}{2\pi}\widehat{j_{r}}\ast\widehat{\chi_{\ell}}$. Using the identity $\widehat{j_{r}(1-\chi_{\ell})}=\widehat{j_{r}}-\frac{1}{2\pi}\widehat{j_{r}}\ast\widehat{\chi_{\ell}}$ (the $2\pi$ factor is due to our Fourier transform convention), we have for all $x$:
$\mathcal{B}_{3}(x)=\mathbb{E}\left[\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)\right].$
For convenience, we denote $K_{\ell}:=\frac{1}{2\pi}\widehat{\chi_{\ell}}$ and use $K_{\ell}$ instead of $\widehat{\chi_{\ell}}$ in the computations below.
Heuristics
The map $\widehat{j_{r}}$ is “essentially” smooth, with a $k$-th derivative of order $r^{-k}$, and on the other hand $K_{\ell}$ is smooth, living “essentially” at scale $\mathcal{O}(\ell^{-1})$ and $\int_{\mathbb{R}^{2}}K_{\ell}=\frac{1}{2\pi}\int\widehat{\chi_{\ell}}=\chi_{\ell}(0)=1$, hence the convolution with $K_{\ell}$ represents a (signed) average of values at distance $\mathcal{O}(\ell^{-1})$. Applying a second order Taylor expansion and using the symmetry of $K_{\ell}$ to kill the first order contribution, one gets for all $t\in\mathbb{R}^{2}$ that $|(\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell})(t)|$ is $\mathcal{O}(\ell^{-2}r^{-2})$ and moreover:
$|\mathrm{Hess}(\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell})(t)|=\mathcal{O}(\ell^{-2}r^{-4}),$ (A.3)
gaining two powers of $r$ thanks to the two derivatives. Then for each $x\in\mathcal{L}$, applying a second order Taylor expansion around $x$ and using the fact that $\mathbb{E}[\mathrm{p}_{x}-\mathrm{p}_{0}]=0$ to kill the first order contribution, we obtain:
$\mathbb{E}\left[(\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell})(x+\mathrm{p}_{x}-\mathrm{p}_{0})-(\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell})(x)\right]=\mathcal{O}(\ell^{-2}r^{-4})\mathbb{E}[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}].$ (A.4)
We can then sum (A.4) over $\mathcal{O}(r^{2})$ terms, yielding an error of size $\mathcal{O}((r\ell)^{-2})=o_{r}(1)$. The main problem with this argument is that $\widehat{j_{r}}$ fails to be differentiable in two ways: at the origin, and along $\partial\mathrm{B}_{2r}$, so that (A.3) is not true everywhere and we need to be more precise.
Regularity of $\mathcal{A}$.
The map $\mathsf{1}_{\mathrm{B}_{1}}\ast\mathsf{1}_{\mathrm{B}_{1}}$ is a radially symmetric function (sometimes called the “Euclid’s hat”) given by (see [Gne99, eq. (4)]):
$\left(\mathsf{1}_{\mathrm{B}_{1}}\ast\mathsf{1}_{\mathrm{B}_{1}}\right)(t)=\int_{\frac{|t|}{2}}^{1}\sqrt{1-v^{2}}\mathtt{d}v\ \text{ if }|t|\leq 2,\quad\ 0\text{ otherwise}.$
This map is continuous on $\mathbb{R}^{2}$ and $C^{\infty}$ on $\mathbb{R}^{2}\setminus\mathrm{B}_{2}$ and on $\mathrm{B}_{2}\setminus\{0\}$. We have $\mathsf{1}_{\mathrm{B}_{r}}\ast\mathsf{1}_{\mathrm{B}_{r}}=r^{2}\left(\mathsf{1}_{\mathrm{B}_{1}}\ast\mathsf{1}_{\mathrm{B}_{1}}\right)(\frac{\cdot}{r})$ and thus the function $\widehat{j_{r}}$ is proportional to $\left(\mathsf{1}_{\mathrm{B}_{1}}\ast\mathsf{1}_{\mathrm{B}_{1}}\right)(\frac{\cdot}{r})$. As a consequence:
  • The map $\widehat{j_{r}}$ is Lipschitz on $\mathbb{R}^{2}$ with a $\mathcal{O}(r^{-1})$ Lipschitz constant. We will simply refer to this as the fact that $\widehat{j_{r}}$ is “$r^{-1}$-Lipschitz”.
  • Since $\widehat{j_{r}}$ is $r^{-1}$-Lipschitz, so is $\widehat{j_{r}}\ast K_{\ell}$ and thus so is $\mathcal{A}:=\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell}$.
  • The map $\widehat{j_{r}}$ is smooth on the punctured open disk $\mathring{\mathrm{B}_{2r}}\setminus\{0\}$ and its $\mathsf{k}$-th derivative is bounded by $C_{k}r^{-\mathsf{k}}$ for $\mathsf{k}\geq 1$.
  • If $x,y$ are in $\mathring{\mathrm{B}_{2r}}\setminus\{0\}$, we have:
    $\widehat{j_{r}}(y)-\widehat{j_{r}}(x)=\nabla\widehat{j_{r}}(x)\cdot(y-x)+\mathrm{Hess}\widehat{j_{r}}(x)(y-x,y-x)+\mathcal{O}(r^{-3}|y-x|^{3}).$ (A.5)
    Indeed, if the line segment joining $x$ and $y$ is contained in $\mathring{\mathrm{B}_{2r}}\setminus\{0\}$ this is a standard second-order Taylor’s expansion, and by continuity it extends to the odd case where $(x,y)$ contains 00. Still by continuity, we can allow $y=0$, but for definitness of $\nabla\widehat{j_{r}}(x)$ and $\mathrm{Hess}\widehat{j_{r}}(x)$, the point $x$ should not be 00.
  • For similar reasons, if $x,y$ are in $\mathring{\mathrm{B}_{2r}}\setminus\{0\}$, we have:
    $\mathrm{Hess}\widehat{j_{r}}(y)-\mathrm{Hess}\widehat{j_{r}}(x)=\mathcal{O}(r^{-3}|y-x|).$ (A.6)
Truncating $\mathrm{p}_{x}-\mathrm{p}_{0}$.
Recall that $\mathcal{B}_{3}(x)=\mathbb{E}\left[\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)\right]$.
Claim A.2.
For all $x\in\mathcal{L}$, denoting $d:=\mathrm{dist}(x,\partial\mathrm{B}_{2r})$, we have (uniformly in $x\in\mathcal{L}$):
$\mathcal{B}_{3}(x)=\mathbb{E}\left[\left(\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\leq{\frac{d}{2}}}\right]{+\mathcal{O}(r^{-1})\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]}.$ (A.7)
Proof of Claim A.2.
Since $\mathcal{A}$ is $r^{-1}$-Lipschitz, we have $\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)=\mathcal{O}(r^{-1})\times|\mathrm{p}_{x}-\mathrm{p}_{0}|$, and thus:
$\mathbb{E}\left[\left(\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq{\frac{d}{2}}}\right]\preceq r^{-1}\times\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq{\frac{d}{2}}}\right].$
Moreover, using an elementary inequality and the fact that $\mathrm{p}_{x}$ and $\mathrm{p}_{0}$ are identically distributed, we have:
$\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq{\frac{d}{2}}}\right]\leq 2\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq{\frac{d}{4}}}\right].$
Cost of discarding bad regions.
Viewing $\widehat{j_{r}}\ast K_{\ell}$ as an average (with signed weights) of $\widehat{j_{r}}$, we want to know the cost of discarding various regions on which estimates are difficult. Recall that $d$ denotes the distance of the lattice point $x$ to $\partial\mathrm{B}_{2r}$ and that $\bar{d}$ is a length-scale chosen in (A.1).
Claim A.3 (Discarding a $p$-neighbhorood of the boundary).
Assume $d\geq\bar{d}$, let $p$ be in $(0,\frac{d}{2})$. We have:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p}\mathtt{d}u=\mathcal{O}\left(\frac{p}{\ell d^{2}}\right).$ (A.8)
Proof of Claim A.3.
Let us re-write the integral over the annulus $\{u,\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p\}$ using angle + distance to the boundary” coordinates as follows:
$\int_{u,\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p}|K_{\ell}(x-u)|\mathtt{d}u{\preceq}\int_{v\in\partial\mathrm{B}_{2r}}\int_{s=-p}^{p}\left|K_{\ell}\left(x-v\left(1+\frac{s}{|v|}\right)\right)\right|\mathtt{d}s\mathtt{d}v.$
Within $\partial\mathrm{B}_{2r}$ there is an arc of length $\mathcal{O}(d)$ of points $v$ such that $d\leq|x-v|\leq 10d$, and since $p\leq\frac{d}{2}$, if $|s|\leq p$, then we can guarantee that: $|x-v\left(1+\frac{s}{|v|}\right)|\geq\frac{d}{2}$. Moreover, we assume that $d{\geq\bar{d}\geq}10\ell^{-1}$, so we can use the decay properties of $\widehat{\chi_{\ell}}$ (thus of $K_{\ell}$) as in (3.7) (taking $m=3$), which give:
$K_{\ell}\left(x-v\left(1+\frac{s}{|v|}\right)\right)\preceq\frac{\ell^{2}}{\ell^{3}\left|x-v\left(1+\frac{s}{|v|}\right)\right|^{3}}\preceq\frac{1}{\ell d^{3}}.$
We can thus evaluate the contribution of this arc as:
$\int_{v\in\partial\mathrm{B}_{2r},|x-v|\leq 10d}\int_{s=-p}^{p}\left|K_{\ell}\left(x-v\left(1+\frac{s}{|v|}\right)\right)\right|\mathtt{d}s\mathtt{d}v\preceq d\times p\times\frac{1}{\ell d^{3}}=\mathcal{O}\left(\frac{p}{\ell d^{2}}\right).$ (A.9)
On the other hand, if $T:=|x-v|\geq 10d$ and $|s|\leq p\leq\frac{d}{2}$ we have $\left|x-v\left(1+\frac{s}{|v|}\right)\right|\geq\frac{T}{2}$, so the contribution of the remaining part of $\partial\mathrm{B}_{2r}$ is:
$\int_{v\in\partial\mathrm{B}_{2r},|x-v|\geq 10d}\int_{s=-p}^{p}\left|K_{\ell}\left(x-v\left(1+\frac{s}{|v|}\right)\right)\right|\mathtt{d}s\mathtt{d}v\preceq\int_{T=10d}^{r}p\times\frac{\ell^{2}}{\ell^{3}T^{3}}\mathtt{d}T,$
using again (3.7) (with $m=3$), which is of the same order as (A.9). ∎
Claim A.4 (Discarding the other side).
We still assume $d\geq\bar{d}$. We have:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\notin\mathrm{B}_{2r}}\mathtt{d}u=\mathcal{O}\left(\frac{1}{\ell d}\right)\text{ (for $x\in\mathrm{B}_{2r}$)},\quad\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\in\mathrm{B}_{2r}}\mathtt{d}u=\mathcal{O}\left(\frac{1}{\ell d}\right)\text{ (for $x\notin\mathrm{B}_{2r}$)}.$
Proof of Claim A.4.
If $x$ is inside $\mathrm{B}_{2r}$, the points $u\notin\mathrm{B}_{2r}$ are at distance at least $d$ from $x$ (since $d$ represents the distance between $x$ and $\partial\mathrm{B}_{2r}$) and we may thus use the decay of $\widehat{\chi_{\ell}}$ (thus of $K_{\ell}$) to write:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\notin\mathrm{B}_{2r}}\mathtt{d}u\preceq\int_{T\geq d}\frac{\ell^{2}}{\ell^{3}T^{3}}\times T\mathtt{d}T=\mathcal{O}\left(\frac{1}{\ell d}\right),$
as desired. Next, if $x$ is outside $\mathrm{B}_{2r}$ then any $u\in\mathrm{B}_{2r}$ is at distance at least $d$ from $x$. If $d\leq 10r$, we proceed as before, and write:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\in\mathrm{B}_{2r}}\mathtt{d}u\preceq\int_{T\geq d}\frac{\ell^{2}}{\ell^{3}T^{3}}\times T\mathtt{d}T=\mathcal{O}\left(\frac{1}{\ell d}\right),$
using again the decay of $K_{\ell}$. If $d\geq 10r$ we simply write:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\in\mathrm{B}_{2r}}\mathtt{d}u\preceq r^{2}\times\frac{\ell^{2}}{\ell^{3}d^{3}}=\mathcal{O}\left(\frac{1}{\ell d}\right).$
Finally, we record a simple scaling estimate.
Claim A.5.
Let $x$ be arbitrary, we have:
$\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)||x-u|\mathtt{d}u=\mathcal{O}(\ell^{-1}).$
Proof of Claim A.5.
Applying the obvious change of variables and using the fact that $\widehat{\chi_{\ell}}=\ell^{2}\widehat{\chi}(\cdot\ell)$ for some Schwartz function $\widehat{\chi}$ yields the result. ∎
Analysis for $x$ inside $\mathrm{B}_{2r}$.
Let $x\neq 0$ be a point in $\mathrm{B}_{2r}$ with $d:=2r-|x|\geq\bar{d}\geq 10\ell^{-1}$. We want to bound the first term in the right-hand side of (A.7), namely:
$\mathbb{E}\left[\left(\mathcal{A}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\mathcal{A}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\leq{\frac{d}{2}}}\right].$
For simplicity, we denote here the random variable $\mathrm{p}_{x}-\mathrm{p}_{0}$ by $p$, and our goal is thus to compare $\mathcal{A}(x+p)$ and $\mathcal{A}(x)$, in expectation, under the condition $|p|\leq{\frac{d}{2}}$. In this paragraph, we prove:
Claim A.6.
$\mathbb{E}\left[\left(\mathcal{A}(x+p)-\mathcal{A}(x)\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}\right]=\mathcal{O}\left(\frac{1}{r\ell d^{2}}\right)+\mathcal{O}(r^{-1})\times\mathbb{E}\left[{|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}}\right].$ (A.10)
Proof.
Observe that if $|p|\leq{\frac{d}{2}}$, then the points $x$ and $x+p$ are both in $\mathring{B}_{2r}$, and so using (A.5) we know that the difference $\widehat{j_{r}}(x+p)-\widehat{j_{r}}(x)$ can be written as:
$\widehat{j_{r}}(x+p)-\widehat{j_{r}}(x)=\nabla\widehat{j_{r}}(x)\cdot p+\mathrm{Hess}\widehat{j_{r}}(x)(p,p)+\mathcal{O}(r^{-3}|p|^{3}).$ (A.11)
On the other hand, we can write $\widehat{j_{r}}\ast K_{\ell}(x+p)-\widehat{j_{r}}\ast K_{\ell}(x)$ as:
$\widehat{j_{r}}\ast K_{\ell}(x+p)-\widehat{j_{r}}\ast K_{\ell}(x)=\int_{\mathbb{R}^{2}}\left(\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)\right)K_{\ell}(x-u)\mathtt{d}u\\ =\int_{\mathbb{R}^{2}}\left(\nabla\widehat{j_{r}}(u)\cdot p+\mathrm{Rem}(u,p)\right)K_{\ell}(x-u)\mathtt{d}u,$ (A.12)
where $\mathrm{Rem}(u,p)$ is simply defined as the remainder $\mathrm{Rem}(u,p):=\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)-\nabla\widehat{j_{r}}(u)\cdot p$ (note that the fact that $\widehat{j_{r}}$ is not differentiable on $\{0\}\cup\partial\mathrm{B}_{2r}$ is here harmless because this set has measure zero).
Recall that $\mathcal{A}=\widehat{j_{r}}-\widehat{j_{r}}\ast K_{\ell}$. Combining (A.11) and (A.12), using the fact that $\int K_{\ell}(x-u)\mathtt{d}u=1$ for all $x$, we can write:
$\left(\mathcal{A}(x+p)-\mathcal{A}(x)\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}=\left(\nabla\widehat{j_{r}}(x)\cdot p-\int_{\mathbb{R}^{2}}\nabla\widehat{j_{r}}(u)\cdot pK_{\ell}(x-u)\mathtt{d}u\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}\\ +\left(\int_{\mathbb{R}^{2}}\left(\mathrm{Hess}\widehat{j_{r}}(x)(p,p)-\mathrm{Rem}(u,p)\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}K_{\ell}(x-u)\mathtt{d}u\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}+\mathcal{O}(r^{-3}|p|^{2}d),$ (A.13)
bounding $\mathcal{O}(r^{-3}|p|^{3})\mathsf{1}_{|p|\leq{\frac{d}{2}}}$ by $\mathcal{O}(r^{-3}|p|^{2}d)$.
Since the random variable $p$ is centered, we have:
$\mathbb{E}\left[\nabla\widehat{j_{r}}(x)\cdot p-\int_{\mathbb{R}^{2}}\nabla\widehat{j_{r}}(u)\cdot pK_{\ell}(x-u)\mathtt{d}u\right]=0,$
and since $\nabla\widehat{j_{r}}$ is $\mathcal{O}(r^{-1})$ we have:
$\mathbb{E}\left[\left(\nabla\widehat{j_{r}}(x)\cdot p-\int_{\mathbb{R}^{2}}\nabla\widehat{j_{r}}(u)\cdot pK_{\ell}(x-u)\mathtt{d}u\right)\mathsf{1}_{|p|\geq{\frac{d}{2}}}\right]=\mathcal{O}(r^{-1})\times\mathbb{E}\left[|p|\mathsf{1}_{|p|\geq\frac{d}{2}}\right].$
Moreover, by the same elementary inequalities as above, we know that:
$\mathbb{E}\left[|p|\mathsf{1}_{|p|\geq\frac{d}{2}}\right]\preceq\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]$
Taking the expectation in (A.13), we thus obtain:
$\mathbb{E}\left[\left(\mathcal{A}(x+p)-\mathcal{A}(x)\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}\right]=\mathbb{E}\left[\left(\int_{\mathbb{R}^{2}}\left(\mathrm{Hess}\widehat{j_{r}}(x)(p,p)-\mathrm{Rem}(u,p)\right)K_{\ell}(x-u)\mathtt{d}u\right)\mathsf{1}_{|p|\leq{\frac{d}{2}}}\right]\\ +\mathcal{O}(r^{-1})\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]+\mathcal{O}\left(r^{-3}d\right),$ (A.14)
using the fact that $|p|$ has a finite second moment (bounded by $4\mathbb{E}[|\mathrm{p}_{0}|^{2}])$.
We now focus on $\mathrm{Rem}(u,p)=\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)-\nabla\widehat{j_{r}}(u)\cdot p$ and distinguish three cases:
  • If $u\in\mathrm{B}_{2r}$ and $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|$ then we can use (A.5) again and write:
    $\mathrm{Rem}(u,p)=\mathrm{Hess}\widehat{j_{r}}(u)(p,p)+\mathcal{O}\left(r^{-3}|p|^{3}\right)$
    (strictly speaking, this is not valid if $u=0$ but that does not affect the integral). Using moreover (A.6) we can compare the Hessian at $u$ to the Hessian at $x$ and get:
    $\mathrm{Rem}(u,p)=\mathrm{Hess}\widehat{j_{r}}(x)(p,p)+\mathcal{O}\left(r^{-3}|x-u|\right)\times|p|^{2}+\mathcal{O}\left(r^{-3}|p|^{3}\right).$
    This is the “good situation”, as the first term in the right-hand side will cancel the $\mathrm{Hess}\widehat{j_{r}}(x)(p,p)$ term in the integrand of (A.14).
  • If $u\notin\mathrm{B}_{2r}$ and $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|$, then $\widehat{j_{r}}(u)=\widehat{j_{r}}(u+p)=0$ hence $\mathrm{Rem}(u,p)=0$.
  • Finally, if $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p$, the only regularity we can use is the $\mathcal{O}(r^{-1})$ Lipschitz bound and we write:
    $\mathrm{Rem}(u,p)=\mathcal{O}(r^{-1}|p|).$
Using those informations, we get:
$\int_{\mathbb{R}^{2}}\mathrm{Rem}(u,p)K_{\ell}(x-u)\mathtt{d}u\\ =\int_{\mathbb{R}^{2}}\left(\mathrm{Hess}\widehat{j_{r}}(x)(p,p)+\mathcal{O}\left(r^{-3}|x-u|\right)\times|p|^{2}+\mathcal{O}\left(r^{-3}|p|^{3}\right)\right)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}K_{\ell}(x-u)\mathtt{d}u\\ +\mathcal{O}(r^{-1}|p|)\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p}\mathtt{d}u.$
Recall also that $\mathrm{Hess}\widehat{j_{r}}=\mathcal{O}(r^{-2})$, so $\mathrm{Hess}\widehat{j_{r}}(x)(p,p)=\mathcal{O}(r^{-2}|p|^{2})$ for all $x,p$. We can bound the difference $\int_{\mathbb{R}^{2}}\left(\mathrm{Hess}\widehat{j_{r}}(x)(p,p)-\mathrm{Rem}(u,p)\right)K_{\ell}(x-u)$ by:
$\int_{\mathbb{R}^{2}}\left(\mathrm{Hess}\widehat{j_{r}}(x)(p,p)-\mathrm{Rem}(u,p)\right)K_{\ell}(x-u)=\mathcal{O}(r^{-2}|p|^{2})\times\left(1-\int_{\mathbb{R}^{2}}K_{\ell}(x-u)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}\mathtt{d}u\right)\\ +\mathcal{O}\left(r^{-3}|p|^{2}\right)\times\int_{\mathbb{R}^{2}}(|u-x|+|p|)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}|K_{\ell}|(x-u)\mathtt{d}u\\ +\mathcal{O}(r^{-1}|p|)\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p}\mathtt{d}u.$ (A.15)
We now control each term in the right-hand side of (A.15).
On the one hand, we have:
$\left|1-\int_{\mathbb{R}^{2}}K_{\ell}(x-u)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}\mathtt{d}u\right|\leq\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\notin\mathrm{B}_{2r}}\mathtt{d}u+\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq|p|}\mathtt{d}u,$
and those terms are controlled by Claims A.3 and A.4. For $|p|\leq\frac{d}{2}$, the error term $\mathcal{O}\left(\frac{1}{\ell d}\right)$ from Claim A.4 is larger than the one from Claim A.3. We get:
$\mathcal{O}(r^{-2}|p|^{2})\times\left(1-\int_{\mathbb{R}^{2}}K_{\ell}(x-u)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}\mathtt{d}u\right)=\mathcal{O}\left(\frac{|p|^{2}}{r^{2}\ell d}\right).$
On the other hand, using Claim A.3, we find:
$\mathcal{O}(r^{-1}|p|)\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq p}\mathtt{d}u=\mathcal{O}(r^{-1}|p|)\times\mathcal{O}\left(\frac{|p|}{\ell d^{2}}\right)=\mathcal{O}\left(\frac{|p|^{2}}{r\ell d^{2}}\right),$
and finally, using Claim A.5, if ${|p|\leq\frac{d}{2}}$:
$\mathcal{O}\left(r^{-3}|p|^{2}\right)\times\int_{\mathbb{R}^{2}}(|u-x|+|p|)\mathsf{1}_{u\in\mathrm{B}_{2r},\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|}|K_{\ell}(x-u)|\mathtt{d}u=\mathcal{O}\left(\frac{|p|^{2}}{r^{3}\ell}\right)+\mathcal{O}\left(\frac{|p|^{2}d}{r^{3}}\right),$
and since $d\geq\bar{d}\geq\ell^{-1}$ the second error term is the largest one.
Inserting those controls into (A.15), taking the expectation, returning to (A.14), and keeping only the dominant error term (using that $d\geq\bar{d}=\sqrt{r\ell^{-1}}$, we have: $\frac{1}{r\ell d^{2}}\geq\frac{d}{r^{3}}\geq\frac{1}{r^{2}\ell d}$), we obtain as desired:
$\mathbb{E}\left[\left(\mathcal{A}(x+p)-\mathcal{A}(x)\right)\mathsf{1}_{|p|\leq\frac{d}{2}}\right]=\mathcal{O}\left(\frac{1}{r\ell d^{2}}\right)+\mathcal{O}(r^{-1})\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right].$
Analysis outside $\mathrm{B}_{2r}$.
Claim A.7.
If $x\notin\mathrm{B}_{2r}$, with $d:=|x|-2r\geq\bar{d}$, we have:
$\mathbb{E}\left[\left(\mathcal{A}(x+p)-\mathcal{A}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\leq{\frac{d}{2}}}\right]=\mathcal{O}\left(\frac{1}{r\ell d^{2}}\right)+\mathcal{O}(r^{-1})\times\mathbb{E}\left[{|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}}\right].$
Proof of Claim A.7.
The analysis here is similar to, and somewhat simpler than the previous case. Since $|p|\leq{\frac{d}{2}}$ the points $x$ and $x+p$ are both outside $\mathrm{B}_{2r}$ hence the difference $\widehat{j_{r}}(x+p)-\widehat{j_{r}}(x)$ is exactly 00 because both terms are 00. On the other hand, we still have:
$\widehat{j_{r}}\ast K_{\ell}(x+p)-\widehat{j_{r}}\ast K_{\ell}(x)=\int_{\mathbb{R}^{2}}\left(\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)\right)K_{\ell}(x-u)\mathtt{d}u,$
so in order to control the difference $\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)$ for $u\in\mathbb{R}^{2}$ we will again distinguish three cases:
  • If $u$ is outside $\mathrm{B}_{2r}$ and $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|$ then both terms are 00 and so is their difference, which we can certainly write as $\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)=0=\nabla\widehat{j_{r}}(u)\cdot p+0$ since $\nabla\widehat{j_{r}}(u)=0$.
  • If $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq|p|$ we can always use the Lipschitz bound on $\widehat{j_{r}}$, which controls the difference by $\mathcal{O}(r^{-1})\times|p|$. For convenience, we write this as:
    $\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)=\nabla\widehat{j_{r}}(u)\cdot p+\mathcal{O}(r^{-1})\times|p|,$
    where in fact both terms in the right-hand side are $\mathcal{O}(r^{-1})\times|p|$.
  • Finally, if $u$ is outside $\mathrm{B}_{2r}$ and $\mathrm{dist}(u,\partial\mathrm{B}_{2r})\geq|p|$, we can use a second-order Taylor’s expansion of $\widehat{j_{r}}$ and write:
    $\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)=\nabla\widehat{j_{r}}(u)\cdot p+\mathcal{O}(r^{-2})\times|p|^{2}.$
Using these informations we get:
$\int_{\mathbb{R}^{2}}\left(\widehat{j_{r}}(u+p)-\widehat{j_{r}}(u)\right)K_{\ell}(x-u)\mathtt{d}u=\int\nabla\widehat{j_{r}}(u)K_{\ell}(x-u)\mathtt{d}u\cdot p\\ +\mathcal{O}(r^{-1})\times|p|\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq|p|}\mathtt{d}u+\mathcal{O}(r^{-2}|p|^{2})\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\in\mathrm{B}_{2r}}\mathtt{d}u.$
As in (A.13), (A.14), when taking the expectation, the term of first order in $p$ vanishes - up to an error controlled as above by $\mathcal{O}(r^{-1})\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]$). The other error terms are handled as in the previous step: on the one hand, using Claim A.3, we have:
$\mathcal{O}(r^{-1})\times|p|\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{\mathrm{dist}(u,\partial\mathrm{B}_{2r})\leq|p|}\mathtt{d}u=\mathcal{O}(r^{-1})\times|p|\times\mathcal{O}\left(\frac{|p|}{\ell d^{2}}\right)=\mathcal{O}\left(\frac{|p|^{2}}{r\ell d^{2}}\right),$
whose expectation is simply $\mathcal{O}\left(\frac{1}{r\ell d^{2}}\right)$, and on the other hand, using Claim A.4:
$\mathcal{O}(r^{-2}|p|^{2})\times\int_{\mathbb{R}^{2}}|K_{\ell}(x-u)|\mathsf{1}_{u\in\mathrm{B}_{2r}}\mathtt{d}u=\mathcal{O}(r^{-2}|p|^{2})\times\mathcal{O}\left(\frac{1}{\ell d}\right)=\mathcal{O}\left(\frac{|p|^{2}}{\ell dr^{2}}\right),$
whose expectation is smaller than the previous error term since $r\geq d$. ∎
Application: proof of Lemma A.1
Combining Claim A.2 with Claim A.6 (for $x\in\mathrm{B}_{2r}$) or Claim A.7 (for $x\notin\mathrm{B}_{2r}$), we obtain that:
$\mathcal{B}_{3}(x)\preceq\frac{1}{r\ell d^{2}}+{r^{-1}\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]}.$
It now remains to compute the sum over $\{x\in\mathcal{L}\setminus\{0\},d\geq\bar{d},|x|\leq 100r\}$.
The first sum is comparable to an integral:
$\sum_{x\in\mathcal{L}\setminus\{0\},d\geq\bar{d},|x|\leq 100r}\frac{1}{r\ell d^{2}}\preceq\frac{1}{\ell r}\times\int_{u=\bar{d}}^{r}\frac{1}{u^{2}}\times r\mathtt{d}u=\mathcal{O}\left(\frac{1}{\ell\bar{d}}\right).$
The second sum can be compared to:
$\sum_{x\in\mathcal{L}\setminus\{0\},d\geq\bar{d},|x|\leq 100r}r^{-1}\times\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{4}}\right]\preceq\int_{s\geq\frac{\bar{d}}{4}}r^{-1}\times\mathbb{E}\left[{|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq s}}\right]\times r\mathtt{d}s\preceq\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{\bar{d}}{4}}\right].$
This concludes the proof of Lemma A.1. ∎
A.2 Close, but not too close to the boundary - Preparation
From now on, we will assume that $0<d\leq\bar{d}$ (recall that $d=d(x)$ denotes the distance from $x$ to the boundary $\partial\mathrm{B}_{2r}$ and that $\bar{d}$ is a lengthscale chosen in (A.1)). We will often use that this implies that $r\leq|x|\leq 3r$.
Preliminaries.
We separate the characteristic function of $\mathrm{p}_{x}-\mathrm{p}_{0}$ into real and imaginary parts and write:
$\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1=\mathbb{E}\left[e^{-it\cdot(\mathrm{p}_{x}-\mathrm{p}_{0}})-1\right]=:\Psi^{r}_{x}(t)-i\Psi^{i}_{x}(t),$
where $\Psi^{r}_{x},\Psi^{i}_{x}$ are the following functions (when the lattice point $x$ that we are considering is unambiguous, we will simply write $p$ instead of $\mathrm{p}_{x}-\mathrm{p}_{0}$):
$\Psi^{r}_{x}(t):=\mathbb{E}\left[\cos(t\cdot p)-1\right],\quad\Psi^{i}_{x}(t):=\mathbb{E}\left[\sin(t\cdot p)\right].$
Clearly $\Psi^{r}_{x}$ is even, $\Psi^{i}_{x}$ is odd. We now record some properties of $\Psi^{r}_{x},\Psi^{i}_{x}$ which will be needed later.
Lemma A.8.
With implicit multiplicative constants depending only on the distribution of a single perturbation, we have:
  • For $|t|\leq 1$:
    $|\Psi^{r}_{x}(t)|+|\Psi^{i}_{x}(t)|=\mathcal{O}(|t|^{2}),\quad|\nabla\Psi^{r}_{x}(t)|+|\nabla\Psi^{i}_{x}(t)|=\mathcal{O}(|t|),\quad|\mathrm{Hess}\Psi^{r}_{x}(t)|+|\mathrm{Hess}\Psi^{i}_{x}(t)|=\mathcal{O}(1),$ (A.16)
  • For $|t|\geq 1$:
    $|\Psi^{r}_{x}(t)|+|\Psi^{i}_{x}(t)|=\mathcal{O}(1),\quad|\nabla\Psi^{r}_{x}(t)|+|\nabla\Psi^{i}_{x}(t)|=\mathcal{O}(1),\quad\quad|\mathrm{Hess}\Psi^{r}_{x}(t)|+|\mathrm{Hess}\Psi^{i}_{x}(t)|=\mathcal{O}(1).$ (A.17)
  • For $\Psi^{r}_{x}$, working in polar coordinates $(\rho,\theta)$:
    $\left|\frac{\partial}{\partial\theta}\Psi^{r}_{x}(\rho,\theta)\right|=\mathcal{O}(\rho^{2}).$ (A.18)
Proof of Lemma A.8.
For $t$ near 00, we can expand $e^{-it\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})}$ as $1-it\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})+\mathcal{O}\left(\frac{|t|^{2}}{2}(\mathrm{p}_{x}-\mathrm{p}_{0})^{2}\right)$, which gives $|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1|=\mathcal{O}(|t|^{2})$ because $(\mathrm{p}_{x}-\mathrm{p}_{0})$ is centered and has finite second moment, and we proceed similarly for the first and second derivatives (using expansions to lower order each time). This yields the bounds (A.16) on $\Psi^{r}_{x},\Psi^{i}_{x}$.
For $|t|$ large, we can use the fact that $\mathrm{p}_{x}-\mathrm{p}_{0}$ has a second moment to see that $\Psi^{r}_{x},\Psi^{i}_{x}$ as well as their first and second derivatives are bounded, which yields (A.17).
It remains to prove (A.18). Working in polar coordinates $t=(\rho\cos\theta,\rho\sin\theta)$, and letting $P$ be the distribution of $\mathrm{p}_{x}-\mathrm{p}_{0}$, we can write:
$\Psi^{r}_{x}(\rho,\theta)=\int_{s>0}\int_{u\in[0,2\pi]}\cos\left(\rho\cos(\theta-u)s\right)s\mathtt{d}P(s,u),$
and taking a derivative with respect to $\theta$ yields (using the inequality $|\sin x|\leq|x|$ in the last step):
$\frac{\partial}{\partial\theta}\Psi^{r}_{x}(\rho,\theta)=-\int_{s>0}\int_{u}\rho s\sin(\theta-u)\sin\left(\rho\cos(\theta-u)s\right)s\mathtt{d}P(s,u)\preceq\rho^{2}\int_{s>0}\int_{u}s^{3}\mathtt{d}P(s,u)=\mathcal{O}(\rho^{2}),$
because $|\mathrm{p}_{x}-\mathrm{p}_{0}|$ has a finite second moment hence $\iint s^{3}\mathtt{d}P(s,u)$ is finite. Note that in this last computation, if $\rho\geq 1$, it would be beneficial to not use $|\sin x|\leq|x|$ and to simply write:
$\frac{\partial}{\partial\theta}\Psi^{r}_{x}(\rho,\theta)\preceq\rho\iint s^{2}\mathtt{d}P(s,u)=\mathcal{O}(\rho),$
however, using the $\mathcal{O}(\rho^{2})$ bound everywhere will be enough for us and this way we avoid distinguishing more cases. ∎
Denote by $\Im$ the imaginary part of a complex number. We will use later the following slight improvement:
Lemma A.9 (A more accurate bound for $\Psi^{i}_{x}$).
For $t\in\mathbb{R}^{2}$ let $\mathcal{K}^{i}_{x}(t):=\Im\left[e^{-it\cdot p}+it\cdot p\right]$. We have:
$|\mathcal{K}^{i}_{x}(t)|\preceq\min\left(|t||p|,|t|^{3}|p|^{3}\right),\quad|\nabla\mathcal{K}^{i}_{x}(t)|\preceq\min\left(|p|,|t|^{2}|p|^{3}\right),\quad|\mathrm{Hess}\mathcal{K}^{i}_{x}(t)|\preceq\min({|p|^{2}},|t||p|^{3}).$ (A.19)
Proof of Lemma A.9.
[…]
Since $p=\mathrm{p}_{x}-\mathrm{p}_{0}$ is centered, we have $\Psi^{i}_{x}(t)=-\mathbb{E}[\mathcal{K}^{i}_{x}(t)]$ for all $t$ and thus the bounds of (A.19) translate into bounds for $\Psi^{i}_{x}$ (as well as its first and second derivatives), which slightly improve on (A.16) because by the dominated convergence theorem, as $t\to 0$:
${\mathbb{E}\left[\min\left(|t||p|,|t|^{3}|p|^{3}\right)\right]=o(|t|^{2}),\quad\mathbb{E}\left[\min\left(|p|,|t|^{2}|p|^{3}\right)\right]=o(|t|)},\quad\mathbb{E}\left[\min(1,|t||p|^{3})\right]=o(1).$
Decomposition on Fourier side.
Recall that:
$\mathcal{B}_{3}(x)=\int_{t\in\mathbb{R}^{2}}{e^{-it\cdot x}}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)(1-\chi_{\ell}(t))\mathtt{d}t\\ =\int_{t\in\mathbb{R}^{2}}\left(\cos(t\cdot x)-i\sin(t\cdot x)\right)j_{r}(t)\left(\Psi^{r}_{x}(t)-i\Psi^{i}_{x}(t)\right)(1-\chi_{\ell}(t))\mathtt{d}t.$
Since the quantity we ultimately want to bound, namely the rescaled number variance, is real, we can focus on controlling the real part of $\mathcal{B}_{3}(x)$. We have:
$\Re(\mathcal{B}_{3}(x))=\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)j_{r}(t)\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t-\int_{t\in\mathbb{R}^{2}}\sin(t\cdot x)j_{r}(t)\Psi^{i}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$
We denote the two integrals in the right-hand side by $\mathcal{B}^{r},-\mathcal{B}^{i}$ (they are both real but one contains $\Psi^{r}_{x}$, the other contains $\Psi^{i}_{x}$). Since $1-\chi_{\ell}$ vanishes on $\{|t|\leq r^{-1}\}$, we may insert the second order asymptotics for the Bessel function $\mathrm{J}_{1}^{2}$ (see (2.9)) and write:
$\mathcal{B}^{r}(x)=\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\frac{\mathrm{J}_{1}^{2}(rt)}{|t|^{2}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t=I^{r}(x)+II^{r}(x)+III^{r}(x),$ (A.20)
where we set:
$I^{r}(x):=a\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\frac{1}{r|t|^{3}}\sin^{2}\left(r|t|-\frac{\pi}{4}\right)\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,\\ II^{r}(x):=b\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\frac{1}{r^{2}|t|^{4}}\cos(2r|t|)\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,\\ III^{r}(x):=\int_{t\in\mathbb{R}^{2}}\mathcal{O}\left(\frac{1}{r^{3}|t|^{5}}\right)\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$ (A.21)
The term $\mathcal{B}^{i}(x)$ has a similar expression, with $\sin(t\cdot x)$ instead of $\cos(t\cdot x)$ and $\Psi^{i}_{x}$ instead of $\Psi^{r}_{x}$, and we decompose it accordingly into a sum of three terms denoted by $I^{i}(x),II^{i}(x),III^{i}(x)$. For convenience, we will refer to the terms $I^{r}(x),I^{i}(x)$ as “$I(x)$” and similarly for $II(x)$, $III(x)$.
Rewriting $I(x),II(x)$ and bounding $III(x)$.
Rewriting $I(x)$.
By elementary trigonometry, we have:
$\sin^{2}\left(r|t|-\frac{\pi}{4}\right)=\frac{1}{2}\left(1-\cos\left(2r|t|-\frac{\pi}{2}\right)\right)=\frac{1}{2}-\frac{1}{2}\sin\left(2r|t|\right),$
and moreover $\cos(t\cdot x)\sin\left(2r|t|\right)=\frac{1}{2}\left(\sin\left(2r|t|+t\cdot x\right)+\sin\left(2r|t|-t\cdot x\right)\right)$, thus we may re-write $I^{r}(x)$ as:
$I^{r}(x)=\frac{a}{2}\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\frac{1}{r|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t-\frac{a}{2}\int_{t\in\mathbb{R}^{2}}\sin\left(2r|t|-t\cdot x\right)\frac{1}{r|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,$ (A.22)
where we used the parity of $\Psi^{r}_{x}$ and $\chi_{\ell}$ in order to combine $\sin\left(2r|t|\pm t\cdot x\right)$ into a single contribution. In the sequel, we will study both integrals in (A.22) separately and denote:
$I^{r}_{1}(x):=\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\frac{1}{r|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,\\ I^{r}_{2}(x):=\int_{t\in\mathbb{R}^{2}}\sin\left(2r|t|-t\cdot x\right)\frac{1}{r|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$ (A.23)
For $I^{i}(x)$, we obtain instead (using this time that $\Psi^{i}_{x}$ is odd):
$I^{i}_{1}(x):=\int_{t\in\mathbb{R}^{2}}\sin(t\cdot x)\frac{1}{r|t|^{3}}\Psi^{i}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,\quad I^{i}_{2}(x):=\int_{t\in\mathbb{R}^{2}}\cos\left(2r|t|-t\cdot x\right)\frac{1}{r|t|^{3}}\Psi^{i}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$ (A.24)
Rewriting $II(x)$.
Still by elementary trigonometry, we have:
$\cos(t\cdot x)\cos(2r|t|)=\frac{1}{2}\left(\cos(2r|t|+t\cdot x)+\cos(2r|t|-t\cdot x)\right).$
Using the parity of $\Psi^{r}_{x}$ and $\chi_{\ell}$ in order to combine both contributions, we may re-write $II^{r}(x)$ as:
$II^{r}(x)=b\int_{t\in\mathbb{R}^{2}}\cos(2r|t|-t\cdot x)\frac{1}{r^{2}|t|^{4}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$ (A.25)
Similarly, for $II^{i}(x)$ we obtain (using this time that $\Psi^{i}_{x}$ is odd):
$II^{i}(x)=b\int_{t\in\mathbb{R}^{2}}\sin(2r|t|-t\cdot x)\frac{1}{r^{2}|t|^{4}}\Psi^{i}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t.$ (A.26)
Remark A.10.
At this point, applying a rough bound is not quite good enough, as one would get, using Lemma A.8:
$|II^{r}(x)|+|II^{i}(x)|\preceq\int_{t\in\mathbb{R}^{2},\ell\leq|t|\leq 1}\frac{1}{r^{2}|t|^{4}}t^{2}\mathtt{d}t+\int_{t\in\mathbb{R}^{2},|t|\geq 1}\frac{1}{r^{2}|t|^{4}}\mathtt{d}t=\mathcal{O}\left(\frac{1}{r^{2}}\log\ell\right),$ (A.27)
and summing over the $\mathcal{O}(r\bar{d})$ lattice points which are at distance $\leq\bar{d}$ from the boundary, we would obtain a control of size $\mathcal{O}\left(\frac{\bar{d}\log\ell^{-1}}{r}\right)$. Althhough we do know that $\ell^{-1}\ll\bar{d}\ll r$, it is not clear that one can take $\ell^{-1}$ and $\bar{d}$ small enough compared to $r$ to ensure that $\bar{d}\log\ell^{-1}\ll r$ (e.g. if $\ell^{-1}$ has to be at least $\frac{r}{\log r}$ then that comparison is false), cf. the discussion after Lemma 3.6, which ultimately governs the choice of $\ell$.
It is worth noticing that if the perturbations were assumed to have a moment of order $2+\delta$ (one could even ask for less), then we could take $\ell$ as some power $r^{-1+\delta^{\prime}}$, and a rough bound like (A.27) would be enough.
Bounding $III(x)$.
We may immediately bound the last term in (A.20), independently of $x$, by:
$III^{r}(x)=\int_{\ell\leq|t|\leq 1}\mathcal{O}\left(\frac{1}{r^{3}|t|^{5}}\right)\times\mathcal{O}(|t|^{2})\mathtt{d}t+\int_{\ell\leq|t|\leq 1}\mathcal{O}\left(\frac{1}{r^{3}|t|^{5}}\right)\times\mathcal{O}(1)\mathtt{d}t=\mathcal{O}(r^{-3}\ell^{-1}),$ (A.28)
and the exact same bound holds for $III^{i}(x)$, using Lemma A.8. In contrast with the previous paragraph, here a rough bound is enough - one can sum $\mathcal{O}(r^{2})$ such terms and still get $\mathcal{O}((r\ell)^{-1})$, which tends to 00 as $r\to\infty$.
Bounding $I_{1}(x)$.
Integrating by parts twice in $I^{r}_{1}(x)$ (see (A.23)) gives:
$I^{r}_{1}(x)=\frac{-1}{r|x|^{2}}\int_{t\in\mathbb{R}^{2}}\cos(t\cdot x)\Delta\left(\frac{1}{|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\right)\mathtt{d}t.$
We now want to evaluate the size of the Laplacian in the integrand. Since $1-\chi_{\ell}$ is supported on $\{|t|\geq\ell\}$, so is that Laplacian, and we make several observations:
For $\frac{1}{2}\ell\leq|t|\leq 1$.
  • The $k$-th derivatives of $1-\chi_{\ell}(t)$ are 00 for $|t|\leq\frac{1}{2}\ell$ ($\chi_{\ell}\equiv 1$) and $|t|\geq\ell$ ($\chi_{\ell}\equiv 0$) and of order $\ell^{-k}$ for $|t|\in[\frac{1}{2}\ell,\ell]$ and $k\geq 0$. Note that for $|t|\in[\frac{1}{2}\ell,\ell]$, the quantities $\ell^{-k}$ and $|t|^{-k}$ are of the same order.
  • The $k$-th derivatives of $\Psi^{r}_{x}$ are of order $|t|^{2-k}$ for $k=0,1,2$, see Lemma A.8.
  • Of course, the $k$-th derivatives of $\frac{1}{|t|^{3}}$ are of order $|t|^{-3-k}$ for $k\geq 0$.
When taking a second derivative of $\frac{1}{|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))$ for $\frac{1}{2}\ell\leq|t|\leq 1$, we may thus lose up to two powers of $|t|$, and the pointwise bound is thus $\mathcal{O}(|t|^{-3+2-2})=\mathcal{O}(|t|^{-3})$.
For $|t|\geq 1$.
  • $1-\chi_{\ell}$ is identically $1$ and its derivatives are thus 00.
  • $\Psi^{r}_{x}$ is bounded and so are its first and second derivatives, see Lemma A.8.
  • The $k$-th derivatives of $\frac{1}{|t|^{3}}$ are still of order $|t|^{-3-k}$ for $k\geq 0$.
When taking a second derivative of $\frac{1}{|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))$ for $|t|\geq 1$, the worst case is thus when both derivatives hit $\Psi^{r}_{x}$, in which case the pointwise bound in $\mathcal{O}(|t|^{-3})$.
Application to bounding $I^{r}_{1}(x)$.
We can thus bound $\Delta\left(\frac{1}{|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\right)$ pointwise by $\mathcal{O}\left(|t|^{-3}\right)$, uniformly with respect to $x\in\mathcal{L}$. We obtain (keeping in mind that $|x|\geq r$):
$I^{r}_{1}(x)\preceq\frac{1}{r|x|^{2}}\int_{t\in\mathbb{R}^{2},|t|\geq\ell}|t|^{-3}\mathtt{d}t=\mathcal{O}\left(\frac{1}{\ell r^{3}}\right).$ (A.29)
The same integration by parts works with $\sin(t\cdot x)$ instead of $\cos(t\cdot x)$, and the estimates of Lemma A.8 also apply to $\Psi^{i}_{x}$, thus $I^{i}_{1}(x)$ can be bounded the same way. Note that the right-hand side of (A.29) is comparable to the error term we found in (A.28) for $III^{r}(x)$. We move on to studying of $I^{r}_{2}(x),I^{i}_{2}(x)$.
A.3 Preliminaries to the study of $I_{2}(x)$
Recall that:
$I^{r}_{2}(x):=\int_{t\in\mathbb{R}^{2}}\sin\left(2r|t|-t\cdot x\right)\frac{1}{r|t|^{3}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,$
and $I^{i}_{2}(x)$ is similar, with $\cos$ instead of $\sin$ and $\Psi^{i}_{x}$ instead of $\Psi^{r}_{x}$.
Using polar coordinates.
We introduce the angle $\theta_{x}$ as an argument of $x$, and re-write $I^{r}_{2}(x)$ (see (A.23)) in polar coordinates as:
$I^{r}_{2}(x)=\frac{1}{r}\int_{\theta\in[0,2\pi]}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\theta-\theta_{x})\right)\frac{1}{\rho^{3}}\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(\rho,\theta)-1\right)(1-\chi_{\ell}(\rho))\rho\mathtt{d}\rho.$
The $\frac{1}{\rho^{3}}$ term coming from the Bessel function and the $\rho$ term coming from (the Jacobian of) polar coordinates can be combined. We also apply a translation $\theta\mapsto\theta+\theta_{x}$ on the angle coordinate, and re-write $I^{r}_{2}(x)$ as:
$I^{r}_{2}(x)=\frac{1}{r}\int_{-\pi}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho.$ (A.30)
The expression for $I^{i}_{2}(x)$ is similar, with $\cos\left(2r\rho-\rho|x|\cos\theta\right)$ instead of $\sin\left(2r\rho-\rho|x|\cos\theta\right)$ and $\Psi^{i}_{x}$ instead of $\Psi^{r}_{x}$.
Auxiliary lengthscales and associated cut-off.
Let $\rho_{\star}$ be some positive length scale and let $\chi_{\rho_{\star}}=\chi\left(\frac{\cdot}{\rho_{\star}}\right)$ be a cut-off at scale $\rho_{\star}$ (for convenience we use the same reference cutoff $\chi$ as in Section 3.1). We will choose $\rho_{\star}$ later, possibly depending on $x,d$ and on the angle coordinate $\theta$. We can then always split $I^{r}_{2}(x),I^{i}_{2}(x)$ in two parts: the “small-$\rho$” one corresponding to $\chi_{\rho_{\star}}$ and the “large-$\rho$” one corresponding to $1-\chi_{\rho_{\star}}$:
$I^{r}_{2}(x)=\frac{1}{r}\int_{-\pi}^{\pi}\mathtt{d}\theta\Bigg{(}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ +\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\Bigg{)},$ (A.31)
and similarly for $I^{i}_{2}(x)$.
After choosing $\rho_{\star}$ depending on $\theta$ we will treat those two parts differently.
  • The “small-$\rho$” term gets evaluated through a rough bound (see (A.32)), except in one case where we will need to look more closely.
  • The “large-$\rho$” term gets integrated by parts (usually twice), and (most of the time) we then apply a rough bound (see (A.34)) on the resulting expression.
Finally, it will remain to integrate those bounds over $\theta$.
Rough bound on the “small-$\rho$” contribution.
Claim A.11.
Let $\rho_{\star}$ be a lengthscale and $\chi_{\rho_{\star}}$ the associated truncation as introduced above. We have:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=\mathcal{O}\left(\min(\rho_{\star},1)\right).$ (A.32)
Proof.
It follows from a direct computation. If $\rho_{\star}\leq 1$, we have (using Lemma A.8):
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\leq\int_{0}^{\rho_{\star}}\mathcal{O}\left(\frac{1}{\rho^{2}}\right)\times\mathcal{O}(\rho^{2})\mathtt{d}\rho=\mathcal{O}(\rho_{\star}),$
uniformly with respect to $\theta$ and $x$. If $\rho_{\star}$ is larger than $1$, we get $\mathcal{O}(1)$ instead of $\mathcal{O}(\rho_{\star})$ in (A.32), because then we can use that $\Psi^{r}_{x}(\rho,\theta)=\mathcal{O}(1)$ (and not $\mathcal{O}(\rho^{2})$) on $\{|t|\in(1,\rho_{\star})\}$. ∎
This rough bound works the same with $\cos$ instead of $\sin$ and replacing $\Psi^{r}_{x}$ by $\Psi^{i}_{x}$, which is also covered by Lemma A.8.
Integration by parts for “large-$\rho$”.
For any fixed $\theta\in[-\pi,\pi]$, we can integrate by parts twice with respect to $\rho$ within the second line of (A.31), which corresponds to “large” values of $\rho$ ($\rho\geq\rho_{\star}$) and write:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ =\frac{-1}{\left(2r-|x|\cos\theta\right)^{2}}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\\ \times\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho.$ (A.33)
For $I^{i}_{2}(x)$, the integration by parts works similarly with $\cos\left(2r\rho-\rho|x|\cos\theta\right)$ instead of $\sin\left(2r\rho-\rho|x|\cos\theta\right)$.
We stress that in (A.33), the lengthscale $\rho_{\star}$ appearing in $\chi_{\rho_{\star}}$ is allowed to depend on $\theta$.
Rough bound on the “large-$\rho$” contribution.
Claim A.12.
Let $\rho_{\star}$ be some lengthscale and $\chi_{\rho_{\star}}$ the associated truncation as introduced above. We have:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho\\ =\mathcal{O}\left(\min\left(\frac{1}{\rho_{\star}},\frac{1}{\ell}\right)\right).$ (A.34)
Proof.
We first study the size of the second derivative with respect to $\rho$ (this is similar to the argument leading to (A.29)). By definition of the cut-off function, the $k$-th derivatives of $1-\chi_{\rho_{\star}}(\rho)$ are 00 for $|\rho|\leq\frac{1}{2}\rho_{\star}$ and $|\rho|\geq\rho_{\star}$ and of order $\rho_{\star}^{-k}$ for $|\rho|\in[\frac{1}{2}\rho_{\star},\rho_{\star}]$ and $k\geq 0$. Note that for $|\rho|\in[\frac{1}{2}\rho_{\star},\rho_{\star}]$, the quantities $\rho_{\star}^{-k}$ and $|\rho|^{-k}$ are of the same order. The same is true for $1-\chi_{\ell}(\rho)$ (with $\ell$ instead of $\rho_{\star}$). Moreover:
For $|\rho|\leq 1$.
  • The $k$-th derivatives of $\Psi^{r}_{x}(\rho,\theta)$ with respect to $\rho$ are of order $|\rho|^{2-k}$ for $k=0,1,2$ (see Lemma A.8).
  • The $k$-th derivatives of $\frac{1}{\rho^{2}}$ are of course of order $\rho^{-2-k}$ for $k\geq 0$.
When taking the second derivative of $\rho\mapsto\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta)(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))$ with respect to $\rho$ for $\rho\leq 1$, we may thus lose up to two powers of $\rho$ in which case the bound is $\mathcal{O}(\rho^{-2})$.
For $|\rho|\geq 1$.
  • The $k$-th derivatives of $\Psi^{r}_{x}$ are bounded for $k=0,1,2$, see Lemma A.8.
  • The $k$-th derivatives of $\frac{1}{\rho^{2}}$ are still of order $\rho^{-2-k}$ for $k\geq 0$.
When taking the second derivative of $\rho\mapsto\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta)(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))$ with respect to $\rho$ for $\rho\geq 1$, the worst case is when both derivatives hit $\Psi^{r}_{x}$, in which case the bound is $\mathcal{O}(\rho^{-2})$.
Application.
Notice that $(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))$ is supported on $\{|\rho|\geq\frac{1}{2}\max(\rho_{\star},\ell)\}$. We just saw that we can bound the second derivative within the integral by $\mathcal{O}(\rho^{-2})$ hence we obtain, uniformly in $\theta$:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho\\ \preceq\int_{\rho=\frac{1}{2}\max(\rho_{\star},\ell)}^{+\infty}\frac{1}{\rho^{2}}\mathtt{d}\rho=\mathcal{O}\left(\min\left(\frac{1}{\rho_{\star}},\frac{1}{\ell}\right)\right).$
∎ The same computation holds for the corresponding part of $I^{i}_{2}(x)$. We now turn to bounding $I^{r}_{2}(x),I^{i}_{2}(x)$.
A.4 Studying $I_{2}(x)$ inside the disk.
Lemma A.13 (Control of $I_{2}(x)$ for $x$ in $\mathrm{B}_{2r}$).
Assume that $x\in B_{2r}$ and that $d:=2r-|x|\leq\bar{d}$. We have:
$|I^{r}_{2}(x)|+|I^{i}_{2}(x)|=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right).$ (A.35)
Proof of Lemma A.13.
Recall that $d\leq\bar{d}\ll r$ implies $|x|\geq r$. We focus on $I^{r}_{2}(x)$ (see (A.30)) - then $I^{i}_{2}(x)$ can be treated exactly the same way.
Preliminary: the $\theta$-dependent term
Recall (see (A.30)) that in the course of integration by parts, a term $\frac{1}{(2r-|x|\cos\theta)^{2}}$ appears. We start by bounding it, depending on the values of $\theta$.
Since $\cos\theta=1-\frac{\theta^{2}}{2}+o(\theta^{2})$ as $\theta\to 0$, there exists $\delta>0$ such that for $|\theta|\leq\delta$, we have $\cos\theta\leq 1-\frac{\theta^{2}}{3}$ and thus we have the following lower bound:
$2r-|x|\cos\theta\geq 2r-|x|\left(1-\frac{\theta^{2}}{3}\right)=d+\frac{|x|}{3}\theta^{2}.$ (A.36)
In particular, we can write:
$\frac{1}{(2r-|x|\cos\theta)^{2}}\preceq\frac{1}{d^{2}}\ \text{ for $|\theta|\leq\sqrt{\frac{d}{r}}$},\quad\frac{1}{(2r-|x|\cos\theta)^{2}}\preceq\frac{1}{r^{2}\theta^{4}}\ \text{ for $|\theta|\geq\sqrt{\frac{d}{r}}$}.$ (A.37)
Moreover, there exists $\varepsilon>0$ such that $\cos\theta\leq 1-\varepsilon$ if $|\theta|\geq\delta$, and thus if $|\theta|\geq\delta$:
$\frac{1}{(2r-|x|\cos\theta)^{2}}\leq\frac{1}{(2r-(1-\varepsilon)|x|)^{2}}\leq\frac{1}{(d+\varepsilon|x|)^{2}}\leq\frac{1}{\varepsilon^{2}r^{2}}.$ (A.38)
We can choose here $\delta,\varepsilon$ globally, not depending on $r$ or any other parameter. Hence, we do not highlight the dependence of the constants on $\delta$ and $\varepsilon.$
We will distinguish three regimes for $\theta$: large, moderate and small.
Large values of $\theta$.
We start by the “large $\theta$” region where $|\theta|\geq\delta$. On this region, we integrate by parts twice the entire $\rho$-integral (which amounts to choosing $\rho_{\star}=\ell$ in (A.31)), we use the rough bound (A.34) together with our estimate (A.38), and write:
$\frac{1}{r}\int_{\theta\in[-\pi,\pi],|\theta|\geq\delta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\times\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\right)\mathtt{d}\rho\\ =\frac{1}{r}\int_{\theta\in[-\pi,\pi],|\theta|\geq\delta}\frac{-1}{\left(2r-|x|\cos\theta\right)^{2}}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\times\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\right)\mathtt{d}\rho\\ =\frac{1}{r}\times\mathcal{O}\left(\frac{1}{r^{2}}\right)\times\mathcal{O}\left(\frac{1}{\ell}\right)=\mathcal{O}\left(\frac{1}{r^{3}\ell}\right).$ (A.39)
Observe that this has the same order of magnitude the bounds on (A.28), (A.29).
Moderate values of $\theta$.
We now consider a “moderate range” $\sqrt{\frac{d}{r}}\leq|\theta|\leq\delta$. We then choose a $\theta$-dependent truncation at scale $\rho_{\star}=\frac{1}{r\theta^{2}}$.
On the one hand, we get, using the rough bound (A.32) on the “small-$\rho$” part:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{r\theta^{2}}\right).$
On the other hand, integrating by parts (see (A.33)), using the rough bound (A.34) and inserting (A.37):
$\frac{1}{\left(2r-|x|\cos\theta\right)^{2}}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\times\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho\\ \preceq\frac{1}{r^{2}\theta^{4}}\times r\theta^{2}=\mathcal{O}\left(\frac{1}{r\theta^{2}}\right).$
We are left to compute an integral over $\theta$:
$\frac{1}{r}\int_{\sqrt{\frac{d}{r}}\leq|\theta|\leq\delta}\frac{1}{r\theta^{2}}\mathtt{d}\theta=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right).$ (A.40)
Small values of $\theta$.
We are left with the range of “small” values $|\theta|\leq\sqrt{\frac{d}{r}}$. We choose a truncation at scale $\rho_{\star}=\frac{1}{d}$. On the one hand, we get, using the rough bound (A.32) on the “small-$\rho$” contribution:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{d}\right),$
and on the other hand, integrating by parts, using the rough bound (A.34) and inserting (A.37):
$\frac{1}{\left(2r-|x|\cos\theta\right)^{2}}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\times\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho\\ \preceq\frac{1}{d^{2}}\times d=\mathcal{O}\left(\frac{1}{d}\right).$
We are left to compute an integral over $\theta$:
$\frac{1}{r}\int_{|\theta|\leq\sqrt{\frac{d}{r}}}\frac{1}{d}\mathtt{d}\theta=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right).$ (A.41)
Conclusion
Combining (A.39), (A.40) and (A.41) (and keeping only the dominant term) yields (A.35) for $I^{r}_{2}(x)$. The term $I^{i}_{2}(x)$ can be treated exactly the same way, as we only relied on the rough bounds of (A.32), (A.34). This proves Lemma A.13. ∎
A.5 Studying $I_{2}(x)$ outside the disk.
Lemma A.14 (Control of $I_{2}(x)$ for $x$ outside $\mathrm{B}_{2r}$).
Assume that $x\in B_{2r}$ and that $d:=|x|-2r\leq\bar{d}$. We have:
$|I^{r}_{2}(x)|+|I^{i}_{2}(x)|=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right).$ (A.42)
Proof of Lemma A.14.
We proceed in a similar fashion as for the previous lemma. The additional difficulty when studying points outside $\partial\mathrm{B}_{2r}$ is that there is always an angle $\bar{\theta}$, defined by:
$\bar{\theta}:=\arccos\left(\frac{2r}{|x|}\right)=\arccos\left(1-\frac{d}{2r+d}\right),$ (A.43)
such that the denominator of $\frac{1}{\left(2r-|x|\cos\theta\right)^{2}}$, which appears when integrating by parts, vanishes for $\theta=\bar{\theta}$. In fact there are two such angles (one in $[0,\pi]$ and one in $[-\pi,0]$), so it is convenient to restrict ourselves to an interval where $\cos$ is one-to-one. We decompose $I^{r}_{2}(x)$, using the parity of $\cos$ and the parity of $\Psi^{r}_{x}(\rho,\cdot)$, as:
$I^{r}_{2}(x)=\frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho\\ +\frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta-\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho.$
and we will focus on bounding the first term - one would proceed identically for the second one. Similarly, we can write (this time $\Psi^{i}_{x}$ is odd):
$I^{i}_{2}(x)=\frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{i}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho\\ -\frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{i}_{x}(\rho,\theta-\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho.$
Our goal is now to bound:
$\frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho,\\ \frac{1}{r}\int_{0}^{\pi}\mathtt{d}\theta\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos\theta\right)\frac{1}{\rho^{2}}\Psi^{i}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho.$ (A.44)
As we will see below, both terms can be treated the same way to some extent, but there is one region of the parameters $(\rho,\theta)$ where we need to study those integrands more carefully, and separately.
Working with $\widetilde{\theta}$.
To emphasize the role of $\bar{\theta}$ as in (A.43), we set $\theta=:\widetilde{\theta}+\bar{\theta}$, so $\widetilde{\theta}$ is our new angle variable, ranging from $-\bar{\theta}$ to $\pi-\bar{\theta}$. Our goal is thus now to study the following quantity (as well as its counterpart):
$\frac{1}{r}\int_{-\bar{\theta}}^{\pi-\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x}+\widetilde{\theta})(1-\chi_{\ell}(\rho))\mathtt{d}\rho,$ (A.45)
and to bound it as claimed in Lemma A.14.
Recall the expansion $\arccos(1-t)=\sqrt{2t}+\mathcal{O}(t^{3/2})$. Since $d\leq\bar{d}\ll r$, we have $\bar{\theta}\sim\sqrt{\frac{d}{r}}$ as $r\to\infty$, and $\sin\bar{\theta}\sim\bar{\theta}\sim\sqrt{\frac{d}{r}}$.
Preliminary: the main $\theta$-dependent term.
We study here not only the term $\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}$ that appears in the integration by parts, but also the argument $2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})$ appearing in (A.45).
We start with an elementary Taylor’s expansion:
$\cos(\bar{\theta}+\widetilde{\theta})=\cos\bar{\theta}-\widetilde{\theta}\sin\bar{\theta}+\mathcal{O}(\widetilde{\theta}^{2})=\cos\bar{\theta}-\widetilde{\theta}\sin\bar{\theta}-\frac{1}{2}\widetilde{\theta}^{2}\cos\bar{\theta}+\mathcal{O}(\widetilde{\theta}^{3}),$ (A.46)
and keep in mind that $\sin\bar{\theta}\sim\sqrt{\frac{d}{r}}$ while $\cos\bar{\theta}\sim 1$ as $r\to\infty$.
We choose $\delta>0$ such that for all $r$ large enough and for all $|\widetilde{\theta}|\leq\delta$, it holds:
$\cos(\bar{\theta}+\widetilde{\theta})\leq\cos\bar{\theta}-\widetilde{\theta}\sin\bar{\theta}-\frac{1}{4}\widetilde{\theta}^{2}.$ (A.47)
We also observe that for $r$ large enough, if $|\widetilde{\theta}|\leq\bar{\theta}\sim\sqrt{\frac{d}{r}}\to 0$ then:
$\cos(\bar{\theta}+\widetilde{\theta})\leq\cos\bar{\theta}-\frac{1}{2}\widetilde{\theta}\sin\bar{\theta}\quad\text{(for $\widetilde{\theta}\geq 0$)},\quad\cos(\bar{\theta}+\widetilde{\theta})\geq\cos\bar{\theta}-\frac{1}{2}\widetilde{\theta}\sin\bar{\theta}\quad\text{(for $\widetilde{\theta}\leq 0$)}.$ (A.48)
Very large values.
If $\delta\leq\widetilde{\theta}\leq\pi-\bar{\theta}$, then $\cos(\bar{\theta}+\widetilde{\theta})\leq\cos(\bar{\theta}+\delta)\leq(1-\varepsilon)\cos\bar{\theta}$ for some $\varepsilon>0$ (depending only on $\delta$). This does not follow from Taylor’s expansions mentioned above, but simply from the identity $\cos(\bar{\theta}+\widetilde{\theta})=\cos(\bar{\theta})\cos(\widetilde{\theta})-\sin(\bar{\theta})\sin(\widetilde{\theta})$, with $\sin(\bar{\theta}),\sin(\widetilde{\theta})\geq 0$ and $\cos(\widetilde{\theta})\leq(1-\varepsilon)$. As a consequence, we have:
$\text{For }\delta\leq\widetilde{\theta}\leq\pi-\bar{\theta},\quad\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}\leq\frac{1}{\left(2r-|x|\times\frac{2r}{|x|}\times(1-\varepsilon)\right)^{2}}=\mathcal{O}\left(\frac{1}{r^{2}}\right).$ (A.49)
Large values.
If $\bar{\theta}\leq\widetilde{\theta}\leq\delta$, we use (A.47) and write:
$2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\geq|x|\left(\widetilde{\theta}\sin\bar{\theta}+\frac{1}{4}\widetilde{\theta}^{2}\right).$
We used that $02r-|x|\cos(\bar{\theta})=$ 0 by definition of $\bar{\theta}$ in (A.43). The quadratic term dominates (because $\sin\bar{\theta}\sim\bar{\theta}\sim\sqrt{\frac{d}{r}}$ and $\widetilde{\theta}\geq\bar{\theta}$), and we get (since $|x|\geq r$):
$\text{For }\bar{\theta}\leq\widetilde{\theta}\leq\delta,\quad\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}=\mathcal{O}\left(\frac{1}{r^{2}\widetilde{\theta}^{2}}\right).$ (A.50)
For simplicity, we combine (A.49) and (A.50) and use (A.50) as a bound on the whole range $\widetilde{\theta}\in[\bar{\theta},\pi-\bar{\theta}]$.
Moderate values.
If $|\widetilde{\theta}|\leq\bar{\theta}$, we use (A.48), distinguish two cases depending on the sign of $\widetilde{\theta}$, and get that:
$\left|2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right|\geq|x|\times\frac{1}{2}\times|\widetilde{\theta}|\sin\bar{\theta}.$
Using that $|x|\geq r$ and that $\sin\bar{\theta}\sim\sqrt{\frac{d}{r}}$, we obtain:
$\text{For }|\widetilde{\theta}|\leq\bar{\theta},\quad\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}=\mathcal{O}\left(\frac{1}{rd\widetilde{\theta}^{2}}\right).$ (A.51)
We can now start evaluating the integrals in (A.44).
Large & very large values ($\bar{\theta}\leq\widetilde{\theta}\leq\pi-\bar{\theta}$).
In this regime, we integrate by parts twice the entire $\rho$-integral in (A.44) (which amounts to choosing $\rho_{\star}=\ell$ in (A.31)) and use the rough bound from (A.34), which gives:
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\right)\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{\ell}\right),$
then we integrate over $\widetilde{\theta}$ using the control (A.50) on $\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}$ (see (A.33)):
$\frac{1}{r}\int_{\bar{\theta}}^{\pi-\bar{\theta}}\frac{1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}\times\mathcal{O}\left(\frac{1}{\ell}\right)\mathtt{d}\widetilde{\theta}=\frac{1}{r}\times\int_{\bar{\theta}}^{\pi-\bar{\theta}}\mathcal{O}\left(\frac{1}{r^{2}\widetilde{\theta}^{2}}\right)\times\mathcal{O}\left(\frac{1}{\ell}\right)\mathtt{d}\widetilde{\theta}=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{r^{3/2}\sqrt{d}\ell}\right),$
using the fact that $\bar{\theta}\sim\sqrt{\frac{d}{r}}$.
Since $\ell r\geq 1$, we can simplify this bound into $\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right)$, which will be enough for us. In conclusion:
$\frac{1}{r}\int_{\bar{\theta}}^{\pi-\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right).$ (A.52)
This computation works identically for the corresponding part of $I^{i}_{2}(x)$, because we are only relying on the rough bounds (A.32), (A.34).
Moderate values: $|\widetilde{\theta}|\leq\bar{\theta}$ (${r}$-part).
This is the most subtle part of the analysis, and the only moment where there is a significant difference of treatment between the “${r}$-part” (term with $\Psi^{r}_{x}$) and the “${i}$-part” (term with $\Psi^{i}_{x}$).
An important observation is that this range of values for $\widetilde{\theta}$ is symmetric around 00. We choose the following truncation:
$\rho_{\star}:=\frac{1}{|\widetilde{\theta}|^{\frac{4}{3}}r^{\frac{2}{3}}d^{\frac{1}{3}}}.$ (A.53)
It is not hard to check that with this choice of $\rho_{\star}$, and for $\widetilde{\theta}\leq\bar{\theta}\sim\sqrt{\frac{d}{r}}$ we have:
$\rho_{\star}r\widetilde{\theta}^{2}\preceq 1,\quad|\widetilde{\theta}|\preceq 1,\quad\rho_{\star}|\widetilde{\theta}|\preceq r\widetilde{\theta}^{2}\rho_{\star}^{2}.$ (A.54)
Small values of $\rho$.
We first consider the part corresponding to $\chi_{\rho_{\star}}$ in (A.44), namely (for fixed $\widetilde{\theta}$):
$\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho.$
In the rest of this paragraph, we may assume that $\rho_{\star}\geq{\frac{1}{2}}\ell$, otherwise this integral vanishes.
Here we do not use the rough bound of (A.32). Instead, we rely on expanding the integrand and on a parity argument to cancel the first order term. First, we expand the oscillating term, using (A.46), as:
$\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)=\sin\left(\rho{|x|}(\widetilde{\theta}\sin\bar{\theta}+\mathcal{O}(\widetilde{\theta}^{2}))\right)=\sin\left(\rho{|x|}\sin\bar{\theta}\cdot\widetilde{\theta}\right)+\mathcal{O}(\rho r\widetilde{\theta}^{2}).$ (A.55)
We also use Lemma A.8 to expand $\Psi^{r}_{x}$ to first order around $\widetilde{\theta}=0$, which yields:
$\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x}+\widetilde{\theta})=\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x})+\widetilde{\theta}\times\mathcal{O}\left(\rho^{2}\right).$ (A.56)
We know a priori that $\Psi^{r}_{x}$ is $\mathcal{O}(\rho^{2})$ (see Lemma A.8). Inserting both expansions (A.55) and (A.56) (the first two inequalities in (A.54) indicate that these are meaningful expansions) into the integral we find:
$\int_{0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ =\int_{0}^{+\infty}\left(\sin\left(\rho{|x|}\sin\bar{\theta}\cdot\widetilde{\theta}\right)+\mathcal{O}(\rho r\widetilde{\theta}^{2})\right)\times\left(\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x})+\widetilde{\theta}\times\mathcal{O}\left(\rho^{2}\right)\right)\times\frac{1}{\rho^{2}}(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ =\int_{\ell}^{\rho_{\star}}\sin\left(\rho{|x|}\sin\bar{\theta}\cdot\widetilde{\theta}\right)\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x})\frac{1}{\rho^{2}}(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ +\int_{\ell}^{\rho_{\star}}\mathcal{O}(\widetilde{\theta})\mathtt{d}\rho+\int_{\ell}^{\rho_{\star}}\mathcal{O}(\rho r\widetilde{\theta}^{2})\mathtt{d}\rho.$
Remark: it might happen that $\rho_{\star}\geq 1$ this is the case if $|\widetilde{\theta}|$ is smaller than $\frac{1}{r^{\frac{1}{2}}d^{\frac{1}{4}}}$ in which case for the part of the integral corresponding to $1\leq\rho\leq\rho_{\star}$, writing that $\Psi^{r}_{x}=\mathcal{O}(\rho^{2})$ and $\partial_{\theta}\Psi^{r}_{x}=\mathcal{O}(\rho^{2})$ is sub-optimal: in view of Lemma A.8 we could in fact use $\Psi^{r}_{x}=\mathcal{O}(1)$ and $\partial_{\theta}\Psi^{r}_{x}=\mathcal{O}(\rho)$. However, the bounds we use here yield a result that is enough for us and we avoid distinguishing more cases.
The function $\widetilde{\theta}\mapsto\int_{\ell}^{\rho_{\star}}\sin\left(\rho{|x|}\sin\bar{\theta}\cdot\widetilde{\theta}\right)\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x})\frac{1}{\rho^{2}}(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho$ is odd (because $\sin$ is, and the dependency of $\rho_{\star}$ is even) and thus (using that the domain of integration is symmetric around 00):
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\ell}^{\rho_{\star}}\sin\left(\rho{|x|}\sin\bar{\theta}\cdot\widetilde{\theta}\right)\Psi^{r}_{x}(\rho,\bar{\theta}+\theta_{x})\frac{1}{\rho^{2}}(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=0.$
On the other hand, we can directly estimate the two other integrals:
$\int_{\ell}^{\rho_{\star}}\widetilde{\theta}\mathtt{d}\rho=\mathcal{O}(\widetilde{\theta}\rho_{\star}),\quad\int_{\ell}^{\rho_{\star}}\rho r\widetilde{\theta}^{2}\mathtt{d}\rho=\mathcal{O}(r\widetilde{\theta}^{2}\rho_{\star}^{2}).$
In view of (A.54), the second term dominates. We are left to compute (using $\bar{\theta}\sim\sqrt{\frac{d}{r}}$):
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}r\widetilde{\theta}^{2}\rho_{\star}^{2}\mathtt{d}\widetilde{\theta}=\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}r\widetilde{\theta}^{2}\frac{1}{|\widetilde{\theta}|^{\frac{8}{3}}r^{\frac{4}{3}}d^{\frac{2}{3}}}\mathtt{d}\widetilde{\theta}\preceq\frac{1}{r}\times\frac{1}{\sqrt{dr}}.$ (A.57)
In conclusion, we have bounded the contribution coming from “small values” of $\rho$ as follows:
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\Psi^{r}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})\frac{1}{\rho^{2}}(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ =\mathcal{O}\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right).$ (A.58)
High values of $\rho$.
We now consider the part corresponding to $1-\chi_{\rho_{\star}}$, we integrate it twice by parts as in (A.33) and we apply the rough bound from (A.34) together with the estimate (A.51). We then integrate this bound over $\widetilde{\theta}$. We get:
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ =\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\frac{-1}{\left(2r-|x|\cos(\bar{\theta}+\widetilde{\theta})\right)^{2}}\times\mathcal{O}\left(\frac{1}{\rho_{\star}}\right)\preceq\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\frac{1}{dr\widetilde{\theta}^{2}}\times\frac{1}{\rho_{\star}}\mathtt{d}\widetilde{\theta}.$
Remark: if $\rho_{\star}$ happens to be smaller than $\ell$ we could use $\mathcal{O}\left(\frac{1}{\ell}\right)$ instead of $\mathcal{O}\left(\frac{1}{\rho_{\star}}\right)$, but we won’t need it.
Our specific choice (A.53) of $\rho_{\star}$ happens to be such that $\frac{1}{dr\widetilde{\theta}^{2}}\times\frac{1}{\rho_{\star}}=r\widetilde{\theta}^{2}\rho_{\star}^{2}$ and we are thus facing (A.57) again. In conclusion, we also have:
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{r}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ =\mathcal{O}\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right).$ (A.59)
Moderate values: $|\widetilde{\theta}|\leq\bar{\theta}$ (“$i$-part”).
Let us now focus on the corresponding range of values for $\widetilde{\theta}$ in the $i$-part of (A.44), we want to prove:
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathtt{d}\widetilde{\theta}\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\Psi^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right).$ (A.60)
One could try to proceed similarly as above, but the parity argument fails to apply in this case. In order to make the necessary gain we will rely instead on Lemma A.9. We start by writing $\Psi^{i}_{x}$ as $-\mathbb{E}[\mathcal{K}^{i}_{x}(\cdot)]$ with the random variable $\mathcal{K}^{i}_{x}(t):=\Im\left[e^{-it\cdot p}+it\cdot p\right]$ as in Lemma A.9, and we recall that (almost surely):
$|\mathcal{K}^{i}_{x}(t)|\preceq\min\left(\rho|p|,\rho^{3}|p|^{3}\right),\quad|\nabla\mathcal{K}^{i}_{x}(t)|\preceq\min\left(|p|,\rho^{2}|p|^{3}\right),\quad|\mathrm{Hess}\mathcal{K}^{i}_{x}(t)|\preceq\min(|p|^{2},\rho|p|^{3}).$ (A.61)
For any $\widetilde{\theta}$ the $\rho$-integral in (A.60) can be re-written as:
[…]
where we simply introduced a distinction according to the respective sizes of $|p|={|\mathrm{p}_{x}-\mathrm{p}_{0}|}$ and $\sqrt{dr}|\widetilde{\theta}|$.
When $|p|$ is large.
To treat the first expectation in (A.62), we work here on the event $\left\{|p|\geq\sqrt{dr}|\widetilde{\theta}|\right\}$. We want to study
[…]
and then take the expectation. We introduce a cut-off at scale
$\rho_{\star}:=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}.$
This divides the $\rho$-integral of (A.63) into “small-$\rho$” and “large-$\rho$” contributions.
Small-$\rho$ for large $|p|$.
In view of (A.61), it is natural to split the “small-$\rho$” contribution further into:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ \leq\int_{\rho=0}^{\frac{1}{|p|}}\frac{1}{\rho^{2}}|\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})|\mathtt{d}\rho+\int_{\rho=\frac{1}{|p|}}^{\rho_{\star}}\frac{1}{\rho^{2}}|\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})|\mathtt{d}\rho.$
We now use the estimates of Lemma A.9 recalled in (A.61) to bound $\mathcal{K}^{i}_{x}$ and obtain (for almost every $p$):
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\\ \preceq\int_{\rho=0}^{\frac{1}{|p|}}\frac{1}{\rho^{2}}\rho^{3}|p|^{3}\mathtt{d}\rho+\int_{\rho=\frac{1}{|p|}}^{\rho_{\star}}\frac{1}{\rho^{2}}\rho|p|\mathtt{d}\rho\preceq|p|+|p|\log(\rho_{\star}|p|),$
the second term being the dominant one when $|p|\geq\frac{1}{\rho_{\star}}$. We may now insert the value of $\rho_{\star}:=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}$, take the expectation and integrate over $\widetilde{\theta}$, which leads us to compute:
$\frac{1}{r}\int_{|\widetilde{\theta}|\leq\bar{\theta}}\mathbb{E}\left[|p|\log\left(\frac{|p|}{|\widetilde{\theta}|\sqrt{dr}}\right)\mathsf{1}_{|p|\geq|\widetilde{\theta}|\sqrt{dr}}\right]\mathtt{d}\widetilde{\theta}=\frac{1}{r}\mathbb{E}\left[\int_{|\widetilde{\theta}|\leq\bar{\theta}}|p|\log\left(\frac{|p|}{|\widetilde{\theta}|\sqrt{dr}}\right)\mathsf{1}_{|p|\geq|\widetilde{\theta}|\sqrt{dr}}\mathtt{d}\widetilde{\theta}\right]\\ \preceq\frac{1}{r}\mathbb{E}\left[\int_{\widetilde{\theta}=0}^{\frac{|p|}{\sqrt{dr}}}|p|\log\left(\frac{|p|}{\widetilde{\theta}\sqrt{dr}}\right)\mathtt{d}\widetilde{\theta}\right]=\frac{1}{r}\frac{1}{\sqrt{dr}}\mathbb{E}\left[|p|^{2}\right]\times\int_{v=0}^{1}-\log v\mathtt{d}v,$
where we used the change of variables $v=\frac{\sqrt{dr}\widetilde{\theta}}{|p|}$. This is indeed $\mathcal{O}\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right)$.
Large-$\rho$ for large $|p|$.
We now look at the “large-$\rho$” contribution in (A.63), which we integrate by parts but this time only once i.e. we write:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ =\frac{1}{2r-|x|\cos(\bar{\theta}+\widetilde{\theta})}\int_{\rho=\frac{1}{2}\rho_{\star}}^{+\infty}\sin\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{\partial}{\partial\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho.$ (A.64)
We can use (A.51) to bound $\frac{1}{2r-|x|\cos(\bar{\theta}+\widetilde{\theta})}$ by $\mathcal{O}\left(\frac{1}{\sqrt{dr}|\widetilde{\theta}|}\right)$ and our next task is to bound the first derivative appearing in the integrand. To do so, we recall that according to Lemma A.9 (as recalled in (A.61)), $\mathcal{K}^{i}_{x}$ is bounded by $\rho|p|$ and $\nabla\mathcal{K}^{i}_{x}$ by $|p|$ almost surely, so we have the pointwise, almost sure bound (compare with the computation leading to (A.34)):
$\frac{\partial}{\partial\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\preceq\frac{1}{\rho^{2}}|p|.$
As a consequence, we get (using (A.64)) an almost sure bound:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ \preceq\frac{1}{\sqrt{dr}|\widetilde{\theta}|}\times\int_{\rho=\frac{1}{2}\rho_{\star}}^{+\infty}\frac{1}{\rho^{2}}|p|\mathtt{d}\rho=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}\times\frac{2}{\rho_{\star}}\times|p|=\mathcal{O}(|p|)$
because we chose $\rho_{\star}=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}$. We can then easily integrate over $|\widetilde{\theta}|\leq\bar{\theta}$:
[…]
(where we have set $v=|\widetilde{\theta}|\sqrt{dr}$), which is again $\mathcal{O}\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right)$.
When $|p|$ is small.
We return to (A.62) and study the second expectation, namely we work here on the event $\left\{|p|\leq\sqrt{dr}|\widetilde{\theta}|\right\}$, we study
[…]
and then take the expectation. We use the same cut-off $\rho_{\star}:=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}$ as before.
Small-$\rho$ for small $|p|$.
Observe that now, if $\rho\leq\rho_{\star}$, then $\rho\leq\frac{1}{|p|}$. Using (A.61), we can thus bound the “small-$\rho$” contribution by:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho\leq\int_{\rho=0}^{\rho_{\star}}\rho|p|^{3}\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{rd\widetilde{\theta}^{2}}\right)\times|p|^{3},$
which we can then integrate over $\widetilde{\theta}$:
[…]
(where we have set $v=|\widetilde{\theta}|\sqrt{dr}$), which is $\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right)$.
Large-$\rho$ for small $|p|$.
Let us finally study the large-$\rho$” regime: there, we perform a double integration by parts and write:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\mathtt{d}\rho\\ =\frac{-1}{(2r-|x|\cos(\bar{\theta}+\widetilde{\theta}))^{2}}\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho.$ (A.66)
In order to estimate the size of the second derivative, we use (A.61) and distinguish two regions:
  1. 1.
    If $\rho_{\star}\leq\rho\leq\frac{1}{|p|}$, the $k^{th}$ derivative of $\mathcal{K}^{i}_{x}$ (with respect to $\rho$) is bounded by $\rho^{3-k}|p|^{3}$ (for $k=0,1,2$) and we get a pointwise bound of the form:
    $\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\preceq\frac{1}{\rho}|p|^{3},$
  2. 2.
    If $\frac{1}{|p|}\leq\rho$, the $k^{th}$ derivative of $\mathcal{K}^{i}_{x}$ (with respect to $\rho$) is bounded by $\rho^{1-k}|p|$ (for $k=0,1$) and $|p|^{2}$ ($k=2$) and we get instead (there are two different terms, depending on where the derivatives hit):
    $\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\preceq{\frac{|p|}{\rho^{3}}+\frac{|p|^{2}}{\rho^{2}}},$
    the second term being the dominant one since $\rho\leq\frac{1}{|p|}$.
Inserting both informations into the integral, we obtain:
$\int_{\rho=0}^{+\infty}\cos\left(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta})\right)\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{2}}\mathcal{K}^{i}_{x}(\rho,\widetilde{\theta}+\bar{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}}(\rho))\right)\mathtt{d}\rho\\ {\preceq\int_{\rho=\rho_{\star}}^{\frac{1}{|p|}}\frac{1}{\rho}|p|^{3}\mathtt{d}\rho+\int_{\rho=\frac{1}{|p|}}^{+\infty}\frac{|p|^{2}}{\rho^{2}}\mathtt{d}\rho=\mathcal{O}\left(|p|^{3}\left(\log\left(\frac{1}{|p|\rho_{\star}}\right)+1\right)\right)},$
Inserting this into (A.66), taking the expectation and integrating over $\widetilde{\theta}$, we obtain:
[…]
where we used the change of variable $w=\frac{\sqrt{dr}\widetilde{\theta}}{|p|}$. This last expression is again $\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right)$.
We have thus proven (A.60).
Conclusion.
Combining (A.52) (which also applies to the “$i$-part”) with either (A.58) and (A.59) (for the “$r$-part”) or (A.60) (for the “$i$-part”), we obtain (A.42), and prove Lemma A.14. ∎
A.6 Bounding $II(x)$
We now turn to studying $II^{r}(x)$ and $II^{i}(x)$. We will prove the following estimate:
Lemma A.15.
Assume that $x$ is such that $d\leq\bar{d}$. We have:
$|II^{r}(x)|+|II^{i}(x)|=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right)+\mathcal{O}\left(\frac{1}{r^{2}}\right)\times\max\left(\log\frac{1}{\ell d},0\right).$ (A.67)
Preliminaries
We want to study the following integral defining $II^{r}(x)$ (see (A.25)):
$II^{r}(x):=\int_{t\in\mathbb{R}^{2}}\cos(2r|t|-t\cdot x)\frac{1}{r^{2}|t|^{4}}\Psi^{r}_{x}(t)(1-\chi_{\ell}(t))\mathtt{d}t,$
and similarly for $II^{i}(x)$ (strictly speaking, there is a multiplicative constant $b$ in front of the integral, which we can safely forget about). The approach here is very close to the analysis of $I_{2}(x)$ done previously, but much shorter.
We use again polar coordinates and write:
$II^{r}(x):=\frac{1}{r^{2}}\int_{0}^{2\pi}\mathtt{d}\theta\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\theta))\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\mathtt{d}\rho.$ (A.68)
We will use truncations $\chi_{\rho_{\star}}$ at certain lengthscales $\rho_{\star}$ as before. The analogue of (A.32) is now:
$\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\theta))\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=\mathcal{O}\left(\max\left(\log\left(\frac{\rho_{\star}}{\ell}\right),0\right)\right),$ (A.69)
and the analogue of (A.34) becomes:
$\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\theta))\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}})(\rho)\right)\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{\rho_{\star}^{2}}\right).$ (A.70)
(Note that we are, in both cases, losing a power of $\rho$ compared to $I^{r}_{2}(x)$).
Points inside $\mathrm{B}_{2r}$.
To simplify, we take the same truncation length $\rho_{\star}=\frac{1}{d}$ for all $\theta$. On the one hand, using (A.69), we get:
$\frac{1}{r^{2}}\int_{0}^{2\pi}\mathtt{d}\theta\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\theta))\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))\chi_{\rho_{\star}}(\rho)\mathtt{d}\rho=\frac{1}{r^{2}}\max\left(\log\left(\frac{1}{\ell d}\right),0\right).$
On the other hand, using (A.70) and the rough bound $\frac{1}{(2r-|x|\cos\theta)^{2}}\leq\frac{1}{d^{2}}$:
$\frac{1}{r^{2}}\int_{0}^{2\pi}\mathtt{d}\theta\frac{1}{(2r-|x|\cos\theta)^{2}}\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\theta))\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\theta+\theta_{x})(1-\chi_{\ell}(\rho))(1-\chi_{\rho_{\star}})(\rho)\right)\mathtt{d}\rho\\ \preceq\frac{1}{r^{2}}\times\frac{1}{d^{2}}\times d^{2}=\mathcal{O}(r^{-2}).$
In conclusion, for points insde $\mathrm{B}_{2r}$ we obtain:
$II^{r}(x)=\mathcal{O}\left(\frac{1}{r^{2}}\right)\times\left(\max\left(\log\left(\frac{1}{\ell d}\right),0\right)+1\right).$ (A.71)
The same analysis holds for $II^{i}(x)$.
Points outside $\mathrm{B}_{2r}$.
The proof goes along the same lines as Section A.5.
Large values ($\widetilde{\theta}\geq\bar{\theta}$).
As before, on this region we integrate by parts twice the entire $\rho$-integral (i.e. we take $\rho_{\star}=\ell$), then use (A.70) (instead of (A.34)), which gives:
$\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta}))\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\right)\mathtt{d}\rho=\mathcal{O}\left(\frac{1}{\ell^{2}}\right),$
and using the control (A.50) on $\frac{1}{(2r-|x|\cos(\bar{\theta}+\widetilde{\theta}))^{2}}$ we compute the integral over $\widetilde{\theta}$ as:
$\frac{1}{r^{2}}\int_{\bar{\theta}}^{\pi-\bar{\theta}}\mathtt{d}\widetilde{\theta}\frac{1}{(2r-|x|\cos(\bar{\theta}+\widetilde{\theta}))^{2}}\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta}))\frac{\partial^{2}}{\partial^{2}\rho}\left(\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))\right)\mathtt{d}\rho\\ \preceq\frac{1}{r^{2}}\int_{\bar{\theta}}^{\pi-\bar{\theta}}\frac{1}{r^{2}\widetilde{\theta}^{2}}\times\frac{1}{\ell^{2}}\mathtt{d}\widetilde{\theta}=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{r^{\frac{5}{2}}d^{\frac{1}{2}}\ell^{2}}\right)=\mathcal{O}\left(\frac{1}{r}\times\frac{1}{\sqrt{dr}}\right),$ (A.72)
using that $r\ell\gg 1$. This should be compared with the computation leading to (A.52).
Moderate values ($|\widetilde{\theta}|\leq\bar{\theta}$).
Compared to what we did for the same regime when studying $I_{2}(x)$, things are simpler and in particular we do not need to expand the integrand or to make any real distinction between “$r$-part” and “$i$-part”. We choose $\rho_{\star}=\frac{1}{\sqrt{dr}|\widetilde{\theta}|}$, and using (A.69), (A.70) and (A.51) we obtain:
$\frac{1}{r^{2}}\int_{0}^{\bar{\theta}}\mathtt{d}\widetilde{\theta}{\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta}))\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))}\mathtt{d}\rho\\ =\frac{1}{r^{2}}\int_{0}^{\bar{\theta}}\max\left(\log\frac{1}{\sqrt{dr}\widetilde{\theta}\ell},0\right)\mathtt{d}\widetilde{\theta}+\frac{1}{r^{2}}\int_{0}^{\bar{\theta}}\frac{1}{rd\widetilde{\theta}^{2}}\times(dr\widetilde{\theta}^{2})\mathtt{d}\widetilde{\theta}.$ (A.73)
(estimating the “small-$\rho$” and “large-$\rho$” contributions separately).
The last integral gives a $\mathcal{O}\left(\frac{1}{r}\times\frac{d^{\frac{1}{2}}}{r^{\frac{3}{2}}}\right)$ contribution, which is negligible compared to $\frac{1}{r}\times\frac{1}{\sqrt{dr}}$ because $d\leq\bar{d}\ll r$, and we forget about it. The first integral in the right-hand side of (A.73) can be evaluated directly:
  • If $d\leq\ell^{-1}$ then the integrand is always equal to $\log\frac{1}{\sqrt{dr}\widetilde{\theta}\ell}$ and we get, after a natural change of variables:
    $\frac{1}{r^{2}}\int_{0}^{\bar{\theta}}\log\frac{1}{\sqrt{dr}\widetilde{\theta}\ell}\mathtt{d}\widetilde{\theta}\preceq\frac{1}{r}\frac{1}{\sqrt{dr}r\ell}\int_{u=0}^{d\ell}-\log u\mathtt{d}u\leq\frac{1}{r}\frac{1}{\sqrt{dr}r\ell}\int_{u=0}^{1}-\log u\mathtt{d}u=o\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right),$
    where we used that $\bar{\theta}\sim\sqrt{\frac{d}{r}}$.
  • If $d\geq\ell^{-1}$, then we should restrict the integral to $\widetilde{\theta}\leq\frac{1}{\sqrt{dr}\ell}$ which gives again:
    $\frac{1}{r^{2}}\int_{0}^{\frac{1}{\sqrt{dr}\ell}}\log\frac{1}{\sqrt{dr}\widetilde{\theta}\ell}\mathtt{d}\widetilde{\theta}=\frac{1}{r}\frac{1}{\sqrt{dr}r\ell}\int_{u=0}^{1}-\log u\mathtt{d}u=o\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right).$
So in conclusion, we have:
${\frac{1}{r^{2}}\int_{0}^{\bar{\theta}}\mathtt{d}\widetilde{\theta}{\int_{0}^{+\infty}\cos(2r\rho-\rho|x|\cos(\bar{\theta}+\widetilde{\theta}))\frac{1}{\rho^{3}}\Psi^{r}_{x}(\rho,\bar{\theta}+\widetilde{\theta}+\theta_{x})(1-\chi_{\ell}(\rho))}\mathtt{d}\rho}=o\left(\frac{1}{r}\frac{1}{\sqrt{dr}}\right).$ (A.74)
The same computation can be done for the “$i$-part”.
Conclusion
Using (A.71) (if $x\in\mathrm{B}_{2r}$) or (A.72) and (A.74) (if $x\notin\mathrm{B}_{2r}$) we obtain (A.67) and prove Lemma A.15.
A.7 Conclusion: proof of Proposition 3.7
Points that are far from the boundary
By Lemma A.1, we have (using the value $\bar{d}=\sqrt{r\ell^{-1}}$):
$\sum_{x\in\mathcal{L}\setminus\{0\},d\geq\bar{d},|x|\leq 100r}\mathcal{B}_{3}(x)\preceq\frac{1}{\sqrt{r\ell}}+\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{\bar{d}}{2}}\right].$ (A.75)
Points that are close, but not too close to the boundary
Claim A.16.
We have:
$\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}\mathcal{B}_{3}(x)\preceq\frac{1}{(r\ell)^{\frac{1}{4}}}+\frac{1}{r^{0.45}}.$ (A.76)
Proof of Claim A.16.
By (A.20) it is enough to study the sums of $I(x),II(x),III(x)$ separately. Let us start with the latter: using (A.28) and the fact that there are $\mathcal{O}(r^{2})$ points at distance $\leq\bar{d}$ from the boundary, we have:
$\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}III(x)\preceq r^{2}\times r^{-3}\ell^{-1}=\frac{1}{r\ell}.$
We split $I(x)$ further into $I_{1}(x)+I_{2}(x)$. Using (A.29) we see already that $\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}I_{1}(x)\preceq\frac{1}{r\ell}$.
Sum of $I_{2}(x)$.
Using Lemma A.13 and Lemma A.14 we have:
$\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}|I^{r}_{2}(x)|+|I^{i}_{2}(x)|\preceq\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}\frac{1}{r}\times\frac{1}{\sqrt{dr}}.$ (A.77)
First, we restrict ourselves to $d\geq 1$ and compare that part of the sum in to an integral:
$\sum_{x\in\mathcal{L},1\leq d\leq\bar{d}}\frac{1}{r}\times\frac{1}{\sqrt{dr}}\preceq\int_{x=1}^{\bar{d}}\frac{1}{\sqrt{xr}}\mathtt{d}x=\mathcal{O}\left(\sqrt{\frac{\bar{d}}{r}}\right)=\mathcal{O}\left(\frac{1}{(r\ell)^{\frac{1}{4}}}\right)\to 0.$
The part of the sum (A.77) corresponding to $\tau\leq d\leq 1$ can be estimated by saying that there are $\mathcal{O}(r)$ lattice points in that annulus and thus, with our choice of $\tau=r^{-0.1}\gg r^{-1}$:
$\sum_{x\in\mathcal{L},\tau\leq d\leq 1}\frac{1}{r}\times\frac{1}{\sqrt{dr}}\preceq\frac{1}{r^{0.45}}.$
This concludes the evaluation of the sum of $I_{2}(x)$.
Sum of $II(x)$.
We use Lemma A.15 and write:
$\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}{|II^{r}(x)|+|II^{i}(x)|}\preceq\sum_{x\in\mathcal{L},\tau\leq d\leq\bar{d}}\frac{1}{r}\times\frac{1}{\sqrt{dr}}+\frac{1}{r^{2}}\max\left(\log\frac{1}{\ell d},0\right).$
The first part of the sum can be treated exactly as above, the only annoying new term is the $\log$ one. This term only appears if $d\leq\ell^{-1}$ and by comparison to an integral we have:
$\frac{1}{r^{2}}\sum_{x\in\mathcal{L},1\leq d\leq\frac{1}{\ell}}\log\frac{1}{\ell d}\preceq\frac{1}{r}\int_{x=1}^{\ell^{-1}}-\log(x\ell)\mathtt{d}x=\mathcal{O}\left(\frac{1}{r\ell}\right).$
In this last sum we only considered $d\geq 1$ so we are left with:
$\sum_{x\in\mathcal{L},\tau\leq d\leq 1}\frac{1}{r^{2}}\times\log\frac{1}{\ell d}\preceq r\times\frac{1}{r^{2}}\times\log\left(\frac{r}{\tau}\right)=\mathcal{O}\left(\frac{\log(r)}{r}\right),$
using our choice of $\tau\gg r^{-1}$.
Collecting all error terms and discarding negligible ones, we obtain (A.76), which proves Claim A.16. ∎
Points that are very close to the boundary
To treat the contribution of the sum of $\mathcal{B}_{3}(x)$ for $d\leq\tau$, we rely on the “number-theoretic” result stated in Lemma 2.2 to count those lattice points and on the rough bound (3.24). We obtain, since $\tau=r^{-0.1}\ll 1$:
$\sum_{x\in\mathcal{L},d\leq\tau}\mathcal{B}_{3}(x)=\left(\mathcal{O}(\tau r)+\mathcal{O}(r^{1-\gamma})\right)\times\mathcal{O}(r^{-1})=\mathcal{O}(r^{-\delta}),$ (A.78)
for $\delta=\min(\gamma,0.1)>0$.
Combining (A.75), (A.76) and (A.78), and discarding negligible terms, we obtain (for some $\delta>0$):
$\sum_{x\in\mathcal{L},x\neq 0,|x|\leq 100r}\mathcal{B}_{3}(x)\preceq\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{\bar{d}}{2}}\right]+r^{-\delta}\preceq\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\mathsf{1}_{|\mathrm{p}_{0}|\geq\ell^{-1}}\right]+r^{-\delta},$
using $\bar{d}\gg\ell^{-1}$, which gives (3.25) and proves Proposition 3.7.
Appendix B The case of dimension 1
Proposition B.1.
If $\mathbb{E}[|\mathrm{p}_{0}|]$ is finite, then $\mathbf{PL}$ is hyperuniform. If $\mathbb{E}[|\mathrm{p}_{0}|^{2}]$ is finite, then $\mathbf{PL}$ is class-I hyperuniform.
Proof.
We start as in the proof of Proposition 3.1 by writing:
$\frac{\mathrm{Var}\left[|\mathbf{PL}\cap\mathrm{B}_{r}|\right]}{|\mathrm{B}_{r}|}=\mathbb{E}\sum_{x\in\mathcal{L}\setminus\{0\}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})+1-|\mathrm{B}_{r}|=A+B,$
where $A$ is a deterministic contribution coming from the lattice $\mathbb{Z}$ and $B$ is the “error term” due to the perturbations:
$A:=\sum_{x\in\mathbb{Z}\setminus\{0\}}\widehat{j_{r}}(x)+1-|\mathrm{B}_{r}|,\quad B:=\mathbb{E}\sum_{x\in\mathbb{Z}\setminus\{0\}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x).$
The term $A$ is $\mathcal{O}(r^{-1})$ by the same computation as in the $\mathsf{d}=2$ case and we focus on $B$.
We first use Lemma 2.9 and write:
$B=\mathbb{E}\sum_{x\in\mathbb{Z}\setminus\{0\},|x|\leq 100r}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)+\mathcal{O}\left(\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq 10r}\right]\right).$ (B.1)
Note that if $\mathbb{E}[|\mathrm{p}_{0}|]$ is finite then the error term tends to 00 as $r\to\infty$ and if $\mathbb{E}[|\mathrm{p}_{0}|^{2}]$ is finite, then by Markov’s inequality it is actually $\mathcal{O}(r^{-1})$.
The function $\widehat{j_{r}}$ is a “tent” supported on $[-2r,2r]$ and whose expression is given by $\widehat{j_{r}}(t)=1-\frac{|t|}{2r}$ on that interval. In particular, it is continuous, $\mathcal{O}(r^{-1})$-Lipschitz and piecewise affine, with a lack of differentiability at 00 and ±$\pm 2r$. For $x\in\mathbb{Z}$, we let $d:=\mathrm{dist}\left(x,\{0,-2r,2r\}\right)$. We introduce some lengthscale $\bar{d}$ to be chosen later with $1\leq\bar{d}\ll r$. First, using the Lipschitz bound on $\widehat{j_{r}}$ we have:
$\mathbb{E}\sum_{x\in\mathbb{Z},d\leq\bar{d}}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\preceq\frac{\bar{d}}{r}\times\mathbb{E}[|\mathrm{p}_{0}|].$ (B.2)
Next, if $x$ is such that $d\geq\bar{d}$, we write:
$\mathbb{E}[\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)]=\mathbb{E}\left[\left(\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\leq d}\right]+\mathbb{E}\left[\left(\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq d}\right].$
Since $\widehat{j_{r}}$ is affine on $(x-d,x+d)$ with slope $\mathcal{O}(r^{-1})$, and $\mathrm{p}_{x}-\mathrm{p}_{0}$ is centered, we have:
$\mathbb{E}\left[\left(\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\leq d}\right]\leq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq d}\right]\preceq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{2}}\right].$
It remains to see that, since $\widehat{j_{r}}$ is $\mathcal{O}(r^{-1})$-Lipschitz:
$\mathbb{E}\left[\left(\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\right)\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq d}\right]\preceq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{x}-\mathrm{p}_{0}|\geq d}\right]\preceq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{2}}\right],$ (B.3)
from which we can derive the following simple bound:
$\mathbb{E}\sum_{x\in\mathbb{Z},d\geq\bar{d},|x|\leq 100r}\widehat{j_{r}}(x+\mathrm{p}_{x}-\mathrm{p}_{0})-\widehat{j_{r}}(x)\preceq\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{\bar{d}}{2}}\right].$ (B.4)
If $\mathbb{E}[|\mathrm{p}_{0}|]$ is finite, the right-hand side goes to 00 as soon as $\bar{d}\to\infty$. On the other hand, if $\mathbb{E}[|\mathrm{p}_{0}|^{2}]$ is finite, we might prefer to return to (B.3) and compute instead the sum:
$\frac{1}{r}\sum_{x\in\mathbb{Z},|x|\leq 100r,d\geq\bar{d}}\mathbb{E}\left[|\mathrm{p}_{0}|\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{2}}\right]\preceq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{0}|\sum_{d=\bar{d}}^{100r}\mathsf{1}_{|\mathrm{p}_{0}|\geq\frac{d}{2}}\right]\preceq\frac{1}{r}\mathbb{E}\left[|\mathrm{p}_{0}|^{2}\right]=\mathcal{O}(r^{-1}).$ (B.5)
In summary, if $\mathbb{E}[|\mathrm{p}_{0}|]$ is finite, we can choose $1\ll\bar{d}\ll r$ in (B.2), (B.4) and see that the term $B$ goes to 00 as $r\to\infty$, so $\mathbf{PL}$ is hyperuniform. If moreover $\mathbb{E}[|\mathrm{p}_{0}|^{2}]$ is finite then we can choose $\bar{d}=1$ in (B.2), (B.5) and we find that $B=\mathcal{O}(r^{-1})$ thus $\mathbf{PL}$ is class-I hyperuniform. ∎
Acknowledgement
DD and DF were supported in part by the Labex CEMPI (ANR-11-LABX-0007-01), the ANR project RANDOM (ANR-19-CE24-0014) and by the CNRS GdR 3477 GeoSto.
DF is supported by the Czech Science Foundation, project no. 22-15763S and thanks the University of Lille for its hospitality.
MH is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265 Random Geometric Systems as well as under Germany’s Excellence Strategy EXC 2044 -390685587, Mathematics Münster: Dynamics–Geometry–Structure.
TL acknowledges the support of JCJC grant ANR-21-CE40-0009 from Agence Nationale de la Recherche, as well as the hospitality of Mathematics Münster, where this project was initiated.